Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0002.lab vs ../HVITE-M/Rec-mau-M-24/mfG0002.rec
LAB: pau k o0104 u0204 e0304 N0404 pau
REC: pau r o0103 N0203 r     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0005.lab vs ../HVITE-M/Rec-mau-M-24/mfG0005.rec
LAB: pau s o0103 m a0203 ts u0303 pau
REC: pau k o0103 w a0203 s  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0006.lab vs ../HVITE-M/Rec-mau-M-24/mfG0006.rec
LAB: pau z o0104 k u0204 z o0304 k u0404 pau
REC: pau         k o0103 d o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0007.lab vs ../HVITE-M/Rec-mau-M-24/mfG0007.rec
LAB: pau s a0104 z u0204 k a0304 r u0404 pau
REC: pau   a0104 z u0204 k a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0008.lab vs ../HVITE-M/Rec-mau-M-24/mfG0008.rec
LAB: pau k a0103 t a0203 i0303 pau
REC: pau t a0103 g a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0009.lab vs ../HVITE-M/Rec-mau-M-24/mfG0009.rec
LAB: pau          d a0102 i0202 pau
REC: pau sh i0103 g a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0010.lab vs ../HVITE-M/Rec-mau-M-24/mfG0010.rec
LAB: pau   a0106 w a0206 t a0306 d a0406 sh i0506 i0606 pau
REC: pau s a0105 w         a0205 g a0305 sh i0405 i0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0011.lab vs ../HVITE-M/Rec-mau-M-24/mfG0011.rec
LAB: pau d a0103 m a0203 s  u0303 pau
REC: pau n a0103 n a0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0012.lab vs ../HVITE-M/Rec-mau-M-24/mfG0012.rec
LAB: pau s e0104 i0204 h i0304 N0404 pau
REC: pau   g     e0103 h i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0013.lab vs ../HVITE-M/Rec-mau-M-24/mfG0013.rec
LAB: pau k e0104 i0204 y a0304 k u0404 pau
REC: pau s e0104 i0204 y a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0015.lab vs ../HVITE-M/Rec-mau-M-24/mfG0015.rec
LAB: pau   o0103 k o0203 r u0303 pau
REC: pau k o0103   u0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0016.lab vs ../HVITE-M/Rec-mau-M-24/mfG0016.rec
LAB: pau h o0103   ky    u0203 u0303 pau
REC: pau t o0103 k e0203 r     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0017.lab vs ../HVITE-M/Rec-mau-M-24/mfG0017.rec
LAB: pau               i0103 q0203 s o0303 pau

REC: pau k i0104 N0204 z     o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0018.lab vs ../HVITE-M/Rec-mau-M-24/mfG0018.rec
LAB: pau k a0103 s e0203 g u0303 pau
REC: pau k a0103 k e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0020.lab vs ../HVITE-M/Rec-mau-M-24/mfG0020.rec
LAB: pau   ts    u0103 i0203 t e0303 pau
REC: pau m i0103 j     i0203 m e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0021.lab vs ../HVITE-M/Rec-mau-M-24/mfG0021.rec
LAB: pau               h e0102 N0202 pau
REC: pau g e0104 N0204 k i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0023.lab vs ../HVITE-M/Rec-mau-M-24/mfG0023.rec
LAB: pau   o0102   u0202 pau
REC: pau t o0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0024.lab vs ../HVITE-M/Rec-mau-M-24/mfG0024.rec
LAB: pau         m i0101 pau
REC: pau g o0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0025.lab vs ../HVITE-M/Rec-mau-M-24/mfG0025.rec
LAB: pau h a0104 i0204 b o0304 k u0404 pau
REC: pau g a0104 i0204 k o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0026.lab vs ../HVITE-M/Rec-mau-M-24/mfG0026.rec
LAB: pau n e0103 z u0203 m i0303 pau
REC: pau h i0103 z u0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0028.lab vs ../HVITE-M/Rec-mau-M-24/mfG0028.rec
LAB: pau         k a0102 d o0202 pau
REC: pau t a0103 r a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0030.lab vs ../HVITE-M/Rec-mau-M-24/mfG0030.rec
LAB: pau n a0104 n i0204 sh i0304 r     o0404 pau
REC: pau g a0104   i0204    ch    u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0031.lab vs ../HVITE-M/Rec-mau-M-24/mfG0031.rec
LAB: pau k i0103 g e0203 N0303 pau
REC: pau m i0103 r e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0032.lab vs ../HVITE-M/Rec-mau-M-24/mfG0032.rec
LAB: pau s a0106 sh i0206 ts u0306 k a0406 e0506 r u0606 pau
REC: pau                     a0104 t a0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0034.lab vs ../HVITE-M/Rec-mau-M-24/mfG0034.rec
LAB: pau               s       u0101 pau
REC: pau k o0104 u0204 s e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0035.lab vs ../HVITE-M/Rec-mau-M-24/mfG0035.rec
LAB: pau         d o0103 u0203 i0303 pau

REC: pau h i0103 d       o0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0036.lab vs ../HVITE-M/Rec-mau-M-24/mfG0036.rec
LAB: pau k o0104 q0204 k     o0304 u0404 pau
REC: pau   j     o0103 u0203 h     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0037.lab vs ../HVITE-M/Rec-mau-M-24/mfG0037.rec
LAB: pau h i0105 k i0205 k     a0305 e0405 s u0505 pau
REC: pau h a0104 k       i0204 d     a0304 s u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0038.lab vs ../HVITE-M/Rec-mau-M-24/mfG0038.rec
LAB: pau d a0103 m a0203 s u0303 pau
REC: pau h a0103 n a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0039.lab vs ../HVITE-M/Rec-mau-M-24/mfG0039.rec
LAB: pau g e0103 N0203 d       o0303 pau
REC: pau m e0104 N0204 d o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0040.lab vs ../HVITE-M/Rec-mau-M-24/mfG0040.rec
LAB: pau ts u0103 k         u0203 e0303 pau
REC: pau          k i0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0041.lab vs ../HVITE-M/Rec-mau-M-24/mfG0041.rec
LAB: pau   o0103 g a0203 m         u0303 pau
REC: pau t o0104 g a0204 m e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0042.lab vs ../HVITE-M/Rec-mau-M-24/mfG0042.rec
LAB: pau t a0103 t             o0203 e0303 pau
REC: pau t a0104 t o0204 e0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0044.lab vs ../HVITE-M/Rec-mau-M-24/mfG0044.rec
LAB: pau               h i0102 r u0202 pau
REC: pau u0104 r a0204 g i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0045.lab vs ../HVITE-M/Rec-mau-M-24/mfG0045.rec
LAB: pau ch u0104 u0204 sh i0304 N0404 pau
REC: pau    k     u0103 sh i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0046.lab vs ../HVITE-M/Rec-mau-M-24/mfG0046.rec
LAB: pau s a0104 N0204 g a0304 ts u0404 pau
REC: pau d a0104 N0204   a0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0047.lab vs ../HVITE-M/Rec-mau-M-24/mfG0047.rec
LAB: pau k i0103 k     i0203   N0303 pau
REC: pau k       a0103 i0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0048.lab vs ../HVITE-M/Rec-mau-M-24/mfG0048.rec
LAB: pau    sh    o0103 y u0203 u0303 pau
REC: pau ky o0104 u0204 y o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0050.lab vs ../HVITE-M/Rec-mau-M-24/mfG0050.rec
LAB: pau k u0103 r u0203 u0303 pau

REC: pau t e0102 r       u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0053.lab vs ../HVITE-M/Rec-mau-M-24/mfG0053.rec
LAB: pau n i0104 N0204 t a0304 i0404 pau
REC: pau         i0103 g a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0054.lab vs ../HVITE-M/Rec-mau-M-24/mfG0054.rec
LAB: pau   i0102   N0202 pau
REC: pau h i0102 r u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0056.lab vs ../HVITE-M/Rec-mau-M-24/mfG0056.rec
LAB: pau k o0104 q0204 k         a0304 k u0404 pau
REC: pau g o0104       k u0204 r a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0057.lab vs ../HVITE-M/Rec-mau-M-24/mfG0057.rec
LAB: pau k a0103 g         a0203 k u0303 pau
REC: pau n a0104 g a0204 b i0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0059.lab vs ../HVITE-M/Rec-mau-M-24/mfG0059.rec
LAB: pau k a0104 N0204 k a0304 k u0404 pau
REC: pau   d     a0103 r a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0060.lab vs ../HVITE-M/Rec-mau-M-24/mfG0060.rec
LAB: pau         t a0102 r u0202 pau
REC: pau k u0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0061.lab vs ../HVITE-M/Rec-mau-M-24/mfG0061.rec
LAB: pau s       a0103 N0203 p     o0303 pau
REC: pau s a0104 N0204 k     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0063.lab vs ../HVITE-M/Rec-mau-M-24/mfG0063.rec
LAB: pau       m a0102 r u0202 pau
REC: pau u0103 m a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0064.lab vs ../HVITE-M/Rec-mau-M-24/mfG0064.rec
LAB: pau         sh i0102 r  i0202 pau
REC: pau k i0103 j  i0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0066.lab vs ../HVITE-M/Rec-mau-M-24/mfG0066.rec
LAB: pau b e0103 N0203 ch i0303 pau
REC: pau g e0103 N0203 sh i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0067.lab vs ../HVITE-M/Rec-mau-M-24/mfG0067.rec
LAB: pau a0103 N0203 j i0303 pau
REC: pau k     a0102 j i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0068.lab vs ../HVITE-M/Rec-mau-M-24/mfG0068.rec
LAB: pau m a0104 g o0204 ts u0304 k u0404 pau
REC: pau s u0103 g          o0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0069.lab vs ../HVITE-M/Rec-mau-M-24/mfG0069.rec
LAB: pau           h     a0102 i0202 pau

REC: pau t a0103 r e0203 r     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0072.lab vs ../HVITE-M/Rec-mau-M-24/mfG0072.rec
LAB: pau       t a0102 r u0202 pau
REC: pau u0103 m a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0077.lab vs ../HVITE-M/Rec-mau-M-24/mfG0077.rec
LAB: pau   o0104 k  u0204 r e0304 r u0404 pau
REC: pau m o0104 ts u0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0078.lab vs ../HVITE-M/Rec-mau-M-24/mfG0078.rec
LAB: pau d o0103 u0203 z       o0303 pau
REC: pau k o0104 u0204 z o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0079.lab vs ../HVITE-M/Rec-mau-M-24/mfG0079.rec
LAB: pau   n     i0102 w     a0202 pau
REC: pau k i0103 b     a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0080.lab vs ../HVITE-M/Rec-mau-M-24/mfG0080.rec
LAB: pau   i0103 ry o0203 k     u0303 pau
REC: pau r i0103    y     u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0081.lab vs ../HVITE-M/Rec-mau-M-24/mfG0081.rec
LAB: pau j i0104 q0204 s e0304 k i0404 pau
REC: pau   r     i0103 r e0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0083.lab vs ../HVITE-M/Rec-mau-M-24/mfG0083.rec
LAB: pau u0102 sh         i0202 pau
REC: pau i0103 sh i0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0084.lab vs ../HVITE-M/Rec-mau-M-24/mfG0084.rec
LAB: pau               d a0102 N0202 pau
REC: pau k a0104 N0204 t e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0087.lab vs ../HVITE-M/Rec-mau-M-24/mfG0087.rec
LAB: pau   sh    i0103 ky u0203 u0303 pau
REC: pau h a0104 i0204 ky u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0089.lab vs ../HVITE-M/Rec-mau-M-24/mfG0089.rec
LAB: pau h o0103 ky u0203   u0303 pau
REC: pau m o0103 g  u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0090.lab vs ../HVITE-M/Rec-mau-M-24/mfG0090.rec
LAB: pau n i0103 g a0203 s u0303 pau
REC: pau k u0103 r a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0091.lab vs ../HVITE-M/Rec-mau-M-24/mfG0091.rec
LAB: pau sh i0104 N0204 k o0304 u0404 pau
REC: pau k  i0104 N0204 k o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0093.lab vs ../HVITE-M/Rec-mau-M-24/mfG0093.rec
LAB: pau ch u0104   u0204 d a0304   N0404 pau

REC: pau f  u0104 k u0204 r a0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0094.lab vs ../HVITE-M/Rec-mau-M-24/mfG0094.rec
LAB: pau s a0103 b o0203 r u0303 pau
REC: pau k a0103 b u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0095.lab vs ../HVITE-M/Rec-mau-M-24/mfG0095.rec
LAB: pau k a0105 t a0205 z u0305 k e0405 r u0505 pau
REC: pau         n a0104 z u0204 k e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0096.lab vs ../HVITE-M/Rec-mau-M-24/mfG0096.rec
LAB: pau t a0103 N0203 b       o0303 pau
REC: pau d a0104 N0204 b o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0097.lab vs ../HVITE-M/Rec-mau-M-24/mfG0097.rec
LAB: pau j i0103 N0203 sh    u0303 pau
REC: pau k i0103 j     u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0098.lab vs ../HVITE-M/Rec-mau-M-24/mfG0098.rec
LAB: pau s a0103 m     a0203 s     u0303 pau
REC: pau k a0104 a0204 t     e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0099.lab vs ../HVITE-M/Rec-mau-M-24/mfG0099.rec
LAB: pau i0104 q0204 t e0304 i0404 pau
REC: pau k     i0103 g a0203 e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0100.lab vs ../HVITE-M/Rec-mau-M-24/mfG0100.rec
LAB: pau           sh    o0103 b     u0203 N0303 pau
REC: pau k u0104 z u0204 r     e0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0101.lab vs ../HVITE-M/Rec-mau-M-24/mfG0101.rec
LAB: pau sh    i0102 r     o0202 pau
REC: pau u0103 d     o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0103.lab vs ../HVITE-M/Rec-mau-M-24/mfG0103.rec
LAB: pau         ts u0102 ch i0202 pau
REC: pau k i0103 r  o0203 k  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0104.lab vs ../HVITE-M/Rec-mau-M-24/mfG0104.rec
LAB: pau   o0103 g a0203 m         u0303 pau
REC: pau t o0104 g a0204 m e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0107.lab vs ../HVITE-M/Rec-mau-M-24/mfG0107.rec
LAB: pau u0105 s u0205 g u0305 r a0405 i0505 pau
REC: pau         u0103 g         a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0110.lab vs ../HVITE-M/Rec-mau-M-24/mfG0110.rec
LAB: pau i0105 r a0205 q0305 sh                 a0405 r u0505 pau
REC: pau         t     a0105 sh i0205 k a0305 m e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0115.lab vs ../HVITE-M/Rec-mau-M-24/mfG0115.rec
LAB: pau          ky    o0102 h     i0202 pau

REC: pau ry o0104 u0204 sh    i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0116.lab vs ../HVITE-M/Rec-mau-M-24/mfG0116.rec
LAB: pau h a0104 k i0204 d     a0304 s  u0404 pau
REC: pau d a0104         N0204 a0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0118.lab vs ../HVITE-M/Rec-mau-M-24/mfG0118.rec
LAB: pau k o0104 k u0204 f u0304 k u0404 pau
REC: pau g u0104 r u0204   u0304 p u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0119.lab vs ../HVITE-M/Rec-mau-M-24/mfG0119.rec
LAB: pau   sh    i0102 r     o0202 pau
REC: pau j i0103 d     o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0120.lab vs ../HVITE-M/Rec-mau-M-24/mfG0120.rec
LAB: pau         h     a0102 sh    i0202 pau
REC: pau b a0104 q0204 k     i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0121.lab vs ../HVITE-M/Rec-mau-M-24/mfG0121.rec
LAB: pau sh       o0103 m i0203 N0303 pau
REC: pau sh o0104 u0204 n i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0122.lab vs ../HVITE-M/Rec-mau-M-24/mfG0122.rec
LAB: pau                 m a0103 s a0203 ts u0303 pau
REC: pau k a0104 w a0204 k a0304 s          u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0123.lab vs ../HVITE-M/Rec-mau-M-24/mfG0123.rec
LAB: pau k a0104 i0204 s e0304 ts u0404 pau
REC: pau k a0104 i0204 k e0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0124.lab vs ../HVITE-M/Rec-mau-M-24/mfG0124.rec
LAB: pau s a0103 g a0203 s u0303 pau
REC: pau h a0103 g a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0125.lab vs ../HVITE-M/Rec-mau-M-24/mfG0125.rec
LAB: pau d e0104 k               a0204 k e0304 r u0404 pau
REC: pau         k u0105 r i0205 a0305 g e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0126.lab vs ../HVITE-M/Rec-mau-M-24/mfG0126.rec
LAB: pau m i0103 n  u0203 k u0303 pau
REC: pau         ch i0102 r u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0127.lab vs ../HVITE-M/Rec-mau-M-24/mfG0127.rec
LAB: pau m u0104 r a0204 s a0304 k i0404 pau
REC: pau t e0104 g a0204 k a0304 r i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0128.lab vs ../HVITE-M/Rec-mau-M-24/mfG0128.rec
LAB: pau   d     a0103 N0203 sh    i0303 pau
REC: pau g a0104 i0204 j     i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0130.lab vs ../HVITE-M/Rec-mau-M-24/mfG0130.rec
LAB: pau   a0104   q0204 s e0304 N0404 pau

REC: pau b a0104 k u0204 z e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0131.lab vs ../HVITE-M/Rec-mau-M-24/mfG0131.rec
LAB: pau s e0104 i0204 ky u0304 u0404 pau
REC: pau k e0104 N0204 ky u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0134.lab vs ../HVITE-M/Rec-mau-M-24/mfG0134.rec
LAB: pau k       i0102 N0202 pau
REC: pau k u0102 g     i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0135.lab vs ../HVITE-M/Rec-mau-M-24/mfG0135.rec
LAB: pau k a0103 t     e0203 i0303 pau
REC: pau k       a0102 b     e0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0136.lab vs ../HVITE-M/Rec-mau-M-24/mfG0136.rec
LAB: pau   ts    u0104 i0204 r a0304 k u0404 pau
REC: pau k u0103 d     a0203 r         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0137.lab vs ../HVITE-M/Rec-mau-M-24/mfG0137.rec
LAB: pau m o0103 g u0203 r u0303 pau
REC: pau k o0103   u0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0138.lab vs ../HVITE-M/Rec-mau-M-24/mfG0138.rec
LAB: pau   r     i0102 k a0202 pau
REC: pau t e0103 i0203 k a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0139.lab vs ../HVITE-M/Rec-mau-M-24/mfG0139.rec
LAB: pau          o0102 k     u0202 pau
REC: pau sh o0103 r     u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0140.lab vs ../HVITE-M/Rec-mau-M-24/mfG0140.rec
LAB: pau s a0103 s a0203 i0303 pau
REC: pau s a0103 k a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0143.lab vs ../HVITE-M/Rec-mau-M-24/mfG0143.rec
LAB: pau   k     i0102 k u0202 pau
REC: pau p o0103 N0203 p u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0144.lab vs ../HVITE-M/Rec-mau-M-24/mfG0144.rec
LAB: pau s a0106 sh i0206 ts u0306 k a0406 e0506 r u0606 pau
REC: pau         m  a0105 ch i0205 g a0305 e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0146.lab vs ../HVITE-M/Rec-mau-M-24/mfG0146.rec
LAB: pau sh i0103 m a0203 u0303 pau
REC: pau n  i0103 n a0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0147.lab vs ../HVITE-M/Rec-mau-M-24/mfG0147.rec
LAB: pau t o0104 r i0204 ts    u0304 g     u0404 pau
REC: pau         b u0104 N0204 m     e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0148.lab vs ../HVITE-M/Rec-mau-M-24/mfG0148.rec
LAB: pau i0102 k                 u0202 pau

REC: pau       k u0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0149.lab vs ../HVITE-M/Rec-mau-M-24/mfG0149.rec
LAB: pau t  o0103 m o0203 n     i0303 pau
REC: pau sh o0103 m       i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0150.lab vs ../HVITE-M/Rec-mau-M-24/mfG0150.rec
LAB: pau t o0103 q0203 s a0303 pau
REC: pau s o0103 u0203 b a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0151.lab vs ../HVITE-M/Rec-mau-M-24/mfG0151.rec
LAB: pau k i0103 k a0203 k u0303 pau
REC: pau n a0103 r a0203 b u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0152.lab vs ../HVITE-M/Rec-mau-M-24/mfG0152.rec
LAB: pau k i0103 m e0203 r u0303 pau
REC: pau n i0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0153.lab vs ../HVITE-M/Rec-mau-M-24/mfG0153.rec
LAB: pau   o0103 o0203 u0303 pau
REC: pau k o0102 m     u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0154.lab vs ../HVITE-M/Rec-mau-M-24/mfG0154.rec
LAB: pau ch i0103 k a0203 u0303 pau
REC: pau ch i0103 g a0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0155.lab vs ../HVITE-M/Rec-mau-M-24/mfG0155.rec
LAB: pau ky o0102   u0202 pau
REC: pau t  o0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0156.lab vs ../HVITE-M/Rec-mau-M-24/mfG0156.rec
LAB: pau         k u0103 d     a0203 s     u0303 pau
REC: pau b a0104 k       u0204 z     e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0157.lab vs ../HVITE-M/Rec-mau-M-24/mfG0157.rec
LAB: pau m o0104 ch    i0204 i0304 r u0404 pau
REC: pau   d     e0103 k     i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0158.lab vs ../HVITE-M/Rec-mau-M-24/mfG0158.rec
LAB: pau m i0103 e0203 r       u0303 pau
REC: pau m e0104 i0204 r o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0159.lab vs ../HVITE-M/Rec-mau-M-24/mfG0159.rec
LAB: pau t o0103 k e0203 r u0303 pau
REC: pau m o0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0160.lab vs ../HVITE-M/Rec-mau-M-24/mfG0160.rec
LAB: pau m o0105 u0205 sh i0305 k o0405 m u0505 pau
REC: pau m                o0103 g u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0161.lab vs ../HVITE-M/Rec-mau-M-24/mfG0161.rec
LAB: pau r o0104 N0204 s o0304 u0404 pau

REC: pau t o0104 u0204 s o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0162.lab vs ../HVITE-M/Rec-mau-M-24/mfG0162.rec
LAB: pau sh u0104 u0204 d a0304 N0404 pau
REC: pau k  e0104 i0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0165.lab vs ../HVITE-M/Rec-mau-M-24/mfG0165.rec
LAB: pau    t     e0104 q0204 p o0304 u0404 pau
REC: pau ts u0104 k     u0204 r o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0166.lab vs ../HVITE-M/Rec-mau-M-24/mfG0166.rec
LAB: pau               i0102 n     u0202 pau
REC: pau b u0104 N0204 r     u0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0167.lab vs ../HVITE-M/Rec-mau-M-24/mfG0167.rec
LAB: pau t o0104 u0204 k e0304 i0404 pau
REC: pau t o0104 u0204 t e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0168.lab vs ../HVITE-M/Rec-mau-M-24/mfG0168.rec
LAB: pau         j i0101 pau
REC: pau k i0102 j i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0169.lab vs ../HVITE-M/Rec-mau-M-24/mfG0169.rec
LAB: pau h e0104 i0204 k     o0304 u0404 pau
REC: pau   g     e0103 N0203 g     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0170.lab vs ../HVITE-M/Rec-mau-M-24/mfG0170.rec
LAB: pau                 f     u0102 i0202 pau
REC: pau k u0104 r i0204 i0304 m     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0172.lab vs ../HVITE-M/Rec-mau-M-24/mfG0172.rec
LAB: pau         s     u0102 u0202 pau
REC: pau k o0104 u0204 u0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0173.lab vs ../HVITE-M/Rec-mau-M-24/mfG0173.rec
LAB: pau j i0104 q0204 s     a0304 i0404 pau
REC: pau k a0103 r     i0203 r     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0174.lab vs ../HVITE-M/Rec-mau-M-24/mfG0174.rec
LAB: pau t e0104 i0204 n e0304 i0404 pau
REC: pau r e0104 N0204 m e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0175.lab vs ../HVITE-M/Rec-mau-M-24/mfG0175.rec
LAB: pau ch u0104 u0204   o0304 u0404 pau
REC: pau y  u0104 u0204 k o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0178.lab vs ../HVITE-M/Rec-mau-M-24/mfG0178.rec
LAB: pau t a0103 i0203 k o0303 pau
REC: pau m a0103 i0203 g o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0180.lab vs ../HVITE-M/Rec-mau-M-24/mfG0180.rec
LAB: pau m o0103 t               o0203 d e0303 pau

REC: pau         t e0104 o0204 k u0304 r e0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0181.lab vs ../HVITE-M/Rec-mau-M-24/mfG0181.rec
LAB: pau s u0103 g a0203 t a0303 pau
REC: pau n a0103 k a0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0182.lab vs ../HVITE-M/Rec-mau-M-24/mfG0182.rec
LAB: pau               y o0102 k o0202 pau
REC: pau h o0104 u0204 k o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0183.lab vs ../HVITE-M/Rec-mau-M-24/mfG0183.rec
LAB: pau t a0104 n     o0204 sh    i0304 i0404 pau
REC: pau   g     a0103 k     u0203 h     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0184.lab vs ../HVITE-M/Rec-mau-M-24/mfG0184.rec
LAB: pau k  a0103 q0203 t e0303 pau
REC: pau ts u0103 i0203 t e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0185.lab vs ../HVITE-M/Rec-mau-M-24/mfG0185.rec
LAB: pau n i0103 q0203 k     i0303 pau
REC: pau k i0103 ts    u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0186.lab vs ../HVITE-M/Rec-mau-M-24/mfG0186.rec
LAB: pau k a0103 i0203 w         a0303 pau
REC: pau   i0104 i0204 w a0304 k e0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0189.lab vs ../HVITE-M/Rec-mau-M-24/mfG0189.rec
LAB: pau k a0105 k e0205 ts u0305 k e0405 r u0505 pau
REC: pau   u0105 k e0205 ts u0305 k e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0190.lab vs ../HVITE-M/Rec-mau-M-24/mfG0190.rec
LAB: pau s a0104 q0204 s u0304 r u0404 pau
REC: pau   k     a0103 s u0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0191.lab vs ../HVITE-M/Rec-mau-M-24/mfG0191.rec
LAB: pau a0103 o0203 g u0303 pau
REC: pau k     o0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0192.lab vs ../HVITE-M/Rec-mau-M-24/mfG0192.rec
LAB: pau         k a0103 r       u0203 i0303 pau
REC: pau k a0104 k u0204 r i0304 ts    u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0193.lab vs ../HVITE-M/Rec-mau-M-24/mfG0193.rec
LAB: pau ts u0104 k a0204 r e0304 r u0404 pau
REC: pau    u0104 k a0204 b e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0194.lab vs ../HVITE-M/Rec-mau-M-24/mfG0194.rec
LAB: pau h i0103 d     a0203 r     i0303 pau
REC: pau b a0104 i0204 t     e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0195.lab vs ../HVITE-M/Rec-mau-M-24/mfG0195.rec
LAB: pau d a0103 N0203 sh i0303 pau

REC: pau g a0103 q0203 k  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0197.lab vs ../HVITE-M/Rec-mau-M-24/mfG0197.rec
LAB: pau   i0104 q0204 sh o0304 u0404 pau
REC: pau t e0104 i0204 ky o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0198.lab vs ../HVITE-M/Rec-mau-M-24/mfG0198.rec
LAB: pau t a0103 t e0203 r u0303 pau
REC: pau k a0103 s e0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0199.lab vs ../HVITE-M/Rec-mau-M-24/mfG0199.rec
LAB: pau k         a0102 s  u0202 pau
REC: pau k u0103 g a0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0200.lab vs ../HVITE-M/Rec-mau-M-24/mfG0200.rec
LAB: pau t  o0104 b i0204 ts u0304 k u0404 pau
REC: pau sh o0104   u0204 d  o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0201.lab vs ../HVITE-M/Rec-mau-M-24/mfG0201.rec
LAB: pau f u0105 r i0205 k a0305 e0405 r u0505 pau
REC: pau t o0105 r i0205 k a0305 e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0203.lab vs ../HVITE-M/Rec-mau-M-24/mfG0203.rec
LAB: pau g  e0104 N0204 d a0304 i0404 pau
REC: pau sh i0104 N0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0204.lab vs ../HVITE-M/Rec-mau-M-24/mfG0204.rec
LAB: pau k       o0103 o0203 s     u0303 pau
REC: pau k o0104 u0204 ts    u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0205.lab vs ../HVITE-M/Rec-mau-M-24/mfG0205.rec
LAB: pau k  u0102 z u0202 pau
REC: pau ts u0102 g u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0206.lab vs ../HVITE-M/Rec-mau-M-24/mfG0206.rec
LAB: pau sh u0104 u0204 s       a0304 i0404 pau
REC: pau    k     u0103 s a0203 r     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0207.lab vs ../HVITE-M/Rec-mau-M-24/mfG0207.rec
LAB: pau                 m e0101 pau
REC: pau n i0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0208.lab vs ../HVITE-M/Rec-mau-M-24/mfG0208.rec
LAB: pau m a0104   N0204 n a0304 k a0404 pau
REC: pau   a0104 z a0204 y a0304 k a0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0209.lab vs ../HVITE-M/Rec-mau-M-24/mfG0209.rec
LAB: pau ts u0103 y o0203 i0303 pau
REC: pau k  i0103 y o0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0211.lab vs ../HVITE-M/Rec-mau-M-24/mfG0211.rec
LAB: pau sh i0103 N0203 g i0303 pau

REC: pau    k     u0102 g i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0212.lab vs ../HVITE-M/Rec-mau-M-24/mfG0212.rec
LAB: pau k e0103 i0203 sh    a0303 pau
REC: pau t e0104 i0204 a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0213.lab vs ../HVITE-M/Rec-mau-M-24/mfG0213.rec
LAB: pau k o0104 N0204 d a0304 N0404 pau
REC: pau b u0104 N0204 t a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0214.lab vs ../HVITE-M/Rec-mau-M-24/mfG0214.rec
LAB: pau       r e0104 N0204 g o0304   u0404 pau
REC: pau i0104 r       e0204 m o0304 n o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0215.lab vs ../HVITE-M/Rec-mau-M-24/mfG0215.rec
LAB: pau ch i0103 k a0203   u0303 pau
REC: pau m  i0103 g a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0216.lab vs ../HVITE-M/Rec-mau-M-24/mfG0216.rec
LAB: pau h o0104 u0204 r i0304 ts u0404 pau
REC: pau k o0104 u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0217.lab vs ../HVITE-M/Rec-mau-M-24/mfG0217.rec
LAB: pau k       a0103 i0203 w     a0303 pau
REC: pau k a0104 i0204 g     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0218.lab vs ../HVITE-M/Rec-mau-M-24/mfG0218.rec
LAB: pau k       u0103 r a0203 i0303 pau
REC: pau k a0104 i0204 h o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0219.lab vs ../HVITE-M/Rec-mau-M-24/mfG0219.rec
LAB: pau j i0102 k u0202 pau
REC: pau m i0102 g i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0220.lab vs ../HVITE-M/Rec-mau-M-24/mfG0220.rec
LAB: pau         k o0104 u0204 ts u0304 u0404 pau
REC: pau r o0103 k       o0203 ts       u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0221.lab vs ../HVITE-M/Rec-mau-M-24/mfG0221.rec
LAB: pau o0103 m         a0203 k e0303 pau
REC: pau u0104 m a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0222.lab vs ../HVITE-M/Rec-mau-M-24/mfG0222.rec
LAB: pau r     o0104 N0204 s     o0304 u0404 pau
REC: pau i0104 m     o0204 u0304 t     o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0223.lab vs ../HVITE-M/Rec-mau-M-24/mfG0223.rec
LAB: pau   i0103   e0203 d e0303 pau
REC: pau r i0103 r e0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0224.lab vs ../HVITE-M/Rec-mau-M-24/mfG0224.rec
LAB: pau         m i0102 r u0202 pau

REC: pau s u0103 b e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0225.lab vs ../HVITE-M/Rec-mau-M-24/mfG0225.rec
LAB: pau s o0104 u0204 z o0304 k u0404 pau
REC: pau   k     o0103 d o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0226.lab vs ../HVITE-M/Rec-mau-M-24/mfG0226.rec
LAB: pau k a0104 N0204 p e0304 k i0404 pau
REC: pau k a0104 N0204 g e0304 k i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0229.lab vs ../HVITE-M/Rec-mau-M-24/mfG0229.rec
LAB: pau h o0103 k e0203 ts u0303 pau
REC: pau h i0103 k e0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0230.lab vs ../HVITE-M/Rec-mau-M-24/mfG0230.rec
LAB: pau f u0104 t o0204 k       o0304 r     o0404 pau
REC: pau                 k o0103 m     o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0231.lab vs ../HVITE-M/Rec-mau-M-24/mfG0231.rec
LAB: pau m o0104 ts u0204 r e0304 r u0404 pau
REC: pau m          o0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0232.lab vs ../HVITE-M/Rec-mau-M-24/mfG0232.rec
LAB: pau k a0103 sh i0203 r a0303 pau
REC: pau r a0103    q0203 p a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0233.lab vs ../HVITE-M/Rec-mau-M-24/mfG0233.rec
LAB: pau k e0104 i0204 z a0304 i0404 pau
REC: pau   g     e0103 h i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0235.lab vs ../HVITE-M/Rec-mau-M-24/mfG0235.rec
LAB: pau k o0104 u0204 k a0304 i0404 pau
REC: pau   g     o0103 k a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0236.lab vs ../HVITE-M/Rec-mau-M-24/mfG0236.rec
LAB: pau t o0104 k u0204 t e0304 i0404 pau
REC: pau g o0104   u0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0237.lab vs ../HVITE-M/Rec-mau-M-24/mfG0237.rec
LAB: pau k a0104 k     e0204 a0304 sh i0404 pau
REC: pau   h     i0103 g     a0203 sh i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0238.lab vs ../HVITE-M/Rec-mau-M-24/mfG0238.rec
LAB: pau j i0103 N0203 sh       u0303 pau
REC: pau m i0104 N0204 sh u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0239.lab vs ../HVITE-M/Rec-mau-M-24/mfG0239.rec
LAB: pau k u0103 k u0203 r u0303 pau
REC: pau         k i0102 r i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0240.lab vs ../HVITE-M/Rec-mau-M-24/mfG0240.rec
LAB: pau y o0104 r o0204 sh    i0304 i0404 pau

REC: pau y         u0103 u0203 k     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0241.lab vs ../HVITE-M/Rec-mau-M-24/mfG0241.rec
LAB: pau         h a0102 k u0202 pau
REC: pau s a0103 b a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0242.lab vs ../HVITE-M/Rec-mau-M-24/mfG0242.rec
LAB: pau g e0104 i0204 j u0304 ts u0404 pau
REC: pau h e0104 i0204 j i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0244.lab vs ../HVITE-M/Rec-mau-M-24/mfG0244.rec
LAB: pau                 sh i0102 o0202 pau
REC: pau r i0104 k u0204 j  o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0245.lab vs ../HVITE-M/Rec-mau-M-24/mfG0245.rec
LAB: pau         s a0102 o0202 pau
REC: pau m i0103 h o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0246.lab vs ../HVITE-M/Rec-mau-M-24/mfG0246.rec
LAB: pau t o0104 k u0204 b e0304 ts u0404 pau
REC: pau         k u0103 b e0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0247.lab vs ../HVITE-M/Rec-mau-M-24/mfG0247.rec
LAB: pau sh u0104 u0204 n e0304 N0404 pau
REC: pau sh i0104 N0204 n e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0249.lab vs ../HVITE-M/Rec-mau-M-24/mfG0249.rec
LAB: pau f u0104 m i0204 k i0304 r i0404 pau
REC: pau         m         i0102 g i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0250.lab vs ../HVITE-M/Rec-mau-M-24/mfG0250.rec
LAB: pau z a0104 N0204 gy o0304 u0404 pau
REC: pau r a0104 N0204 y  o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0252.lab vs ../HVITE-M/Rec-mau-M-24/mfG0252.rec
LAB: pau   e0104 N0204 k a0304   i0404 pau
REC: pau r i0104 N0204 k a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0253.lab vs ../HVITE-M/Rec-mau-M-24/mfG0253.rec
LAB: pau         d e0103 N0203 sh i0303 pau
REC: pau k u0104 r i0204 i0304 m  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0254.lab vs ../HVITE-M/Rec-mau-M-24/mfG0254.rec
LAB: pau s a0104 N0204 my a0304 k u0404 pau
REC: pau   k     a0103 n  a0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0255.lab vs ../HVITE-M/Rec-mau-M-24/mfG0255.rec
LAB: pau k u0103 r e0203 r u0303 pau
REC: pau t a0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0256.lab vs ../HVITE-M/Rec-mau-M-24/mfG0256.rec
LAB: pau s e0104 N0204 s e0304 i0404 pau

REC: pau z e0104 N0204 z e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0259.lab vs ../HVITE-M/Rec-mau-M-24/mfG0259.rec
LAB: pau   sh    o0103 m i0203 N0303 pau
REC: pau k o0104 u0204 n i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0260.lab vs ../HVITE-M/Rec-mau-M-24/mfG0260.rec
LAB: pau a0103 t     e0203 n       a0303 pau
REC: pau s     a0104 i0204 n a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0261.lab vs ../HVITE-M/Rec-mau-M-24/mfG0261.rec
LAB: pau sh         u0103 t a0203 i0303 pau
REC: pau sh u0104 k u0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0262.lab vs ../HVITE-M/Rec-mau-M-24/mfG0262.rec
LAB: pau m     u0103 y o0203 u0303 pau
REC: pau u0104 N0204 d o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0264.lab vs ../HVITE-M/Rec-mau-M-24/mfG0264.rec
LAB: pau                 m a0102 ts u0202 pau
REC: pau k a0103 s u0203 m          u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0265.lab vs ../HVITE-M/Rec-mau-M-24/mfG0265.rec
LAB: pau                 o0102 b i0202 pau
REC: pau k u0104 r i0204 i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0268.lab vs ../HVITE-M/Rec-mau-M-24/mfG0268.rec
LAB: pau t o0103 k         a0203 k u0303 pau
REC: pau         k u0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0270.lab vs ../HVITE-M/Rec-mau-M-24/mfG0270.rec
LAB: pau k o0103 u0203 i0303 pau
REC: pau s o0102 r     u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0271.lab vs ../HVITE-M/Rec-mau-M-24/mfG0271.rec
LAB: pau k i0103 r o0203 k u0303 pau
REC: pau y u0103 s u0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0272.lab vs ../HVITE-M/Rec-mau-M-24/mfG0272.rec
LAB: pau m         a0102 n e0202 pau
REC: pau m a0103 g u0203 r e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0274.lab vs ../HVITE-M/Rec-mau-M-24/mfG0274.rec
LAB: pau          k i0102 k i0202 pau
REC: pau sh i0103 k i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0276.lab vs ../HVITE-M/Rec-mau-M-24/mfG0276.rec
LAB: pau k       o0102 r     o0202 pau
REC: pau k o0103 m     o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0277.lab vs ../HVITE-M/Rec-mau-M-24/mfG0277.rec
LAB: pau sh o0104 u0204 b o0304 u0404 pau

REC: pau k  o0104 u0204 d o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0278.lab vs ../HVITE-M/Rec-mau-M-24/mfG0278.rec
LAB: pau g e0104 N0204 s o0304 u0404 pau
REC: pau r e0104 N0204 s o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0279.lab vs ../HVITE-M/Rec-mau-M-24/mfG0279.rec
LAB: pau z e0104 N0204 ch o0304 u0404 pau
REC: pau b e0104 N0204 sh o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0280.lab vs ../HVITE-M/Rec-mau-M-24/mfG0280.rec
LAB: pau                 s a0102 y a0202 pau
REC: pau k u0103 r a0203 s         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0281.lab vs ../HVITE-M/Rec-mau-M-24/mfG0281.rec
LAB: pau d       a0102 m u0202 pau
REC: pau d a0103 N0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0284.lab vs ../HVITE-M/Rec-mau-M-24/mfG0284.rec
LAB: pau         s a0102 s u0202 pau
REC: pau k a0103 s u0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0285.lab vs ../HVITE-M/Rec-mau-M-24/mfG0285.rec
LAB: pau f u0103 r     u0203   u0303 pau
REC: pau   d     o0103 u0203 m o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0289.lab vs ../HVITE-M/Rec-mau-M-24/mfG0289.rec
LAB: pau k a0103 w a0203 r     i0303 pau
REC: pau k a0104 w a0204 i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0292.lab vs ../HVITE-M/Rec-mau-M-24/mfG0292.rec
LAB: pau o0103 b i0203 r                  u0303 pau
REC: pau       t o0104 r i0204 ts u0304 g u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0293.lab vs ../HVITE-M/Rec-mau-M-24/mfG0293.rec
LAB: pau s a0103 k e0203 b u0303 pau
REC: pau k a0103 k e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0294.lab vs ../HVITE-M/Rec-mau-M-24/mfG0294.rec
LAB: pau   m     u0103 g e0203 N0303 pau
REC: pau k i0104 N0204 b e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0295.lab vs ../HVITE-M/Rec-mau-M-24/mfG0295.rec
LAB: pau k       o0103 r i0203 r u0303 pau
REC: pau k o0104 u0204 g e0304 k i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0296.lab vs ../HVITE-M/Rec-mau-M-24/mfG0296.rec
LAB: pau   j     o0103 ch u0203 u0303 pau
REC: pau k u0104 u0204 ch u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0298.lab vs ../HVITE-M/Rec-mau-M-24/mfG0298.rec
LAB: pau h a0103 sh i0203 r u0303 pau

REC: pau b a0103 k  e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0299.lab vs ../HVITE-M/Rec-mau-M-24/mfG0299.rec
LAB: pau s u0106 k u0206 n a0306 k  u0406 t o0506 m o0606 pau
REC: pau         k         u0104 ts u0204 r o0304 g u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0300.lab vs ../HVITE-M/Rec-mau-M-24/mfG0300.rec
LAB: pau y o0104 u0204 s e0304 k i0404 pau
REC: pau y       u0103 s u0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0301.lab vs ../HVITE-M/Rec-mau-M-24/mfG0301.rec
LAB: pau k     u0103 y     a0203 m     u0303 pau
REC: pau u0104 r     a0204 n     a0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0303.lab vs ../HVITE-M/Rec-mau-M-24/mfG0303.rec
LAB: pau   t     o0102 sh    i0202 pau
REC: pau b o0103 k     i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0304.lab vs ../HVITE-M/Rec-mau-M-24/mfG0304.rec
LAB: pau                 k a0102 s u0202 pau
REC: pau h i0103 r a0203 k         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0305.lab vs ../HVITE-M/Rec-mau-M-24/mfG0305.rec
LAB: pau n i0103   q0203 k i0303 pau
REC: pau m i0103 b u0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0307.lab vs ../HVITE-M/Rec-mau-M-24/mfG0307.rec
LAB: pau t e0103 g o0203 r o0303 pau
REC: pau s e0103 r o0203   N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0308.lab vs ../HVITE-M/Rec-mau-M-24/mfG0308.rec
LAB: pau         j i0103 sh u0203 k u0303 pau
REC: pau k u0103 j          i0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0310.lab vs ../HVITE-M/Rec-mau-M-24/mfG0310.rec
LAB: pau   y     o0103 k e0203 i0303 pau
REC: pau h o0104 u0204 g e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0311.lab vs ../HVITE-M/Rec-mau-M-24/mfG0311.rec
LAB: pau k o0103   o0203 r u0303 pau
REC: pau   o0103 m o0203   u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0312.lab vs ../HVITE-M/Rec-mau-M-24/mfG0312.rec
LAB: pau n u0103 k         e0203 r u0303 pau
REC: pau         k o0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0313.lab vs ../HVITE-M/Rec-mau-M-24/mfG0313.rec
LAB: pau h  a0104 N0204 t     a0304 i0404 pau
REC: pau ts u0104 r     i0204 a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0314.lab vs ../HVITE-M/Rec-mau-M-24/mfG0314.rec
LAB: pau       s a0103 s u0203 g a0303 pau

REC: pau i0104 t a0204 z u0304 r a0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0317.lab vs ../HVITE-M/Rec-mau-M-24/mfG0317.rec
LAB: pau   sh    o0103 m i0203 N0303 pau
REC: pau h o0104 N0204 n i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0318.lab vs ../HVITE-M/Rec-mau-M-24/mfG0318.rec
LAB: pau r o0103 N0203 r     i0303 pau
REC: pau d o0103       u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0319.lab vs ../HVITE-M/Rec-mau-M-24/mfG0319.rec
LAB: pau d o0103 u0203 m         o0303 pau
REC: pau   k     i0103 m o0203 n o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0320.lab vs ../HVITE-M/Rec-mau-M-24/mfG0320.rec
LAB: pau         ts u0102 r u0202 pau
REC: pau m o0103 g  u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0321.lab vs ../HVITE-M/Rec-mau-M-24/mfG0321.rec
LAB: pau   o0102 j i0202 pau
REC: pau m u0102 k i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0322.lab vs ../HVITE-M/Rec-mau-M-24/mfG0322.rec
LAB: pau b u0103 j o0203 k u0303 pau
REC: pau t o0103 j i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0324.lab vs ../HVITE-M/Rec-mau-M-24/mfG0324.rec
LAB: pau sh o0104 u0204 d o0304 k u0404 pau
REC: pau    k     o0103 d o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0325.lab vs ../HVITE-M/Rec-mau-M-24/mfG0325.rec
LAB: pau n         o0103 z o0203 m i0303 pau
REC: pau n o0104 k o0204 g i0304 r i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0326.lab vs ../HVITE-M/Rec-mau-M-24/mfG0326.rec
LAB: pau m e0103 d a0203 ts u0303 pau
REC: pau k u0103 g a0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0327.lab vs ../HVITE-M/Rec-mau-M-24/mfG0327.rec
LAB: pau         sh    i0103 a0203 i0303 pau
REC: pau k o0104 u0204 g     a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0329.lab vs ../HVITE-M/Rec-mau-M-24/mfG0329.rec
LAB: pau h a0104 ts u0204 r a0304 ts u0404 pau
REC: pau h a0104 k  i0204 d a0304 s  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0330.lab vs ../HVITE-M/Rec-mau-M-24/mfG0330.rec
LAB: pau k o0103 N0203 d       o0303 pau
REC: pau d o0104 N0204 d o0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0331.lab vs ../HVITE-M/Rec-mau-M-24/mfG0331.rec
LAB: pau t a0104 ch i0204 m     a0304 ch i0404 pau

REC: pau k a0104    k     e0204 a0304 sh i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0334.lab vs ../HVITE-M/Rec-mau-M-24/mfG0334.rec
LAB: pau         y o0102 s o0202 pau
REC: pau h o0103 k o0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0335.lab vs ../HVITE-M/Rec-mau-M-24/mfG0335.rec
LAB: pau j u0104 u0204 t a0304 k u0404 pau
REC: pau j u0104 N0204 k a0304   N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0336.lab vs ../HVITE-M/Rec-mau-M-24/mfG0336.rec
LAB: pau         b a0102 k a0202 pau
REC: pau m a0103 t a0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0337.lab vs ../HVITE-M/Rec-mau-M-24/mfG0337.rec
LAB: pau   y     a0103 s     u0203 r i0303 pau
REC: pau h a0104 ts    u0204 i0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0338.lab vs ../HVITE-M/Rec-mau-M-24/mfG0338.rec
LAB: pau                    u0102 s o0202 pau
REC: pau k u0104 ts u0204 r o0304 g u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0340.lab vs ../HVITE-M/Rec-mau-M-24/mfG0340.rec
LAB: pau k a0104       k u0204 sh i0304 N0404 pau
REC: pau d a0104 i0204 k          i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0341.lab vs ../HVITE-M/Rec-mau-M-24/mfG0341.rec
LAB: pau               s o0104 r o0204 e0304 r u0404 pau
REC: pau u0105 n u0205 b o0305 r       e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0342.lab vs ../HVITE-M/Rec-mau-M-24/mfG0342.rec
LAB: pau sh i0103 s     o0203 u0303 pau
REC: pau sh       i0102 r     o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0344.lab vs ../HVITE-M/Rec-mau-M-24/mfG0344.rec
LAB: pau j u0104 N0204 t o0304 u0404 pau
REC: pau k e0104 N0204 t o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0345.lab vs ../HVITE-M/Rec-mau-M-24/mfG0345.rec
LAB: pau sh i0104 N0204 d o0304 u0404 pau
REC: pau    k     i0103 b o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0346.lab vs ../HVITE-M/Rec-mau-M-24/mfG0346.rec
LAB: pau g i0103 N0203 m i0303 pau
REC: pau t a0103 N0203   i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0348.lab vs ../HVITE-M/Rec-mau-M-24/mfG0348.rec
LAB: pau a0103 m     a0203 r       i0303 pau
REC: pau g     a0104 N0204 r a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0349.lab vs ../HVITE-M/Rec-mau-M-24/mfG0349.rec
LAB: pau h i0103 s a0203 sh i0303 pau

REC: pau         k a0102 j  i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0351.lab vs ../HVITE-M/Rec-mau-M-24/mfG0351.rec
LAB: pau s e0104 i0204 k a0304 k u0404 pau
REC: pau s e0104 i0204 y a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0352.lab vs ../HVITE-M/Rec-mau-M-24/mfG0352.rec
LAB: pau           o0103 t  o0203 s u0303 pau
REC: pau b o0104 k u0204 ch i0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0354.lab vs ../HVITE-M/Rec-mau-M-24/mfG0354.rec
LAB: pau t a0104 y a0204 s       u0304 i0404 pau
REC: pau         y u0103 s u0203 g     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0355.lab vs ../HVITE-M/Rec-mau-M-24/mfG0355.rec
LAB: pau m o0103 r a0203 u0303 pau
REC: pau h a0103 r a0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0356.lab vs ../HVITE-M/Rec-mau-M-24/mfG0356.rec
LAB: pau k o0103 u0203 sh i0303 pau
REC: pau   h     o0102 sh i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0357.lab vs ../HVITE-M/Rec-mau-M-24/mfG0357.rec
LAB: pau t o0104 k u0204 t e0304 i0404 pau
REC: pau b o0104   u0204 k e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0358.lab vs ../HVITE-M/Rec-mau-M-24/mfG0358.rec
LAB: pau j o0104 u0204 o0304 u0404 pau
REC: pau   k     o0102 r     o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0360.lab vs ../HVITE-M/Rec-mau-M-24/mfG0360.rec
LAB: pau s e0103 b i0203 r         o0303 pau
REC: pau k e0103         r e0203 d o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0363.lab vs ../HVITE-M/Rec-mau-M-24/mfG0363.rec
LAB: pau   t     o0102 m u0202 pau
REC: pau s o0103 u0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0364.lab vs ../HVITE-M/Rec-mau-M-24/mfG0364.rec
LAB: pau   n     e0103 m u0203 r u0303 pau
REC: pau t e0104 e0204 b u0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0365.lab vs ../HVITE-M/Rec-mau-M-24/mfG0365.rec
LAB: pau j u0104 u0204 by    o0304 u0404 pau
REC: pau         e0103 N0203 ry    o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0366.lab vs ../HVITE-M/Rec-mau-M-24/mfG0366.rec
LAB: pau k o0103 u0203 ch a0303 pau
REC: pau r o0103 u0203 k  a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0368.lab vs ../HVITE-M/Rec-mau-M-24/mfG0368.rec
LAB: pau           sh    o0103 b     u0203 N0303 pau

REC: pau k u0104 z u0204 r     e0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0369.lab vs ../HVITE-M/Rec-mau-M-24/mfG0369.rec
LAB: pau y                 o0101 pau
REC: pau y u0103 r u0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0371.lab vs ../HVITE-M/Rec-mau-M-24/mfG0371.rec
LAB: pau sh i0103 m e0203   i0303 pau
REC: pau sh i0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0372.lab vs ../HVITE-M/Rec-mau-M-24/mfG0372.rec
LAB: pau m i0104 y a0204 b u0304 r         u0404 pau
REC: pau                         r a0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0373.lab vs ../HVITE-M/Rec-mau-M-24/mfG0373.rec
LAB: pau           s     e0102 i0202 pau
REC: pau h a0103 g e0203 m     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0374.lab vs ../HVITE-M/Rec-mau-M-24/mfG0374.rec
LAB: pau           p     e0102 N0202 pau
REC: pau t e0103 g a0203 r     a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0377.lab vs ../HVITE-M/Rec-mau-M-24/mfG0377.rec
LAB: pau   o0102   u0202 pau
REC: pau t o0102 m o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0378.lab vs ../HVITE-M/Rec-mau-M-24/mfG0378.rec
LAB: pau         k     i0103 n     o0203 u0303 pau
REC: pau t a0104 N0204 d     o0304 k     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0379.lab vs ../HVITE-M/Rec-mau-M-24/mfG0379.rec
LAB: pau j o0104 u0204 o0304 u0404 pau
REC: pau m u0103 k     o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0380.lab vs ../HVITE-M/Rec-mau-M-24/mfG0380.rec
LAB: pau sh i0102 n         a0202 pau
REC: pau    i0103 n a0203 k a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0381.lab vs ../HVITE-M/Rec-mau-M-24/mfG0381.rec
LAB: pau                 d a0103 k a0203 r a0303 pau
REC: pau k u0105 w a0205 d a0305 t e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0382.lab vs ../HVITE-M/Rec-mau-M-24/mfG0382.rec
LAB: pau b u0103 ts u0203 r i0303 pau
REC: pau s u0103 s  u0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0383.lab vs ../HVITE-M/Rec-mau-M-24/mfG0383.rec
LAB: pau   e0103 N0203 ry       o0303 pau
REC: pau s e0104 N0204 ry o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0384.lab vs ../HVITE-M/Rec-mau-M-24/mfG0384.rec
LAB: pau k e0104 i0204 s o0304 ts u0404 pau

REC: pau k       e0103 z u0203 r  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0385.lab vs ../HVITE-M/Rec-mau-M-24/mfG0385.rec
LAB: pau y a0103 sh i0203 N0303 pau
REC: pau d a0103 sh i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0386.lab vs ../HVITE-M/Rec-mau-M-24/mfG0386.rec
LAB: pau n a0104 g a0204 m e0304 r u0404 pau
REC: pau n a0104 g a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0388.lab vs ../HVITE-M/Rec-mau-M-24/mfG0388.rec
LAB: pau t o0102 k               u0202 pau
REC: pau k o0104 k u0204 d e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0389.lab vs ../HVITE-M/Rec-mau-M-24/mfG0389.rec
LAB: pau   e0103 N0203 ry         o0303 pau
REC: pau k e0104 N0204 ry o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0391.lab vs ../HVITE-M/Rec-mau-M-24/mfG0391.rec
LAB: pau f u0103 z     o0203 k       u0303 pau
REC: pau   k     o0104 u0204 k e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0392.lab vs ../HVITE-M/Rec-mau-M-24/mfG0392.rec
LAB: pau k u0103 sh    i0203 N0303 pau
REC: pau g i0103 N0203 m     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0393.lab vs ../HVITE-M/Rec-mau-M-24/mfG0393.rec
LAB: pau k u0102 r         o0202 pau
REC: pau k o0103 r o0203 b u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0394.lab vs ../HVITE-M/Rec-mau-M-24/mfG0394.rec
LAB: pau k a0104 k u0204 b e0304 ts u0404 pau
REC: pau r a0104   i0204 g e0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0395.lab vs ../HVITE-M/Rec-mau-M-24/mfG0395.rec
LAB: pau t a0104 N0204 j o0304 u0404 pau
REC: pau d a0104 i0204 k o0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0396.lab vs ../HVITE-M/Rec-mau-M-24/mfG0396.rec
LAB: pau ts u0104 u0204 k             o0304 u0404 pau
REC: pau                k o0103 u0203 h     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0397.lab vs ../HVITE-M/Rec-mau-M-24/mfG0397.rec
LAB: pau y o0104 u0204 ky u0304 u0404 pau
REC: pau t o0104 q0204 ky u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0398.lab vs ../HVITE-M/Rec-mau-M-24/mfG0398.rec
LAB: pau         m o0103 d e0203 r u0303 pau
REC: pau k u0104 z u0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0399.lab vs ../HVITE-M/Rec-mau-M-24/mfG0399.rec
LAB: pau d  e0104 sh a0204 b a0304 r u0404 pau

REC: pau ts u0104 k  a0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0400.lab vs ../HVITE-M/Rec-mau-M-24/mfG0400.rec
LAB: pau k             e0103 N0203 ky o0303 pau
REC: pau k e0104 N0204 ry    o0304 k  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0401.lab vs ../HVITE-M/Rec-mau-M-24/mfG0401.rec
LAB: pau sh o0103 t a0203   i0303 pau
REC: pau k  u0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0402.lab vs ../HVITE-M/Rec-mau-M-24/mfG0402.rec
LAB: pau d o0103 u0203 i0303 pau
REC: pau m o0102 r     i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0404.lab vs ../HVITE-M/Rec-mau-M-24/mfG0404.rec
LAB: pau t a0104 z u0204 n e0304 r u0404 pau
REC: pau s a0104 k u0204 n e0304   N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0406.lab vs ../HVITE-M/Rec-mau-M-24/mfG0406.rec
LAB: pau   sh    i0103 a0203   N0303 pau
REC: pau k u0103 y     a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0408.lab vs ../HVITE-M/Rec-mau-M-24/mfG0408.rec
LAB: pau         b a0102 n     e0202 pau
REC: pau k i0103 b       a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0409.lab vs ../HVITE-M/Rec-mau-M-24/mfG0409.rec
LAB: pau sh i0103 g o0203 t o0303 pau
REC: pau k  i0103 m o0203 n o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0410.lab vs ../HVITE-M/Rec-mau-M-24/mfG0410.rec
LAB: pau          r     i0103 y     u0203 u0303 pau
REC: pau ry u0104 u0204 g     a0304 k     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0411.lab vs ../HVITE-M/Rec-mau-M-24/mfG0411.rec
LAB: pau sh         i0102 t         a0202 pau
REC: pau sh i0104 t a0204 t e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0412.lab vs ../HVITE-M/Rec-mau-M-24/mfG0412.rec
LAB: pau o0105 ch i0205 b u0305 r e0405 r u0505 pau
REC: pau                k i0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0413.lab vs ../HVITE-M/Rec-mau-M-24/mfG0413.rec
LAB: pau k e0104       i0204 h     i0304 N0404 pau
REC: pau n e0104 N0204 g     e0304 ts    u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0414.lab vs ../HVITE-M/Rec-mau-M-24/mfG0414.rec
LAB: pau         o0103 u0203 b     o0303 pau
REC: pau d o0104 N0204 d     o0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0416.lab vs ../HVITE-M/Rec-mau-M-24/mfG0416.rec
LAB: pau   m     i0103 r a0203 i0303 pau

REC: pau g a0104 i0204 j i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0417.lab vs ../HVITE-M/Rec-mau-M-24/mfG0417.rec
LAB: pau k i0105 r i0205 a0305 g e0405 r u0505 pau
REC: pau h i0105 k i0205 a0305 g e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0418.lab vs ../HVITE-M/Rec-mau-M-24/mfG0418.rec
LAB: pau         s u0103 s u0203 g u0303 pau
REC: pau t e0104 k i0204 s u0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0419.lab vs ../HVITE-M/Rec-mau-M-24/mfG0419.rec
LAB: pau n u0103 r a0203 s u0303 pau
REC: pau m o0103 r a0203   u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0420.lab vs ../HVITE-M/Rec-mau-M-24/mfG0420.rec
LAB: pau    y     o0103 s a0203 N0303 pau
REC: pau ch u0104 u0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0421.lab vs ../HVITE-M/Rec-mau-M-24/mfG0421.rec
LAB: pau h e0104 i0204 k i0304 N0404 pau
REC: pau t e0104 q0204 k i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0422.lab vs ../HVITE-M/Rec-mau-M-24/mfG0422.rec
LAB: pau sh       i0103 k i0203 r u0303 pau
REC: pau sh i0104 i0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0423.lab vs ../HVITE-M/Rec-mau-M-24/mfG0423.rec
LAB: pau o0105 s o0205 r     o0305 sh i0405 i0505 pau
REC: pau         k     o0103 u0203 sh       i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0424.lab vs ../HVITE-M/Rec-mau-M-24/mfG0424.rec
LAB: pau k o0103 t     e0203 N0303 pau
REC: pau k o0104 u0204 e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0427.lab vs ../HVITE-M/Rec-mau-M-24/mfG0427.rec
LAB: pau k a0103 g e0203   N0303 pau
REC: pau m a0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0428.lab vs ../HVITE-M/Rec-mau-M-24/mfG0428.rec
LAB: pau         b u0103 j  o0203 k u0303 pau
REC: pau g a0104 k u0204 ry o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0429.lab vs ../HVITE-M/Rec-mau-M-24/mfG0429.rec
LAB: pau k o0103 n o0203 m         u0303 pau
REC: pau   o0104 m o0204 m u0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0430.lab vs ../HVITE-M/Rec-mau-M-24/mfG0430.rec
LAB: pau          t o0102 o0202 pau
REC: pau ch o0103 k i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0431.lab vs ../HVITE-M/Rec-mau-M-24/mfG0431.rec
LAB: pau k o0104 u0204 k     o0304 k u0404 pau

REC: pau k o0104 g     o0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0433.lab vs ../HVITE-M/Rec-mau-M-24/mfG0433.rec
LAB: pau m a0102 g         o0202 pau
REC: pau t a0103 g u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0434.lab vs ../HVITE-M/Rec-mau-M-24/mfG0434.rec
LAB: pau h i0103   e0203 r u0303 pau
REC: pau   i0103 j i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0436.lab vs ../HVITE-M/Rec-mau-M-24/mfG0436.rec
LAB: pau y         a0102 d o0202 pau
REC: pau y u0103 r u0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0437.lab vs ../HVITE-M/Rec-mau-M-24/mfG0437.rec
LAB: pau n a0103 y a0203 m u0303 pau
REC: pau n a0103 r a0203 b u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0438.lab vs ../HVITE-M/Rec-mau-M-24/mfG0438.rec
LAB: pau s a0102 y a0202 pau
REC: pau g a0102 r a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0439.lab vs ../HVITE-M/Rec-mau-M-24/mfG0439.rec
LAB: pau m  i0103 r a0203 i0303 pau
REC: pau sh i0103   a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0440.lab vs ../HVITE-M/Rec-mau-M-24/mfG0440.rec
LAB: pau f u0103 r     u0203 u0303 pau
REC: pau   n     u0102 g     u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0441.lab vs ../HVITE-M/Rec-mau-M-24/mfG0441.rec
LAB: pau k i0104 y o0204 m e0304 r u0404 pau
REC: pau h i0104 r o0204 m e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0442.lab vs ../HVITE-M/Rec-mau-M-24/mfG0442.rec
LAB: pau         a0104 t a0204 e0304 r u0404 pau
REC: pau k a0105 N0205 g a0305 e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0443.lab vs ../HVITE-M/Rec-mau-M-24/mfG0443.rec
LAB: pau s u0103 s     u0203 g     u0303 pau
REC: pau         i0103 z     e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0444.lab vs ../HVITE-M/Rec-mau-M-24/mfG0444.rec
LAB: pau b u0102 k                     i0202 pau
REC: pau         k o0104 o0204 h i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0446.lab vs ../HVITE-M/Rec-mau-M-24/mfG0446.rec
LAB: pau ts u0104 k         i0204 a0304 u0404 pau
REC: pau          k u0103 y a0203 m     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0448.lab vs ../HVITE-M/Rec-mau-M-24/mfG0448.rec
LAB: pau   a0103 t e0203 n a0303 pau

REC: pau k u0103 r u0203 m a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0449.lab vs ../HVITE-M/Rec-mau-M-24/mfG0449.rec
LAB: pau m i0104 n     o0204 g a0304 s u0404 pau
REC: pau   k     o0104 u0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0450.lab vs ../HVITE-M/Rec-mau-M-24/mfG0450.rec
LAB: pau s o0104 r o0204 b a0304 N0404 pau
REC: pau k o0104   N0204 n a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0451.lab vs ../HVITE-M/Rec-mau-M-24/mfG0451.rec
LAB: pau k a0104    i0204 sh a0304 k u0404 pau
REC: pau m a0104 ch i0204 g  a0304   i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0452.lab vs ../HVITE-M/Rec-mau-M-24/mfG0452.rec
LAB: pau       k         e0101 pau
REC: pau u0103 k e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0453.lab vs ../HVITE-M/Rec-mau-M-24/mfG0453.rec
LAB: pau         h e0102 y a0202 pau
REC: pau k o0103 g a0203 t a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0454.lab vs ../HVITE-M/Rec-mau-M-24/mfG0454.rec
LAB: pau ts u0104 b u0204 y a0304 k u0404 pau
REC: pau          k u0103 y a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0456.lab vs ../HVITE-M/Rec-mau-M-24/mfG0456.rec
LAB: pau a0103 r a0203 t         a0303 pau
REC: pau               t a0102 m a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0457.lab vs ../HVITE-M/Rec-mau-M-24/mfG0457.rec
LAB: pau o0104 t o0204 t o0304 sh i0404 pau
REC: pau       k o0103 d o0203 k  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0460.lab vs ../HVITE-M/Rec-mau-M-24/mfG0460.rec
LAB: pau                         n a0101 pau
REC: pau f u0104 k u0204 r a0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0461.lab vs ../HVITE-M/Rec-mau-M-24/mfG0461.rec
LAB: pau         i0103 s a0203 N0303 pau
REC: pau k a0104 i0204 g a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0463.lab vs ../HVITE-M/Rec-mau-M-24/mfG0463.rec
LAB: pau   sh    u0103 u0203 r i0303 pau
REC: pau n i0103 k     u0203   i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0464.lab vs ../HVITE-M/Rec-mau-M-24/mfG0464.rec
LAB: pau          j o0103 u0203 h     o0303 pau
REC: pau ch i0103 j             o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0465.lab vs ../HVITE-M/Rec-mau-M-24/mfG0465.rec
LAB: pau           i0104 q0204 t e0304 i0404 pau

REC: pau k o0104 g i0204 q0304 t       e0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0466.lab vs ../HVITE-M/Rec-mau-M-24/mfG0466.rec
LAB: pau sh i0104 N0204 g a0304 i0404 pau
REC: pau sh       i0103 g a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0467.lab vs ../HVITE-M/Rec-mau-M-24/mfG0467.rec
LAB: pau sh i0103   N0203 r i0303 pau
REC: pau j  i0103 k i0203 n i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0469.lab vs ../HVITE-M/Rec-mau-M-24/mfG0469.rec
LAB: pau a0104 q0204 p a0304 k u0404 pau
REC: pau k     a0103 g a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0471.lab vs ../HVITE-M/Rec-mau-M-24/mfG0471.rec
LAB: pau j u0104 N0204 k a0304 N0404 pau
REC: pau h i0104 N0204 p a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0473.lab vs ../HVITE-M/Rec-mau-M-24/mfG0473.rec
LAB: pau o0104 N0204 s e0304 N0404 pau
REC: pau g     o0103 z e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0474.lab vs ../HVITE-M/Rec-mau-M-24/mfG0474.rec
LAB: pau s u0103   e0203 r u0303 pau
REC: pau z u0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0475.lab vs ../HVITE-M/Rec-mau-M-24/mfG0475.rec
LAB: pau         b u0103 N0203 p u0303 pau
REC: pau k a0103 b       u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0476.lab vs ../HVITE-M/Rec-mau-M-24/mfG0476.rec
LAB: pau h a0103 sh i0203 g o0303 pau
REC: pau b a0103 k  u0203 r o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0478.lab vs ../HVITE-M/Rec-mau-M-24/mfG0478.rec
LAB: pau m e0103 d a0203 ts u0303 pau
REC: pau k a0103 g a0203 k  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0480.lab vs ../HVITE-M/Rec-mau-M-24/mfG0480.rec
LAB: pau         t a0102 b i0202 pau
REC: pau k a0103 g i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0481.lab vs ../HVITE-M/Rec-mau-M-24/mfG0481.rec
LAB: pau       r e0103 j a0203   a0303 pau
REC: pau a0103 r         a0203 t a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0483.lab vs ../HVITE-M/Rec-mau-M-24/mfG0483.rec
LAB: pau k i0104 k o0204 e0304 r u0404 pau
REC: pau k o0104 g o0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0484.lab vs ../HVITE-M/Rec-mau-M-24/mfG0484.rec
LAB: pau          d o0103 n a0203 r u0303 pau

REC: pau sh i0104 z u0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0486.lab vs ../HVITE-M/Rec-mau-M-24/mfG0486.rec
LAB: pau          n i0103 a0203 u0303 pau
REC: pau ts u0104 k i0204 a0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0487.lab vs ../HVITE-M/Rec-mau-M-24/mfG0487.rec
LAB: pau h a0104       r i0204 a0304 i0404 pau
REC: pau g a0104 N0204 r       a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0488.lab vs ../HVITE-M/Rec-mau-M-24/mfG0488.rec
LAB: pau   y     o0103 k e0203 i0303 pau
REC: pau g o0104 u0204 k e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0489.lab vs ../HVITE-M/Rec-mau-M-24/mfG0489.rec
LAB: pau s e0104 i0204 j i0304 N0404 pau
REC: pau   g     e0103 h i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0490.lab vs ../HVITE-M/Rec-mau-M-24/mfG0490.rec
LAB: pau   y     u0103 k  i0203 k i0303 pau
REC: pau m u0104 i0204 sh i0304 k i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0491.lab vs ../HVITE-M/Rec-mau-M-24/mfG0491.rec
LAB: pau s e0103 m a0203 r u0303 pau
REC: pau k u0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0492.lab vs ../HVITE-M/Rec-mau-M-24/mfG0492.rec
LAB: pau         t       o0102 sh o0202 pau
REC: pau s u0104 t o0204 o0304 b  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0493.lab vs ../HVITE-M/Rec-mau-M-24/mfG0493.rec
LAB: pau s  o0103 k       u0203 d     o0303 pau
REC: pau ts u0104 k u0204 r     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0494.lab vs ../HVITE-M/Rec-mau-M-24/mfG0494.rec
LAB: pau                 k u0101 pau
REC: pau m a0103 r u0203 d e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0495.lab vs ../HVITE-M/Rec-mau-M-24/mfG0495.rec
LAB: pau k       i0103 k i0203 N0303 pau
REC: pau k i0104 N0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0496.lab vs ../HVITE-M/Rec-mau-M-24/mfG0496.rec
LAB: pau ky o0104 u0204 y o0304 u0404 pau
REC: pau d  o0104 u0204 y o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0497.lab vs ../HVITE-M/Rec-mau-M-24/mfG0497.rec
LAB: pau my o0103 u0203 j       i0303 pau
REC: pau y  o0104 u0204 j i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0499.lab vs ../HVITE-M/Rec-mau-M-24/mfG0499.rec
LAB: pau n i0103 g a0203 s  u0303 pau

REC: pau g o0103 g a0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0500.lab vs ../HVITE-M/Rec-mau-M-24/mfG0500.rec
LAB: pau sh i0103 b     o0203 u0303 pau
REC: pau k  o0103 u0203 g     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0502.lab vs ../HVITE-M/Rec-mau-M-24/mfG0502.rec
LAB: pau k  u0103 r o0203   u0303 pau
REC: pau ts u0103 m o0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0503.lab vs ../HVITE-M/Rec-mau-M-24/mfG0503.rec
LAB: pau i0104 t         a0204 m e0304 r u0404 pau
REC: pau       t o0104 g a0204 m e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0506.lab vs ../HVITE-M/Rec-mau-M-24/mfG0506.rec
LAB: pau z a0104 ts u0204 y  o0304   u0404 pau
REC: pau g a0104 k  u0204 ry o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0507.lab vs ../HVITE-M/Rec-mau-M-24/mfG0507.rec
LAB: pau       d o0103 k a0203 s u0303 pau
REC: pau o0104 d o0204 k a0304 s u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0508.lab vs ../HVITE-M/Rec-mau-M-24/mfG0508.rec
LAB: pau i0104 t         a0204 w a0304 r     u0404 pau
REC: pau       t a0104 m a0204   sh    i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0510.lab vs ../HVITE-M/Rec-mau-M-24/mfG0510.rec
LAB: pau h o0103 sh o0203   u0303 pau
REC: pau m o0103 t  o0203 d e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0511.lab vs ../HVITE-M/Rec-mau-M-24/mfG0511.rec
LAB: pau   i0104 q0204 t a0304 N0404 pau
REC: pau b i0104 N0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0512.lab vs ../HVITE-M/Rec-mau-M-24/mfG0512.rec
LAB: pau j i0102 k       u0202 pau
REC: pau h i0103 k u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0513.lab vs ../HVITE-M/Rec-mau-M-24/mfG0513.rec
LAB: pau h a0104       g e0204 sh    i0304 i0404 pau
REC: pau k a0104 N0204 g       e0304 k     i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0515.lab vs ../HVITE-M/Rec-mau-M-24/mfG0515.rec
LAB: pau    k     e0103 z u0203 r     u0303 pau
REC: pau sh i0104 i0204 z       u0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0516.lab vs ../HVITE-M/Rec-mau-M-24/mfG0516.rec
LAB: pau   e0103 N0203 ry         o0303 pau
REC: pau k e0104 N0204 ry o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0518.lab vs ../HVITE-M/Rec-mau-M-24/mfG0518.rec
LAB: pau sh u0104 q0204 p a0304 ts u0404 pau

REC: pau h  a0104 N0204 p a0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0519.lab vs ../HVITE-M/Rec-mau-M-24/mfG0519.rec
LAB: pau sh    i0102 r     o0202 pau
REC: pau u0103 d     o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0520.lab vs ../HVITE-M/Rec-mau-M-24/mfG0520.rec
LAB: pau s a0103 s a0203   i0303 pau
REC: pau k a0103 g a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0523.lab vs ../HVITE-M/Rec-mau-M-24/mfG0523.rec
LAB: pau m o0104 N0204 d a0304 i0404 pau
REC: pau g u0104 N0204 t a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0527.lab vs ../HVITE-M/Rec-mau-M-24/mfG0527.rec
LAB: pau f u0104 u0204 z o0304 k u0404 pau
REC: pau   k     a0103 b u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0529.lab vs ../HVITE-M/Rec-mau-M-24/mfG0529.rec
LAB: pau sh i0103 sh u0203 ts u0303 pau
REC: pau k  u0103 ts u0203    u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0530.lab vs ../HVITE-M/Rec-mau-M-24/mfG0530.rec
LAB: pau         k a0103 e0203 r u0303 pau
REC: pau m u0104 k a0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0531.lab vs ../HVITE-M/Rec-mau-M-24/mfG0531.rec
LAB: pau   t     a0103 y o0203 r     u0303 pau
REC: pau d a0104 i0204 y       o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0533.lab vs ../HVITE-M/Rec-mau-M-24/mfG0533.rec
LAB: pau         sh    u0102 t     o0202 pau
REC: pau k o0104 u0204 ts    u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0534.lab vs ../HVITE-M/Rec-mau-M-24/mfG0534.rec
LAB: pau ky o0104 u0204 d a0304 i0404 pau
REC: pau t  o0104 u0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0537.lab vs ../HVITE-M/Rec-mau-M-24/mfG0537.rec
LAB: pau h i0103 k a0203 N0303 pau
REC: pau k i0103 b a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0540.lab vs ../HVITE-M/Rec-mau-M-24/mfG0540.rec
LAB: pau sh u0103 t a0203 i0303 pau
REC: pau k  u0103 r a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0541.lab vs ../HVITE-M/Rec-mau-M-24/mfG0541.rec
LAB: pau   sh    i0103 ts    u0203 d     o0303 pau
REC: pau t a0104 k     u0204 sh    i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0542.lab vs ../HVITE-M/Rec-mau-M-24/mfG0542.rec
LAB: pau               k a0102 b  i0202 pau

REC: pau r a0104 i0204 g e0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0543.lab vs ../HVITE-M/Rec-mau-M-24/mfG0543.rec
LAB: pau m u0103 sh i0203 r o0303 pau
REC: pau m i0103 b  u0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0544.lab vs ../HVITE-M/Rec-mau-M-24/mfG0544.rec
LAB: pau n i0104 N0204 t a0304 i0404 pau
REC: pau g u0104 N0204 t a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0546.lab vs ../HVITE-M/Rec-mau-M-24/mfG0546.rec
LAB: pau k a0102 g     e0202 pau
REC: pau t a0103 N0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0547.lab vs ../HVITE-M/Rec-mau-M-24/mfG0547.rec
LAB: pau h o0102 r u0202 pau
REC: pau t o0102   u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0548.lab vs ../HVITE-M/Rec-mau-M-24/mfG0548.rec
LAB: pau               y o0102 u0202 pau
REC: pau b o0104 u0204 d o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0549.lab vs ../HVITE-M/Rec-mau-M-24/mfG0549.rec
LAB: pau         t e0103 m     a0203 e0303 pau
REC: pau k u0104 w a0204 e0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0550.lab vs ../HVITE-M/Rec-mau-M-24/mfG0550.rec
LAB: pau sh i0103 n         o0203 b u0303 pau
REC: pau k  i0104 n o0204 d o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0551.lab vs ../HVITE-M/Rec-mau-M-24/mfG0551.rec
LAB: pau h a0104 N0204 p a0304 ts u0404 pau
REC: pau k a0104 N0204 k a0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0552.lab vs ../HVITE-M/Rec-mau-M-24/mfG0552.rec
LAB: pau t       e0104 k i0204 t o0304 u0404 pau
REC: pau t a0103 i0203 k               o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0553.lab vs ../HVITE-M/Rec-mau-M-24/mfG0553.rec
LAB: pau k o0104 u0204 b     a0304 i0404 pau
REC: pau         g     u0103 a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0554.lab vs ../HVITE-M/Rec-mau-M-24/mfG0554.rec
LAB: pau ts u0103 i0203 n i0303 pau
REC: pau r  i0103 N0203 j i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0555.lab vs ../HVITE-M/Rec-mau-M-24/mfG0555.rec
LAB: pau s e0104 N0204 t a0304 k u0404 pau
REC: pau s       e0103 m a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0556.lab vs ../HVITE-M/Rec-mau-M-24/mfG0556.rec
LAB: pau         s a0102 k i0202 pau

REC: pau k u0103 s a0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0557.lab vs ../HVITE-M/Rec-mau-M-24/mfG0557.rec
LAB: pau k o0104 N0204 j     o0304 u0404 pau
REC: pau k       o0103 N0203 ky    o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0558.lab vs ../HVITE-M/Rec-mau-M-24/mfG0558.rec
LAB: pau ch i0103 g i0203 r               u0303 pau
REC: pau          k u0104 r i0204 i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0560.lab vs ../HVITE-M/Rec-mau-M-24/mfG0560.rec
LAB: pau m a0104 k i0204 k o0304 m u0404 pau
REC: pau         k a0103 k o0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0561.lab vs ../HVITE-M/Rec-mau-M-24/mfG0561.rec
LAB: pau m a0102   N0202 pau
REC: pau m a0102 g o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0563.lab vs ../HVITE-M/Rec-mau-M-24/mfG0563.rec
LAB: pau                 i0102 d o0202 pau
REC: pau k i0104 N0204 z o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0564.lab vs ../HVITE-M/Rec-mau-M-24/mfG0564.rec
LAB: pau h i0102 k         u0202 pau
REC: pau b a0103 k u0203 r o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0567.lab vs ../HVITE-M/Rec-mau-M-24/mfG0567.rec
LAB: pau h o0104 N0204 m o0304 n o0404 pau
REC: pau               m o0102 n o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0568.lab vs ../HVITE-M/Rec-mau-M-24/mfG0568.rec
LAB: pau n e0103 u0203 ch i0303 pau
REC: pau   k     u0102 sh i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0570.lab vs ../HVITE-M/Rec-mau-M-24/mfG0570.rec
LAB: pau i0103 g a0203 i0303 pau
REC: pau u0103 g a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0571.lab vs ../HVITE-M/Rec-mau-M-24/mfG0571.rec
LAB: pau j o0104 u0204 s e0304 i0404 pau
REC: pau g o0104 u0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0573.lab vs ../HVITE-M/Rec-mau-M-24/mfG0573.rec
LAB: pau f     u0102 r     i0202 pau
REC: pau u0103 g     a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0574.lab vs ../HVITE-M/Rec-mau-M-24/mfG0574.rec
LAB: pau k o0103 ky u0203 u0303 pau
REC: pau k u0103 r  u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0575.lab vs ../HVITE-M/Rec-mau-M-24/mfG0575.rec
LAB: pau       n a0103 b i0203 k u0303 pau

REC: pau u0104 r a0204 g i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0576.lab vs ../HVITE-M/Rec-mau-M-24/mfG0576.rec
LAB: pau k a0103 k e0203 r u0303 pau
REC: pau k a0103 s e0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0577.lab vs ../HVITE-M/Rec-mau-M-24/mfG0577.rec
LAB: pau ky o0104 u0204 d o0304 u0404 pau
REC: pau d  o0104 u0204 d o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0579.lab vs ../HVITE-M/Rec-mau-M-24/mfG0579.rec
LAB: pau m         i0103 k       a0203 N0303 pau
REC: pau m e0104 g a0204 k e0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0581.lab vs ../HVITE-M/Rec-mau-M-24/mfG0581.rec
LAB: pau f u0104 k i0204 d a0304 s u0404 pau
REC: pau s u0104   u0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0582.lab vs ../HVITE-M/Rec-mau-M-24/mfG0582.rec
LAB: pau   gy    o0104 u0204 r e0304 ts u0404 pau
REC: pau m o0103 d     o0203 r          u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0583.lab vs ../HVITE-M/Rec-mau-M-24/mfG0583.rec
LAB: pau k u0103 m     o0203 r u0303 pau
REC: pau   g     o0103 u0203 k a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0585.lab vs ../HVITE-M/Rec-mau-M-24/mfG0585.rec
LAB: pau n a0103 s a0203 k e0303 pau
REC: pau h a0103 s a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0586.lab vs ../HVITE-M/Rec-mau-M-24/mfG0586.rec
LAB: pau k i0103 ts             u0203 i0303 pau
REC: pau         ts u0103 i0203 t     e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0587.lab vs ../HVITE-M/Rec-mau-M-24/mfG0587.rec
LAB: pau y o0104 q0204 ky    u0304 u0404 pau
REC: pau f u0103 r     o0203 k     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0588.lab vs ../HVITE-M/Rec-mau-M-24/mfG0588.rec
LAB: pau j i0103 m e0203 N0303 pau
REC: pau m i0103 r e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0590.lab vs ../HVITE-M/Rec-mau-M-24/mfG0590.rec
LAB: pau       s o0103 d a0203 ts u0303 pau
REC: pau u0104 z u0204 m a0304 r  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0592.lab vs ../HVITE-M/Rec-mau-M-24/mfG0592.rec
LAB: pau n o0103 b o0203 r u0303 pau
REC: pau m o0103 d o0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0593.lab vs ../HVITE-M/Rec-mau-M-24/mfG0593.rec
LAB: pau          d a0102 m     u0202 pau

REC: pau sh u0103 d       a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0595.lab vs ../HVITE-M/Rec-mau-M-24/mfG0595.rec
LAB: pau               b o0102 N0202 pau
REC: pau z o0104 u0204 k i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0596.lab vs ../HVITE-M/Rec-mau-M-24/mfG0596.rec
LAB: pau   o0103   u0203 b o0303 pau
REC: pau t o0103 m o0203 n i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0598.lab vs ../HVITE-M/Rec-mau-M-24/mfG0598.rec
LAB: pau m e0103 d a0203 ts u0303 pau
REC: pau k u0103 g a0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0602.lab vs ../HVITE-M/Rec-mau-M-24/mfG0602.rec
LAB: pau t o0104 k u0204 t e0304 i0404 pau
REC: pau g o0104   u0204 r e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0603.lab vs ../HVITE-M/Rec-mau-M-24/mfG0603.rec
LAB: pau h i0103 r         a0203 k u0303 pau
REC: pau   u0104 r a0204 g i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0604.lab vs ../HVITE-M/Rec-mau-M-24/mfG0604.rec
LAB: pau   o0103 d o0203 r i0303 pau
REC: pau m i0103 b u0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0605.lab vs ../HVITE-M/Rec-mau-M-24/mfG0605.rec
LAB: pau k a0103 k a0203 r u0303 pau
REC: pau m a0103 z a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0606.lab vs ../HVITE-M/Rec-mau-M-24/mfG0606.rec
LAB: pau   sh    o0104 u0204 d o0304 k u0404 pau
REC: pau k i0104 n     o0204 d o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0607.lab vs ../HVITE-M/Rec-mau-M-24/mfG0607.rec
LAB: pau   h     e0102 r     i0202 pau
REC: pau g e0104 N0204 i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0610.lab vs ../HVITE-M/Rec-mau-M-24/mfG0610.rec
LAB: pau ts u0103 i0203 n i0303 pau
REC: pau k  a0103 i0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0611.lab vs ../HVITE-M/Rec-mau-M-24/mfG0611.rec
LAB: pau k a0103 k     e0203 r     u0303 pau
REC: pau r e0104 N0204 g     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0612.lab vs ../HVITE-M/Rec-mau-M-24/mfG0612.rec
LAB: pau ky o0104 u0204 ry o0304 k u0404 pau
REC: pau d  o0104 u0204 y  o0304   u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0614.lab vs ../HVITE-M/Rec-mau-M-24/mfG0614.rec
LAB: pau h o0104 u0204 s     o0304 u0404 pau

REC: pau         m     o0102 t     o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0615.lab vs ../HVITE-M/Rec-mau-M-24/mfG0615.rec
LAB: pau   t     a0103 i0203 k     o0303 pau
REC: pau k a0104 i0204 h     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0616.lab vs ../HVITE-M/Rec-mau-M-24/mfG0616.rec
LAB: pau h i0103 r a0203 k               u0303 pau
REC: pau         y a0104 k u0204 h i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0618.lab vs ../HVITE-M/Rec-mau-M-24/mfG0618.rec
LAB: pau h i0103 r o0203 b a0303 pau
REC: pau k u0103 r u0203 m a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0620.lab vs ../HVITE-M/Rec-mau-M-24/mfG0620.rec
LAB: pau n a0103 y a0203 m u0303 pau
REC: pau   u0103 r a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0621.lab vs ../HVITE-M/Rec-mau-M-24/mfG0621.rec
LAB: pau f u0104 k u0204 sh    u0304 u0404 pau
REC: pau k u0103 k       u0203 r     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0623.lab vs ../HVITE-M/Rec-mau-M-24/mfG0623.rec
LAB: pau s a0104 i0204 k e0304 ts u0404 pau
REC: pau k a0104 i0204 k e0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0624.lab vs ../HVITE-M/Rec-mau-M-24/mfG0624.rec
LAB: pau   a0103 N0203 j       i0303 pau
REC: pau t a0104 N0204 j o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0627.lab vs ../HVITE-M/Rec-mau-M-24/mfG0627.rec
LAB: pau z e0104 N0204 ch    o0304 u0404 pau
REC: pau   k     e0103 N0203 ky    o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0629.lab vs ../HVITE-M/Rec-mau-M-24/mfG0629.rec
LAB: pau   k     o0104 u0204 d o0304 u0404 pau
REC: pau y o0103 h     o0203 d       o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0630.lab vs ../HVITE-M/Rec-mau-M-24/mfG0630.rec
LAB: pau b o0104 u0204   e0304 i0404 pau
REC: pau g o0104 u0204 r e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0631.lab vs ../HVITE-M/Rec-mau-M-24/mfG0631.rec
LAB: pau s u0102 k i0202 pau
REC: pau b u0102 k i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0632.lab vs ../HVITE-M/Rec-mau-M-24/mfG0632.rec
LAB: pau    y     u0103 k i0203 k  i0303 pau
REC: pau ch u0104 u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0634.lab vs ../HVITE-M/Rec-mau-M-24/mfG0634.rec
LAB: pau         j u0103 u0203 j i0303 pau

REC: pau k u0104 r i0204 i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0635.lab vs ../HVITE-M/Rec-mau-M-24/mfG0635.rec
LAB: pau         y  o0102 ch i0202 pau
REC: pau k a0103 ts u0203 j  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0636.lab vs ../HVITE-M/Rec-mau-M-24/mfG0636.rec
LAB: pau   s     a0103 s     u0203 g     a0303 pau
REC: pau b a0104 k     u0204 d     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0637.lab vs ../HVITE-M/Rec-mau-M-24/mfG0637.rec
LAB: pau y a0104 b u0204 r e0304 r u0404 pau
REC: pau k a0104 k u0204 m e0304   i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0638.lab vs ../HVITE-M/Rec-mau-M-24/mfG0638.rec
LAB: pau n u0103 r a0203 s u0303 pau
REC: pau p u0103 r a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0640.lab vs ../HVITE-M/Rec-mau-M-24/mfG0640.rec
LAB: pau b e0103 N0203 r     i0303 pau
REC: pau g e0104 N0204 i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0641.lab vs ../HVITE-M/Rec-mau-M-24/mfG0641.rec
LAB: pau   sh    i0104 i0204 r e0304 r u0404 pau
REC: pau m o0104 ch    i0204   i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0642.lab vs ../HVITE-M/Rec-mau-M-24/mfG0642.rec
LAB: pau h a0103 n a0203 s u0303 pau
REC: pau k a0103 n a0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0644.lab vs ../HVITE-M/Rec-mau-M-24/mfG0644.rec
LAB: pau sh i0103 s a0203 ts u0303 pau
REC: pau          z a0102 ts u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0645.lab vs ../HVITE-M/Rec-mau-M-24/mfG0645.rec
LAB: pau b u0102 k i0202 pau
REC: pau m u0102 g i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0646.lab vs ../HVITE-M/Rec-mau-M-24/mfG0646.rec
LAB: pau                             i0102 t a0202 pau
REC: pau sh i0105 g a0205 m i0305 ts u0405 k u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0647.lab vs ../HVITE-M/Rec-mau-M-24/mfG0647.rec
LAB: pau   o0103 e0203 r u0303 pau
REC: pau s u0103 e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0648.lab vs ../HVITE-M/Rec-mau-M-24/mfG0648.rec
LAB: pau   o0104 b o0204 e0304 r u0404 pau
REC: pau k o0104 g o0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0651.lab vs ../HVITE-M/Rec-mau-M-24/mfG0651.rec
LAB: pau s u0102 g i0202 pau

REC: pau k u0102 g i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0653.lab vs ../HVITE-M/Rec-mau-M-24/mfG0653.rec
LAB: pau   ry    o0104 u0204 y o0304 u0404 pau
REC: pau r i0104 k     u0204 j o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0654.lab vs ../HVITE-M/Rec-mau-M-24/mfG0654.rec
LAB: pau k a0104 N0204 g e0304 k     i0404 pau
REC: pau b a0104 N0204 g       o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0655.lab vs ../HVITE-M/Rec-mau-M-24/mfG0655.rec
LAB: pau m i0103 e0203 r u0303 pau
REC: pau k o0103 u0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0656.lab vs ../HVITE-M/Rec-mau-M-24/mfG0656.rec
LAB: pau t o0103 n a0203 r                  i0303 pau
REC: pau t         o0104 r i0204 ts u0304 g u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0657.lab vs ../HVITE-M/Rec-mau-M-24/mfG0657.rec
LAB: pau n u0103 k e0203 r u0303 pau
REC: pau t a0103 b e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0658.lab vs ../HVITE-M/Rec-mau-M-24/mfG0658.rec
LAB: pau a0103 r  a0203 s  u0303 pau
REC: pau       sh a0102 ts u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0659.lab vs ../HVITE-M/Rec-mau-M-24/mfG0659.rec
LAB: pau h a0102 r u0202 pau
REC: pau k a0102   N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0660.lab vs ../HVITE-M/Rec-mau-M-24/mfG0660.rec
LAB: pau ch o0104 k     u0204 s e0304 N0404 pau
REC: pau    y     u0104 u0204 s e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0661.lab vs ../HVITE-M/Rec-mau-M-24/mfG0661.rec
LAB: pau k a0104 N0204 t e0304 N0404 pau
REC: pau k a0104 N0204 s e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0662.lab vs ../HVITE-M/Rec-mau-M-24/mfG0662.rec
LAB: pau s a0102 k i0202 pau
REC: pau k a0102 g i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0664.lab vs ../HVITE-M/Rec-mau-M-24/mfG0664.rec
LAB: pau s       e0104 i0204 t o0304   u0404 pau
REC: pau s o0104 r     e0204 t o0304 m o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0667.lab vs ../HVITE-M/Rec-mau-M-24/mfG0667.rec
LAB: pau h a0104 j i0204 m e0304 r u0404 pau
REC: pau s a0104 b i0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0668.lab vs ../HVITE-M/Rec-mau-M-24/mfG0668.rec
LAB: pau sh u0103 ry o0203 k u0303 pau

REC: pau m  i0103 b  u0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0669.lab vs ../HVITE-M/Rec-mau-M-24/mfG0669.rec
LAB: pau k o0104 u0204 d o0304 u0404 pau
REC: pau h o0104 u0204 d o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0670.lab vs ../HVITE-M/Rec-mau-M-24/mfG0670.rec
LAB: pau y o0103 s o0203 k u0303 pau
REC: pau n o0103 z o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0671.lab vs ../HVITE-M/Rec-mau-M-24/mfG0671.rec
LAB: pau j u0104 N0204 t     o0304 u0404 pau
REC: pau k i0103 m     o0203 n     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0673.lab vs ../HVITE-M/Rec-mau-M-24/mfG0673.rec
LAB: pau   e0104 i0204 s e0304   i0404 pau
REC: pau k e0104 q0204 s e0304 k i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0674.lab vs ../HVITE-M/Rec-mau-M-24/mfG0674.rec
LAB: pau   j     i0103 t a0203 i0303 pau
REC: pau b o0104 u0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0675.lab vs ../HVITE-M/Rec-mau-M-24/mfG0675.rec
LAB: pau sh o0104 u0204 m o0304 u0404 pau
REC: pau    t     o0102 m       o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0677.lab vs ../HVITE-M/Rec-mau-M-24/mfG0677.rec
LAB: pau   a0104 i0204 m a0304 i0404 pau
REC: pau s a0104 i0204 g a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0678.lab vs ../HVITE-M/Rec-mau-M-24/mfG0678.rec
LAB: pau   j     u0103 k e0203 N0303 pau
REC: pau b o0104 u0204 k e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0679.lab vs ../HVITE-M/Rec-mau-M-24/mfG0679.rec
LAB: pau r a0104 N0204 y  o0304 u0404 pau
REC: pau z a0104 N0204 gy o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0680.lab vs ../HVITE-M/Rec-mau-M-24/mfG0680.rec
LAB: pau k a0103 s e0203 g u0303 pau
REC: pau k a0103 k e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0683.lab vs ../HVITE-M/Rec-mau-M-24/mfG0683.rec
LAB: pau   g     a0103 N0203 sh         o0303 pau
REC: pau n a0104 n     i0204 sh i0304 r o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0684.lab vs ../HVITE-M/Rec-mau-M-24/mfG0684.rec
LAB: pau sh o0103 r u0203 i0303 pau
REC: pau h  i0103 k u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0685.lab vs ../HVITE-M/Rec-mau-M-24/mfG0685.rec
LAB: pau   g     u0103 a0203   i0303 pau

REC: pau k a0103 w     a0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0686.lab vs ../HVITE-M/Rec-mau-M-24/mfG0686.rec
LAB: pau          ts u0102 r u0202 pau
REC: pau sh i0103 z  u0203 k a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0687.lab vs ../HVITE-M/Rec-mau-M-24/mfG0687.rec
LAB: pau m  o0104 t a0204 r e0304 r u0404 pau
REC: pau ts u0104 k a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0691.lab vs ../HVITE-M/Rec-mau-M-24/mfG0691.rec
LAB: pau w a0103 z a0203 t o0303 pau
REC: pau m a0103 z a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0693.lab vs ../HVITE-M/Rec-mau-M-24/mfG0693.rec
LAB: pau y a0104 s a0204 sh    i0304 i0404 pau
REC: pau         g a0103 q0203 k     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0694.lab vs ../HVITE-M/Rec-mau-M-24/mfG0694.rec
LAB: pau ts u0104 t a0204 e0304 r u0404 pau
REC: pau ts u0104 k a0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0696.lab vs ../HVITE-M/Rec-mau-M-24/mfG0696.rec
LAB: pau o0103 k               u0203 r u0303 pau
REC: pau       k o0104 u0204 m u0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0697.lab vs ../HVITE-M/Rec-mau-M-24/mfG0697.rec
LAB: pau         s a0104 b i0204 r e0304 r u0404 pau
REC: pau k u0105 t a0205 b i0305 r e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0700.lab vs ../HVITE-M/Rec-mau-M-24/mfG0700.rec
LAB: pau s a0106 sh i0206 ts u0306 k a0406 e0506 r u0606 pau
REC: pau            a0104 z  u0204 k       a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0701.lab vs ../HVITE-M/Rec-mau-M-24/mfG0701.rec
LAB: pau k o0104 k u0204 r     i0304 ts u0404 pau
REC: pau         k       o0103 u0203 h  o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0704.lab vs ../HVITE-M/Rec-mau-M-24/mfG0704.rec
LAB: pau          j a0102 m a0202 pau
REC: pau sh i0103 r a0203 g a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0706.lab vs ../HVITE-M/Rec-mau-M-24/mfG0706.rec
LAB: pau h e0103 k o0203 m u0303 pau
REC: pau k a0103 k o0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0707.lab vs ../HVITE-M/Rec-mau-M-24/mfG0707.rec
LAB: pau a0106 r u0206 m i0306 n     i0406 u0506 m u0606 pau
REC: pau u0104 r               a0204 g     i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0709.lab vs ../HVITE-M/Rec-mau-M-24/mfG0709.rec
LAB: pau j u0103 N0203 b i0303 pau

REC: pau h e0103 N0203 j i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0710.lab vs ../HVITE-M/Rec-mau-M-24/mfG0710.rec
LAB: pau   t     e0103 i0203 sh    i0303 pau
REC: pau s e0104 i0204 k     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0711.lab vs ../HVITE-M/Rec-mau-M-24/mfG0711.rec
LAB: pau w a0103 r i0203 n i0303 pau
REC: pau k a0103   i0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0714.lab vs ../HVITE-M/Rec-mau-M-24/mfG0714.rec
LAB: pau n a0104 N0204 d a0304 k a0404 pau
REC: pau k a0104 N0204 k a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0716.lab vs ../HVITE-M/Rec-mau-M-24/mfG0716.rec
LAB: pau s o0104 k o0204 n a0304   u0404 pau
REC: pau k o0104 r o0204 g a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0717.lab vs ../HVITE-M/Rec-mau-M-24/mfG0717.rec
LAB: pau h  i0103 r u0203 m         a0303 pau
REC: pau sh i0104 z u0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0718.lab vs ../HVITE-M/Rec-mau-M-24/mfG0718.rec
LAB: pau         ry o0103 u0203 sh i0303 pau
REC: pau k i0103 ry       o0203 k  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0720.lab vs ../HVITE-M/Rec-mau-M-24/mfG0720.rec
LAB: pau k e0104 i0204 s a0304 ts u0404 pau
REC: pau k e0104 q0204 s a0304 k  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0721.lab vs ../HVITE-M/Rec-mau-M-24/mfG0721.rec
LAB: pau u0104 ch i0204 k o0304 m u0404 pau
REC: pau       m  i0103 k o0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0722.lab vs ../HVITE-M/Rec-mau-M-24/mfG0722.rec
LAB: pau u0104 k  e0204 t o0304 r u0404 pau
REC: pau       ts u0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0723.lab vs ../HVITE-M/Rec-mau-M-24/mfG0723.rec
LAB: pau k u0102 m i0202 pau
REC: pau n u0102 g u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0724.lab vs ../HVITE-M/Rec-mau-M-24/mfG0724.rec
LAB: pau       sh u0103 gy    o0203 u0303 pau
REC: pau u0103 sh       i0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0725.lab vs ../HVITE-M/Rec-mau-M-24/mfG0725.rec
LAB: pau    i0103 t a0203 m                  i0303 pau
REC: pau sh i0105 g a0205 m i0305 ts u0405 k u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0726.lab vs ../HVITE-M/Rec-mau-M-24/mfG0726.rec
LAB: pau h o0104 u0204 t a0304 i0404 pau

REC: pau   sh    o0103 t a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0727.lab vs ../HVITE-M/Rec-mau-M-24/mfG0727.rec
LAB: pau         k u0102 n i0202 pau
REC: pau m i0103 d a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0728.lab vs ../HVITE-M/Rec-mau-M-24/mfG0728.rec
LAB: pau   ts    u0103 k u0203 e0303 pau
REC: pau s e0104 N0204 g e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0729.lab vs ../HVITE-M/Rec-mau-M-24/mfG0729.rec
LAB: pau e0103 N0203 k i0303 pau
REC: pau e0103 N0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0730.lab vs ../HVITE-M/Rec-mau-M-24/mfG0730.rec
LAB: pau          k o0104 u0204 sh i0304 N0404 pau
REC: pau ch o0103 k                i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0731.lab vs ../HVITE-M/Rec-mau-M-24/mfG0731.rec
LAB: pau s a0104 i0204 t e0304 i0404 pau
REC: pau b a0104 i0204 t e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0733.lab vs ../HVITE-M/Rec-mau-M-24/mfG0733.rec
LAB: pau h i0103 k a0203 k u0303 pau
REC: pau         d a0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0734.lab vs ../HVITE-M/Rec-mau-M-24/mfG0734.rec
LAB: pau s a0104 q0204 s u0304 r u0404 pau
REC: pau   ch    i0103 g i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0735.lab vs ../HVITE-M/Rec-mau-M-24/mfG0735.rec
LAB: pau m a0103 s  u0203 i0303 pau
REC: pau   a0103 ts u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0736.lab vs ../HVITE-M/Rec-mau-M-24/mfG0736.rec
LAB: pau s a0103 i0203 f u0303 pau
REC: pau t a0103 i0203 p u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0737.lab vs ../HVITE-M/Rec-mau-M-24/mfG0737.rec
LAB: pau g a0104 q0204 p e0304 i0404 pau
REC: pau   t     a0103 t o0203 e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0738.lab vs ../HVITE-M/Rec-mau-M-24/mfG0738.rec
LAB: pau   o0103 r o0203 s u0303 pau
REC: pau s o0103 m u0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0739.lab vs ../HVITE-M/Rec-mau-M-24/mfG0739.rec
LAB: pau n i0104 N0204 m e0304 i0404 pau
REC: pau         u0102 m       e0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0743.lab vs ../HVITE-M/Rec-mau-M-24/mfG0743.rec
LAB: pau k a0103 k o0203 m u0303 pau

REC: pau k a0103 k u0203 d o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0744.lab vs ../HVITE-M/Rec-mau-M-24/mfG0744.rec
LAB: pau m                i0102 ch i0202 pau
REC: pau m u0104 i0204 sh i0304 k  i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0745.lab vs ../HVITE-M/Rec-mau-M-24/mfG0745.rec
LAB: pau   u0103 ch i0203 w a0303 pau
REC: pau k i0103 m  a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0746.lab vs ../HVITE-M/Rec-mau-M-24/mfG0746.rec
LAB: pau i0104 n e0204 m     u0304 r i0404 pau
REC: pau         g     e0103 N0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0748.lab vs ../HVITE-M/Rec-mau-M-24/mfG0748.rec
LAB: pau k a0104 N0204 ch    o0304 u0404 pau
REC: pau   g     a0103 N0203 sh    o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0749.lab vs ../HVITE-M/Rec-mau-M-24/mfG0749.rec
LAB: pau s               o0102 b o0202 pau
REC: pau s o0104 u0204 z o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0750.lab vs ../HVITE-M/Rec-mau-M-24/mfG0750.rec
LAB: pau m o0103 d e0203 r u0303 pau
REC: pau k a0103 b u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0752.lab vs ../HVITE-M/Rec-mau-M-24/mfG0752.rec
LAB: pau k  o0104 u0204 ts u0304 u0404 pau
REC: pau ch o0104 u0204 sh u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0753.lab vs ../HVITE-M/Rec-mau-M-24/mfG0753.rec
LAB: pau m a0103 g e0203 r u0303 pau
REC: pau s a0103 g u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0755.lab vs ../HVITE-M/Rec-mau-M-24/mfG0755.rec
LAB: pau ky o0104 u0204 s e0304 i0404 pau
REC: pau g  o0104 u0204 k e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0757.lab vs ../HVITE-M/Rec-mau-M-24/mfG0757.rec
LAB: pau   i0104 q0204 t a0304 N0404 pau
REC: pau s e0104 i0204 s a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0758.lab vs ../HVITE-M/Rec-mau-M-24/mfG0758.rec
LAB: pau    u0104 t a0204 g a0304 i0404 pau
REC: pau sh i0104 h a0204 r a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0760.lab vs ../HVITE-M/Rec-mau-M-24/mfG0760.rec
LAB: pau h i0104 q0204 k o0304 sh i0404 pau
REC: pau y u0103 u0203 k          i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0761.lab vs ../HVITE-M/Rec-mau-M-24/mfG0761.rec
LAB: pau y  a0103 k e0203 d o0303 pau

REC: pau sh a0103 k u0203 d o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0762.lab vs ../HVITE-M/Rec-mau-M-24/mfG0762.rec
LAB: pau               d o0102 a0202 pau
REC: pau j o0104 u0204 h i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0763.lab vs ../HVITE-M/Rec-mau-M-24/mfG0763.rec
LAB: pau   a0102 r u0202 pau
REC: pau h a0102 r u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0764.lab vs ../HVITE-M/Rec-mau-M-24/mfG0764.rec
LAB: pau k a0103 ch i0203 k u0303 pau
REC: pau d a0103    i0203 b u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0765.lab vs ../HVITE-M/Rec-mau-M-24/mfG0765.rec
LAB: pau k i0103 k i0203 N0303 pau
REC: pau b o0103 k i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0766.lab vs ../HVITE-M/Rec-mau-M-24/mfG0766.rec
LAB: pau         sh o0102 u0202 pau
REC: pau r i0103 k  o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0767.lab vs ../HVITE-M/Rec-mau-M-24/mfG0767.rec
LAB: pau   k     e0102 ch i0202 pau
REC: pau n e0103 u0203 ch i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0768.lab vs ../HVITE-M/Rec-mau-M-24/mfG0768.rec
LAB: pau k o0103 u0203 g i0303 pau
REC: pau   m     o0102 r i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0769.lab vs ../HVITE-M/Rec-mau-M-24/mfG0769.rec
LAB: pau s o0103 m a0203 ts u0303 pau
REC: pau t o0103 m a0203 r  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0770.lab vs ../HVITE-M/Rec-mau-M-24/mfG0770.rec
LAB: pau k a0103 k     o0203 u0303 pau
REC: pau k a0103 N0203 g     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0771.lab vs ../HVITE-M/Rec-mau-M-24/mfG0771.rec
LAB: pau k o0102 b  u0202 pau
REC: pau m u0102 ch i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0773.lab vs ../HVITE-M/Rec-mau-M-24/mfG0773.rec
LAB: pau         g e0102 sh a0202 pau
REC: pau h i0103 g e0203 k  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0774.lab vs ../HVITE-M/Rec-mau-M-24/mfG0774.rec
LAB: pau k u0103 g a0203 ts u0303 pau
REC: pau k u0103 r a0203 s  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0776.lab vs ../HVITE-M/Rec-mau-M-24/mfG0776.rec
LAB: pau           b     a0102 N0202 pau

REC: pau k a0103 n a0203 g     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0777.lab vs ../HVITE-M/Rec-mau-M-24/mfG0777.rec
LAB: pau         m     e0102 i0202 pau
REC: pau g e0104 N0204 i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0778.lab vs ../HVITE-M/Rec-mau-M-24/mfG0778.rec
LAB: pau n  a0103 y a0203 m u0303 pau
REC: pau sh a0103 g a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0779.lab vs ../HVITE-M/Rec-mau-M-24/mfG0779.rec
LAB: pau          k  e0103 m o0203 n     o0303 pau
REC: pau sh i0104 ts u0204 m       o0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0781.lab vs ../HVITE-M/Rec-mau-M-24/mfG0781.rec
LAB: pau m a0103 ts u0203 r u0303 pau
REC: pau         w  a0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0782.lab vs ../HVITE-M/Rec-mau-M-24/mfG0782.rec
LAB: pau s o0104 k u0204 ry    o0304 u0404 pau
REC: pau f u0103 k       u0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0783.lab vs ../HVITE-M/Rec-mau-M-24/mfG0783.rec
LAB: pau y o0103 t e0203 i0303 pau
REC: pau y u0103 k u0203 e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0784.lab vs ../HVITE-M/Rec-mau-M-24/mfG0784.rec
LAB: pau m a0103 g e0203 r u0303 pau
REC: pau         g a0102 r a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0785.lab vs ../HVITE-M/Rec-mau-M-24/mfG0785.rec
LAB: pau s o0102 r a0202 pau
REC: pau h a0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0786.lab vs ../HVITE-M/Rec-mau-M-24/mfG0786.rec
LAB: pau         y     o0103 u0203 k i0303 pau
REC: pau r o0104 N0204 j     i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0787.lab vs ../HVITE-M/Rec-mau-M-24/mfG0787.rec
LAB: pau ch o0104 k u0204 s e0304 N0404 pau
REC: pau ky o0104 k u0204 s e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0789.lab vs ../HVITE-M/Rec-mau-M-24/mfG0789.rec
LAB: pau m o0104 N0204 d a0304 i0404 pau
REC: pau k o0104 u0204 g a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0790.lab vs ../HVITE-M/Rec-mau-M-24/mfG0790.rec
LAB: pau s o0102 t               o0202 pau
REC: pau s u0104 t o0204 o0304 b u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0791.lab vs ../HVITE-M/Rec-mau-M-24/mfG0791.rec
LAB: pau s e0104 i0204 sh i0304 N0404 pau

REC: pau k e0104 q0204 sh i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0792.lab vs ../HVITE-M/Rec-mau-M-24/mfG0792.rec
LAB: pau   o0103 k o0203 r u0303 pau
REC: pau k o0103   o0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0794.lab vs ../HVITE-M/Rec-mau-M-24/mfG0794.rec
LAB: pau t o0105 r i0205 sh i0305 m a0405 r     u0505 pau
REC: pau                 ts u0104 k i0204 a0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0795.lab vs ../HVITE-M/Rec-mau-M-24/mfG0795.rec
LAB: pau         o0103 m o0203 n  i0303 pau
REC: pau k o0104 u0204 b u0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0797.lab vs ../HVITE-M/Rec-mau-M-24/mfG0797.rec
LAB: pau a0103 r a0203 s u0303 pau
REC: pau a0103   i0203 z u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0798.lab vs ../HVITE-M/Rec-mau-M-24/mfG0798.rec
LAB: pau               o0103 o0203 i0303 pau
REC: pau d o0104 u0204 y     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0799.lab vs ../HVITE-M/Rec-mau-M-24/mfG0799.rec
LAB: pau h o0104 u0204 r i0304 ts u0404 pau
REC: pau k o0104 u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0800.lab vs ../HVITE-M/Rec-mau-M-24/mfG0800.rec
LAB: pau   a0103 N0203 j       i0303 pau
REC: pau g a0104 N0204 j o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0802.lab vs ../HVITE-M/Rec-mau-M-24/mfG0802.rec
LAB: pau               a0102 i0202 pau
REC: pau t a0103 i0203 r     a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0803.lab vs ../HVITE-M/Rec-mau-M-24/mfG0803.rec
LAB: pau   t     a0103 b e0203 r     u0303 pau
REC: pau d a0104 i0204 b       e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0804.lab vs ../HVITE-M/Rec-mau-M-24/mfG0804.rec
LAB: pau   n     i0102 j     i0202 pau
REC: pau g e0103 h     i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0805.lab vs ../HVITE-M/Rec-mau-M-24/mfG0805.rec
LAB: pau h a0104 i0204 b o0304 k u0404 pau
REC: pau g a0104 i0204 k o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0806.lab vs ../HVITE-M/Rec-mau-M-24/mfG0806.rec
LAB: pau         h a0102 sh i0202 pau
REC: pau g a0103 k u0203 sh i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0807.lab vs ../HVITE-M/Rec-mau-M-24/mfG0807.rec
LAB: pau s a0105 k a0205 n o0305 b o0405 r u0505 pau

REC: pau         k a0104 t a0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0808.lab vs ../HVITE-M/Rec-mau-M-24/mfG0808.rec
LAB: pau t e0104 i0204 a0304 N0404 pau
REC: pau   d     e0103 a0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0813.lab vs ../HVITE-M/Rec-mau-M-24/mfG0813.rec
LAB: pau sh u0104 u0204 sh o0304 k     u0404 pau
REC: pau ky u0104 u0204 sh       u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0814.lab vs ../HVITE-M/Rec-mau-M-24/mfG0814.rec
LAB: pau sh a0103 k u0203 d o0303 pau
REC: pau k  a0103 k u0203 d o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0815.lab vs ../HVITE-M/Rec-mau-M-24/mfG0815.rec
LAB: pau b  u0103 j     o0203 k     u0303 pau
REC: pau ts u0104 i0204 ky    u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0816.lab vs ../HVITE-M/Rec-mau-M-24/mfG0816.rec
LAB: pau k a0102 k u0202 pau
REC: pau k a0102   N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0817.lab vs ../HVITE-M/Rec-mau-M-24/mfG0817.rec
LAB: pau n u0103 r u0203 i0303 pau
REC: pau j i0103 b u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0818.lab vs ../HVITE-M/Rec-mau-M-24/mfG0818.rec
LAB: pau y a0104 k u0204 n i0304 N0404 pau
REC: pau y a0104 k u0204 h i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0819.lab vs ../HVITE-M/Rec-mau-M-24/mfG0819.rec
LAB: pau n e0103 z  u0203 m i0303 pau
REC: pau   i0103 ts u0203 m o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0820.lab vs ../HVITE-M/Rec-mau-M-24/mfG0820.rec
LAB: pau k o0104 u0204 m u0304 r u0404 pau
REC: pau   s     o0103 m u0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0821.lab vs ../HVITE-M/Rec-mau-M-24/mfG0821.rec
LAB: pau h a0104 N0204 s u0304 r       u0404 pau
REC: pau               k u0103 r u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0822.lab vs ../HVITE-M/Rec-mau-M-24/mfG0822.rec
LAB: pau o0103 k o0203 s  u0303 pau
REC: pau       k o0102 ts u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0823.lab vs ../HVITE-M/Rec-mau-M-24/mfG0823.rec
LAB: pau s u0103 t     e0203 r     u0303 pau
REC: pau   b     o0104 u0204 e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0825.lab vs ../HVITE-M/Rec-mau-M-24/mfG0825.rec
LAB: pau k a0102 w a0202 pau

REC: pau n a0102 m a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0828.lab vs ../HVITE-M/Rec-mau-M-24/mfG0828.rec
LAB: pau k o0103 r i0203 r u0303 pau
REC: pau m o0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0829.lab vs ../HVITE-M/Rec-mau-M-24/mfG0829.rec
LAB: pau   o0103 e0203 r u0303 pau
REC: pau s u0103 e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0830.lab vs ../HVITE-M/Rec-mau-M-24/mfG0830.rec
LAB: pau         u0102 m e0202 pau
REC: pau t e0103 i0203 r e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0831.lab vs ../HVITE-M/Rec-mau-M-24/mfG0831.rec
LAB: pau m a0103 q0203 k a0303 pau
REC: pau   b     a0102 k a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0832.lab vs ../HVITE-M/Rec-mau-M-24/mfG0832.rec
LAB: pau           a0102 k u0202 pau
REC: pau k u0103 d a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0833.lab vs ../HVITE-M/Rec-mau-M-24/mfG0833.rec
LAB: pau         k o0102 t o0202 pau
REC: pau s o0103 n o0203 t a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0834.lab vs ../HVITE-M/Rec-mau-M-24/mfG0834.rec
LAB: pau a0104 k a0204 r       u0304 i0404 pau
REC: pau       b a0103 r i0203 k     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0837.lab vs ../HVITE-M/Rec-mau-M-24/mfG0837.rec
LAB: pau                 i0102 n u0202 pau
REC: pau k i0104 N0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0838.lab vs ../HVITE-M/Rec-mau-M-24/mfG0838.rec
LAB: pau o0104 k     u0204 by    o0304 u0404 pau
REC: pau m     u0103 sh    i0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0839.lab vs ../HVITE-M/Rec-mau-M-24/mfG0839.rec
LAB: pau         f u0102 r i0202 pau
REC: pau d e0103 k i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0840.lab vs ../HVITE-M/Rec-mau-M-24/mfG0840.rec
LAB: pau sh u0103 gy    o0203 u0303 pau
REC: pau k  o0103 u0203 ry    o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0841.lab vs ../HVITE-M/Rec-mau-M-24/mfG0841.rec
LAB: pau h e0104 d a0204 t a0304 r u0404 pau
REC: pau         k a0103 z a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0842.lab vs ../HVITE-M/Rec-mau-M-24/mfG0842.rec
LAB: pau m i0104 k a0204 k e0304 r u0404 pau

REC: pau m e0104 g a0204 k e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0844.lab vs ../HVITE-M/Rec-mau-M-24/mfG0844.rec
LAB: pau          s a0103 g u0203 r u0303 pau
REC: pau ch i0104 k a0204 g o0304 r o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0845.lab vs ../HVITE-M/Rec-mau-M-24/mfG0845.rec
LAB: pau k a0104 t a0204 m e0304 r u0404 pau
REC: pau   a0104 b a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0846.lab vs ../HVITE-M/Rec-mau-M-24/mfG0846.rec
LAB: pau           u0103 g  o0203 k i0303 pau
REC: pau t a0104 n o0204 sh i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0847.lab vs ../HVITE-M/Rec-mau-M-24/mfG0847.rec
LAB: pau           sh    i0101 pau
REC: pau k i0103 k i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0848.lab vs ../HVITE-M/Rec-mau-M-24/mfG0848.rec
LAB: pau h i0103 k         a0203 r i0303 pau
REC: pau         k a0103 r i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0849.lab vs ../HVITE-M/Rec-mau-M-24/mfG0849.rec
LAB: pau         y u0102 m i0202 pau
REC: pau t a0103 k u0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0850.lab vs ../HVITE-M/Rec-mau-M-24/mfG0850.rec
LAB: pau k o0103 u0203 ry    o0303 pau
REC: pau m o0103 y     o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0851.lab vs ../HVITE-M/Rec-mau-M-24/mfG0851.rec
LAB: pau         k o0104 u0204 ts    u0304 u0404 pau
REC: pau h i0104 t o0204 m     a0304 z     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0852.lab vs ../HVITE-M/Rec-mau-M-24/mfG0852.rec
LAB: pau sh u0103 j     u0203 ts    u0303 pau
REC: pau    r     i0103 k     o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0853.lab vs ../HVITE-M/Rec-mau-M-24/mfG0853.rec
LAB: pau   t     o0102 b          u0202 pau
REC: pau k o0104 u0204 b u0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0854.lab vs ../HVITE-M/Rec-mau-M-24/mfG0854.rec
LAB: pau   a0104 j i0204 w a0304   u0404 pau
REC: pau m a0104 j i0204 w a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0855.lab vs ../HVITE-M/Rec-mau-M-24/mfG0855.rec
LAB: pau u0105 s u0205 g u0305 r     a0405 i0505 pau
REC: pau               g a0103 i0203 b     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0856.lab vs ../HVITE-M/Rec-mau-M-24/mfG0856.rec
LAB: pau ts u0102 n                 a0202 pau

REC: pau          n a0103 k a0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0858.lab vs ../HVITE-M/Rec-mau-M-24/mfG0858.rec
LAB: pau h a0104 s a0204 m a0304 r u0404 pau
REC: pau k a0104 t a0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0859.lab vs ../HVITE-M/Rec-mau-M-24/mfG0859.rec
LAB: pau g e0103 N0203 sh i0303 pau
REC: pau d e0103 N0203 sh i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0860.lab vs ../HVITE-M/Rec-mau-M-24/mfG0860.rec
LAB: pau k e0104 N0204 t o0304 u0404 pau
REC: pau k e0104 N0204 s o0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0862.lab vs ../HVITE-M/Rec-mau-M-24/mfG0862.rec
LAB: pau   k     a0103 y o0203 u0303 pau
REC: pau s a0104 i0204 b o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0863.lab vs ../HVITE-M/Rec-mau-M-24/mfG0863.rec
LAB: pau b e0103 q0203 d     o0303 pau
REC: pau g e0103 j     u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0866.lab vs ../HVITE-M/Rec-mau-M-24/mfG0866.rec
LAB: pau y u0104 u0204 k a0304 i0404 pau
REC: pau b u0104 N0204 k a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0867.lab vs ../HVITE-M/Rec-mau-M-24/mfG0867.rec
LAB: pau y u0103 s  u0203 g u0303 pau
REC: pau         ts u0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0868.lab vs ../HVITE-M/Rec-mau-M-24/mfG0868.rec
LAB: pau y u0103 k                i0203 k i0303 pau
REC: pau         k e0104 N0204 sh i0304 k i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0872.lab vs ../HVITE-M/Rec-mau-M-24/mfG0872.rec
LAB: pau t  o0104 u0204 k e0304 i0404 pau
REC: pau ch u0104 u0204 k e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0873.lab vs ../HVITE-M/Rec-mau-M-24/mfG0873.rec
LAB: pau g o0103 k a0203 i0303 pau
REC: pau k o0103 t a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0874.lab vs ../HVITE-M/Rec-mau-M-24/mfG0874.rec
LAB: pau                 t     o0102 o0202 pau
REC: pau k o0104 u0204 m u0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0877.lab vs ../HVITE-M/Rec-mau-M-24/mfG0877.rec
LAB: pau                 h o0102 j o0202 pau
REC: pau k o0104 k u0204 h a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0879.lab vs ../HVITE-M/Rec-mau-M-24/mfG0879.rec
LAB: pau   k     o0103 t e0203 N0303 pau

REC: pau h o0104 u0204 g e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0880.lab vs ../HVITE-M/Rec-mau-M-24/mfG0880.rec
LAB: pau k  i0102 z         u0202 pau
REC: pau ts u0103 z u0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0881.lab vs ../HVITE-M/Rec-mau-M-24/mfG0881.rec
LAB: pau s e0104 i0204 s a0304 N0404 pau
REC: pau g e0104 N0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0883.lab vs ../HVITE-M/Rec-mau-M-24/mfG0883.rec
LAB: pau ch o0104 u0204 m e0304 N0404 pau
REC: pau h  o0104 u0204 g e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0885.lab vs ../HVITE-M/Rec-mau-M-24/mfG0885.rec
LAB: pau y o0104 r o0204 sh    i0304 i0404 pau
REC: pau         j o0103 u0203 k     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0886.lab vs ../HVITE-M/Rec-mau-M-24/mfG0886.rec
LAB: pau b o0104 u0204 d a0304 i0404 pau
REC: pau m o0104 N0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0887.lab vs ../HVITE-M/Rec-mau-M-24/mfG0887.rec
LAB: pau t e0103 N0203 k i0303 pau
REC: pau b e0103 N0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0888.lab vs ../HVITE-M/Rec-mau-M-24/mfG0888.rec
LAB: pau         h a0102 k o0202 pau
REC: pau n a0103 g o0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0890.lab vs ../HVITE-M/Rec-mau-M-24/mfG0890.rec
LAB: pau f u0104 N0204 sh i0304 ts u0404 pau
REC: pau j u0104 N0204 sh i0304    N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0891.lab vs ../HVITE-M/Rec-mau-M-24/mfG0891.rec
LAB: pau h a0104 n a0204 r e0304 r u0404 pau
REC: pau n a0104 g a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0893.lab vs ../HVITE-M/Rec-mau-M-24/mfG0893.rec
LAB: pau u0103 r a0203 m u0303 pau
REC: pau o0103 g a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0894.lab vs ../HVITE-M/Rec-mau-M-24/mfG0894.rec
LAB: pau ts u0103 k i0203 h i0303 pau
REC: pau          k i0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0895.lab vs ../HVITE-M/Rec-mau-M-24/mfG0895.rec
LAB: pau k u0103 r     o0203 u0303 pau
REC: pau   b     o0102 ch    i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0896.lab vs ../HVITE-M/Rec-mau-M-24/mfG0896.rec
LAB: pau h i0103 r e0203 i0303 pau

REC: pau         s e0102 N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0897.lab vs ../HVITE-M/Rec-mau-M-24/mfG0897.rec
LAB: pau w a0103 k         a0203 i0303 pau
REC: pau b a0104 k u0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0898.lab vs ../HVITE-M/Rec-mau-M-24/mfG0898.rec
LAB: pau   s     u0102 u0202 pau
REC: pau r i0102 ts    u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0900.lab vs ../HVITE-M/Rec-mau-M-24/mfG0900.rec
LAB: pau s a0103 k         e0203 r u0303 pau
REC: pau   u0104 k a0204 b e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0901.lab vs ../HVITE-M/Rec-mau-M-24/mfG0901.rec
LAB: pau ky o0104 u0204 d o0304 u0404 pau
REC: pau ch o0104 u0204 h o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0902.lab vs ../HVITE-M/Rec-mau-M-24/mfG0902.rec
LAB: pau o0103 e0203 r         u0303 pau
REC: pau b     e0103 r u0203 t o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0903.lab vs ../HVITE-M/Rec-mau-M-24/mfG0903.rec
LAB: pau k a0104 i0204 t e0304    i0404 pau
REC: pau r a0104 i0204 g e0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0904.lab vs ../HVITE-M/Rec-mau-M-24/mfG0904.rec
LAB: pau h e0103 k o0203 m u0303 pau
REC: pau t a0103 n o0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0906.lab vs ../HVITE-M/Rec-mau-M-24/mfG0906.rec
LAB: pau s o0104 u0204 z o0304 u0404 pau
REC: pau k o0104 u0204 z o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0908.lab vs ../HVITE-M/Rec-mau-M-24/mfG0908.rec
LAB: pau t o0104 u0204 k e0304 i0404 pau
REC: pau b o0104 u0204   e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0909.lab vs ../HVITE-M/Rec-mau-M-24/mfG0909.rec
LAB: pau s e0104 N0204 s e0304 i0404 pau
REC: pau   e0104 N0204 g e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0910.lab vs ../HVITE-M/Rec-mau-M-24/mfG0910.rec
LAB: pau ry o0104 u0204 i0304 k i0404 pau
REC: pau    h     o0103 N0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0911.lab vs ../HVITE-M/Rec-mau-M-24/mfG0911.rec
LAB: pau   a0103 i0203 d       a0303 pau
REC: pau k a0104 i0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0912.lab vs ../HVITE-M/Rec-mau-M-24/mfG0912.rec
LAB: pau k o0105 r a0205 sh i0305 m e0405 r u0505 pau

REC: pau   i0105 m a0205 sh i0305 m e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0913.lab vs ../HVITE-M/Rec-mau-M-24/mfG0913.rec
LAB: pau sh i0104 t a0204 g a0304 u0404 pau
REC: pau    u0104 k a0204 g a0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0915.lab vs ../HVITE-M/Rec-mau-M-24/mfG0915.rec
LAB: pau h i0103 n     i0203 ch    i0303 pau
REC: pau k a0104 N0204 sh    i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0918.lab vs ../HVITE-M/Rec-mau-M-24/mfG0918.rec
LAB: pau h     o0103 u0203 k i0303 pau
REC: pau u0103 g     o0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0919.lab vs ../HVITE-M/Rec-mau-M-24/mfG0919.rec
LAB: pau sh i0103 h o0203   N0303 pau
REC: pau sh i0103 k o0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0920.lab vs ../HVITE-M/Rec-mau-M-24/mfG0920.rec
LAB: pau s u0105 r e0205 ch i0305 g a0405 u0505 pau
REC: pau k i0103 r                  a0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0922.lab vs ../HVITE-M/Rec-mau-M-24/mfG0922.rec
LAB: pau   u0102 sh         i0202 pau
REC: pau h a0103 sh i0203 g o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0923.lab vs ../HVITE-M/Rec-mau-M-24/mfG0923.rec
LAB: pau s e0104 i0204 d a0304 i0404 pau
REC: pau g e0104 N0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0924.lab vs ../HVITE-M/Rec-mau-M-24/mfG0924.rec
LAB: pau s o0104 q0204 ch o0304 k u0404 pau
REC: pau s o0104 u0204 sh o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0925.lab vs ../HVITE-M/Rec-mau-M-24/mfG0925.rec
LAB: pau t a0103   e0203 r u0303 pau
REC: pau k a0103 g e0203   N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0926.lab vs ../HVITE-M/Rec-mau-M-24/mfG0926.rec
LAB: pau       t e0103 m a0203 e0303 pau
REC: pau u0104 d e0204 m a0304 e0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0927.lab vs ../HVITE-M/Rec-mau-M-24/mfG0927.rec
LAB: pau n e0103 m u0203 r u0303 pau
REC: pau g o0102 m         u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0928.lab vs ../HVITE-M/Rec-mau-M-24/mfG0928.rec
LAB: pau               o0102 t     o0202 pau
REC: pau r o0104 u0204 d     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0930.lab vs ../HVITE-M/Rec-mau-M-24/mfG0930.rec
LAB: pau m o0103   u0203 k e0303 pau

REC: pau m o0103 t o0203 d e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0931.lab vs ../HVITE-M/Rec-mau-M-24/mfG0931.rec
LAB: pau   a0103 i0203 z u0303 pau
REC: pau t a0103 i0203 h o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0932.lab vs ../HVITE-M/Rec-mau-M-24/mfG0932.rec
LAB: pau h i0104 q0204 p a0304 r u0404 pau
REC: pau   k     u0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0933.lab vs ../HVITE-M/Rec-mau-M-24/mfG0933.rec
LAB: pau                         g o0101 pau
REC: pau m o0104 t a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0934.lab vs ../HVITE-M/Rec-mau-M-24/mfG0934.rec
LAB: pau a0103 sh i0203 t a0303 pau
REC: pau i0103 n  a0203 k a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0936.lab vs ../HVITE-M/Rec-mau-M-24/mfG0936.rec
LAB: pau m i0102 r       u0202 pau
REC: pau m i0103 r e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0937.lab vs ../HVITE-M/Rec-mau-M-24/mfG0937.rec
LAB: pau n e0104 k     u0204 t a0304 i0404 pau
REC: pau n       i0104 N0204 t a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0938.lab vs ../HVITE-M/Rec-mau-M-24/mfG0938.rec
LAB: pau ky u0104 u0204 sh u0304 u0404 pau
REC: pau g  u0104 N0204 sh u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0939.lab vs ../HVITE-M/Rec-mau-M-24/mfG0939.rec
LAB: pau k a0103 r         i0203 r u0303 pau
REC: pau   u0104 r a0204 g i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0940.lab vs ../HVITE-M/Rec-mau-M-24/mfG0940.rec
LAB: pau k a0103 t a0203 ch i0303 pau
REC: pau b u0103 r a0203 sh i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0941.lab vs ../HVITE-M/Rec-mau-M-24/mfG0941.rec
LAB: pau sh o0104 u0204 j i0304 k i0404 pau
REC: pau k  o0104 o0204 h i0304   i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0942.lab vs ../HVITE-M/Rec-mau-M-24/mfG0942.rec
LAB: pau         s a0102 r u0202 pau
REC: pau k u0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0943.lab vs ../HVITE-M/Rec-mau-M-24/mfG0943.rec
LAB: pau e0105 ch i0205 k e0305 q0405 t o0505 pau
REC: pau       r  o0104 k e0204 q0304 t o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0944.lab vs ../HVITE-M/Rec-mau-M-24/mfG0944.rec
LAB: pau s a0103 b  a0203 k                 u0303 pau

REC: pau         ts u0104 k a0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0945.lab vs ../HVITE-M/Rec-mau-M-24/mfG0945.rec
LAB: pau d a0103 i0203 b u0303 pau
REC: pau g a0103 i0203 b u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0947.lab vs ../HVITE-M/Rec-mau-M-24/mfG0947.rec
LAB: pau sh o0104 u0204 gy    o0304 u0404 pau
REC: pau          k     o0102 r     o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0948.lab vs ../HVITE-M/Rec-mau-M-24/mfG0948.rec
LAB: pau y o0104 u0204 j i0304 N0404 pau
REC: pau r o0104 u0204 j i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0950.lab vs ../HVITE-M/Rec-mau-M-24/mfG0950.rec
LAB: pau k u0104 r u0204 sh i0304 m u0404 pau
REC: pau                 ts u0102 g u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0951.lab vs ../HVITE-M/Rec-mau-M-24/mfG0951.rec
LAB: pau h a0103 n e0203 r u0303 pau
REC: pau h a0103 g e0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0952.lab vs ../HVITE-M/Rec-mau-M-24/mfG0952.rec
LAB: pau k       i0103 k a0203 i0303 pau
REC: pau k o0104 u0204 g a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0953.lab vs ../HVITE-M/Rec-mau-M-24/mfG0953.rec
LAB: pau t o0103 b a0203 s  u0303 pau
REC: pau g o0103 g a0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0954.lab vs ../HVITE-M/Rec-mau-M-24/mfG0954.rec
LAB: pau   a0103 N0203   i0303 pau
REC: pau k a0103 N0203 b u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0955.lab vs ../HVITE-M/Rec-mau-M-24/mfG0955.rec
LAB: pau h o0102 k a0202 pau
REC: pau m a0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0956.lab vs ../HVITE-M/Rec-mau-M-24/mfG0956.rec
LAB: pau   u0104 r u0204 o0304 i0404 pau
REC: pau k o0102 r             o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0957.lab vs ../HVITE-M/Rec-mau-M-24/mfG0957.rec
LAB: pau k     a0104 i0204 f u0304 k u0404 pau
REC: pau u0104 r     a0204 g i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0959.lab vs ../HVITE-M/Rec-mau-M-24/mfG0959.rec
LAB: pau k i0102 N0202 pau
REC: pau d a0102 N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0963.lab vs ../HVITE-M/Rec-mau-M-24/mfG0963.rec
LAB: pau f u0104 k e0204 i0304 k i0404 pau

REC: pau         b e0103 N0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0964.lab vs ../HVITE-M/Rec-mau-M-24/mfG0964.rec
LAB: pau k o0104 n     i0204 m o0304 ts u0404 pau
REC: pau   b     a0104 N0204 g o0304    u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0965.lab vs ../HVITE-M/Rec-mau-M-24/mfG0965.rec
LAB: pau           k     a0102 d     o0202 pau
REC: pau g a0104 s o0204 r     i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0966.lab vs ../HVITE-M/Rec-mau-M-24/mfG0966.rec
LAB: pau k o0104 u0204 m u0304 r u0404 pau
REC: pau   g     o0102 m         u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0967.lab vs ../HVITE-M/Rec-mau-M-24/mfG0967.rec
LAB: pau r a0104 N0204 z a0304 ts u0404 pau
REC: pau k a0104 N0204 k a0304 k  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0968.lab vs ../HVITE-M/Rec-mau-M-24/mfG0968.rec
LAB: pau o0104 o0204 k     i0304 i0404 pau
REC: pau r     o0103 N0203 r     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0969.lab vs ../HVITE-M/Rec-mau-M-24/mfG0969.rec
LAB: pau   o0104 b o0204 e0304 r u0404 pau
REC: pau k o0104 g o0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0970.lab vs ../HVITE-M/Rec-mau-M-24/mfG0970.rec
LAB: pau k o0104 u0204 h a0304 N0404 pau
REC: pau   g     o0103 h a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0971.lab vs ../HVITE-M/Rec-mau-M-24/mfG0971.rec
LAB: pau h a0103 d a0203 k a0303 pau
REC: pau         b a0102 k a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0972.lab vs ../HVITE-M/Rec-mau-M-24/mfG0972.rec
LAB: pau   n     o0102 d     o0202 pau
REC: pau t a0103 b     u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0973.lab vs ../HVITE-M/Rec-mau-M-24/mfG0973.rec
LAB: pau   u0104 ch i0204 k o0304 m u0404 pau
REC: pau t o0104 b  i0204 k o0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0976.lab vs ../HVITE-M/Rec-mau-M-24/mfG0976.rec
LAB: pau ry u0104 u0204 g a0304 k u0404 pau
REC: pau h  o0104 u0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0977.lab vs ../HVITE-M/Rec-mau-M-24/mfG0977.rec
LAB: pau       t a0103 r e0203 r u0303 pau
REC: pau u0104 k a0204 b e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0978.lab vs ../HVITE-M/Rec-mau-M-24/mfG0978.rec
LAB: pau k u0103 d a0203 s u0303 pau

REC: pau k u0103 d a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0979.lab vs ../HVITE-M/Rec-mau-M-24/mfG0979.rec
LAB: pau sh i0103 b o0203   u0303 pau
REC: pau m  i0103 g o0203 t o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0980.lab vs ../HVITE-M/Rec-mau-M-24/mfG0980.rec
LAB: pau d o0104 u0204 j     o0304 u0404 pau
REC: pau   k     o0103 u0203 h     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0982.lab vs ../HVITE-M/Rec-mau-M-24/mfG0982.rec
LAB: pau ky u0104 u0204 k u0304 ts u0404 pau
REC: pau y  u0104 u0204   u0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0984.lab vs ../HVITE-M/Rec-mau-M-24/mfG0984.rec
LAB: pau f u0103 t o0203 N0303 pau
REC: pau   u0103 d o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0985.lab vs ../HVITE-M/Rec-mau-M-24/mfG0985.rec
LAB: pau   i0104 ch i0204 g a0304 ts u0404 pau
REC: pau h a0104 ch i0204 g a0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0986.lab vs ../HVITE-M/Rec-mau-M-24/mfG0986.rec
LAB: pau ky o0104 u0204 ry o0304 k u0404 pau
REC: pau sh u0104 u0204 ry o0304   u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0987.lab vs ../HVITE-M/Rec-mau-M-24/mfG0987.rec
LAB: pau b u0103 ts u0203 r i0303 pau
REC: pau m u0103 s  u0203 m e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0989.lab vs ../HVITE-M/Rec-mau-M-24/mfG0989.rec
LAB: pau   t     o0103 u0203 sh    o0303 pau
REC: pau k o0104 u0204 ky    o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0990.lab vs ../HVITE-M/Rec-mau-M-24/mfG0990.rec
LAB: pau k e0104 N0204 p o0304 u0404 pau
REC: pau k e0104 N0204 t o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0991.lab vs ../HVITE-M/Rec-mau-M-24/mfG0991.rec
LAB: pau n a0104 g a0204 r e0304 r u0404 pau
REC: pau t a0104 k a0204 m e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0992.lab vs ../HVITE-M/Rec-mau-M-24/mfG0992.rec
LAB: pau                 o0103 s e0203 j i0303 pau
REC: pau u0105 r i0205 k i0305 r e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0993.lab vs ../HVITE-M/Rec-mau-M-24/mfG0993.rec
LAB: pau                 t     o0102 u0202 pau
REC: pau k o0104 u0204 g a0304 k     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0994.lab vs ../HVITE-M/Rec-mau-M-24/mfG0994.rec
LAB: pau   t     a0103 r e0203 r u0303 pau

REC: pau s a0104 e0204 g i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0995.lab vs ../HVITE-M/Rec-mau-M-24/mfG0995.rec
LAB: pau               r a0103 q0203 p     a0303 pau
REC: pau g a0104 N0204 r             a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0996.lab vs ../HVITE-M/Rec-mau-M-24/mfG0996.rec
LAB: pau d e0103   a0203   u0303 pau
REC: pau n e0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG0999.lab vs ../HVITE-M/Rec-mau-M-24/mfG0999.rec
LAB: pau         i0102 n u0202 pau
REC: pau k a0103 N0203 b u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1000.lab vs ../HVITE-M/Rec-mau-M-24/mfG1000.rec
LAB: pau         u0103 ts u0203 r     u0303 pau
REC: pau k o0104 u0204 ts       u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1001.lab vs ../HVITE-M/Rec-mau-M-24/mfG1001.rec
LAB: pau k e0104 q0204 k a0304 N0404 pau
REC: pau b i0104 N0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1002.lab vs ../HVITE-M/Rec-mau-M-24/mfG1002.rec
LAB: pau b a0103 k u0203 r o0303 pau
REC: pau k a0103 r u0203   i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1004.lab vs ../HVITE-M/Rec-mau-M-24/mfG1004.rec
LAB: pau t e0103 i0203 j       i0303 pau
REC: pau g e0103       j u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1005.lab vs ../HVITE-M/Rec-mau-M-24/mfG1005.rec
LAB: pau t e0103 i0203 r     e0303 pau
REC: pau k e0104 i0204 e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1008.lab vs ../HVITE-M/Rec-mau-M-24/mfG1008.rec
LAB: pau   t     o0102 k u0202 pau
REC: pau d o0103 u0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1009.lab vs ../HVITE-M/Rec-mau-M-24/mfG1009.rec
LAB: pau   o0102 t o0202 pau
REC: pau t o0102 m o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1010.lab vs ../HVITE-M/Rec-mau-M-24/mfG1010.rec
LAB: pau h o0103 k e0203 ts u0303 pau
REC: pau h a0103 g e0203 m  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1012.lab vs ../HVITE-M/Rec-mau-M-24/mfG1012.rec
LAB: pau k a0103 r i0203 r u0303 pau
REC: pau b a0103   i0203   u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1013.lab vs ../HVITE-M/Rec-mau-M-24/mfG1013.rec
LAB: pau sh         i0102 j i0202 pau

REC: pau sh i0103 z u0203 k a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1014.lab vs ../HVITE-M/Rec-mau-M-24/mfG1014.rec
LAB: pau d e0103 k i0203 r u0303 pau
REC: pau g e0103 h i0203   N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1015.lab vs ../HVITE-M/Rec-mau-M-24/mfG1015.rec
LAB: pau   a0103 k u0203 b i0303 pau
REC: pau k a0103 r i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1017.lab vs ../HVITE-M/Rec-mau-M-24/mfG1017.rec
LAB: pau k o0103 o0203 r u0303 pau
REC: pau   m     o0102 r u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1018.lab vs ../HVITE-M/Rec-mau-M-24/mfG1018.rec
LAB: pau g e0104 N0204 k i0304 N0404 pau
REC: pau g e0104 N0204   i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1022.lab vs ../HVITE-M/Rec-mau-M-24/mfG1022.rec
LAB: pau t o0104 r i0204 ts u0304 g u0404 pau
REC: pau s u0104 s u0204 m  e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1023.lab vs ../HVITE-M/Rec-mau-M-24/mfG1023.rec
LAB: pau f u0103 s e0203 r u0303 pau
REC: pau f u0103 k e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1025.lab vs ../HVITE-M/Rec-mau-M-24/mfG1025.rec
LAB: pau s e0104 q0204 t o0304 k u0404 pau
REC: pau s       u0103 s u0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1026.lab vs ../HVITE-M/Rec-mau-M-24/mfG1026.rec
LAB: pau g             u0103 a0203 i0303 pau
REC: pau g a0104 i0204 j     i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1027.lab vs ../HVITE-M/Rec-mau-M-24/mfG1027.rec
LAB: pau s  u0103 m a0203   i0303 pau
REC: pau ts u0103 n a0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1028.lab vs ../HVITE-M/Rec-mau-M-24/mfG1028.rec
LAB: pau n i0103 k u0203 m u0303 pau
REC: pau         g o0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1029.lab vs ../HVITE-M/Rec-mau-M-24/mfG1029.rec
LAB: pau t a0103 h a0203 t a0303 pau
REC: pau d a0103 m a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1031.lab vs ../HVITE-M/Rec-mau-M-24/mfG1031.rec
LAB: pau b a0103 k u0203 r o0303 pau
REC: pau   a0103 r u0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1032.lab vs ../HVITE-M/Rec-mau-M-24/mfG1032.rec
LAB: pau   o0104 N0204 s e0304 N0404 pau

REC: pau g u0104 u0204 z e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1033.lab vs ../HVITE-M/Rec-mau-M-24/mfG1033.rec
LAB: pau         s     u0102 sh    i0202 pau
REC: pau d a0104 i0204 k     i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1034.lab vs ../HVITE-M/Rec-mau-M-24/mfG1034.rec
LAB: pau h a0104 r a0204 h a0304 r a0404 pau
REC: pau k a0104 w a0204 k a0304 s u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1035.lab vs ../HVITE-M/Rec-mau-M-24/mfG1035.rec
LAB: pau ky o0104 u0204 t e0304 i0404 pau
REC: pau g  o0104 u0204 k e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1036.lab vs ../HVITE-M/Rec-mau-M-24/mfG1036.rec
LAB: pau h a0102 r e0202 pau
REC: pau m a0102 r e0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1038.lab vs ../HVITE-M/Rec-mau-M-24/mfG1038.rec
LAB: pau   i0104   i0204 w a0304 k e0404 pau
REC: pau h i0104 k i0204 w a0304 k e0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1039.lab vs ../HVITE-M/Rec-mau-M-24/mfG1039.rec
LAB: pau f u0103 s e0203   i0303 pau
REC: pau f u0103 s e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1040.lab vs ../HVITE-M/Rec-mau-M-24/mfG1040.rec
LAB: pau sh i0103       j o0203 u0303 pau
REC: pau k  i0103 N0203 j       o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1041.lab vs ../HVITE-M/Rec-mau-M-24/mfG1041.rec
LAB: pau                 k i0102 j i0202 pau
REC: pau h i0103 r a0203 k         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1042.lab vs ../HVITE-M/Rec-mau-M-24/mfG1042.rec
LAB: pau    k     i0103 o0203 k u0303 pau
REC: pau sh u0103 ry    o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1043.lab vs ../HVITE-M/Rec-mau-M-24/mfG1043.rec
LAB: pau         k a0103 n o0203 u0303 pau
REC: pau f u0104 k a0204 n o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1044.lab vs ../HVITE-M/Rec-mau-M-24/mfG1044.rec
LAB: pau s e0103 b     i0203 r o0303 pau
REC: pau   k     i0103 i0203 r o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1045.lab vs ../HVITE-M/Rec-mau-M-24/mfG1045.rec
LAB: pau k o0104 u0204 u0304 N0404 pau
REC: pau k o0103 m     o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1047.lab vs ../HVITE-M/Rec-mau-M-24/mfG1047.rec
LAB: pau           a0104 k a0204 r u0304 i0404 pau

REC: pau t e0104 g a0204 k a0304 r       i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1049.lab vs ../HVITE-M/Rec-mau-M-24/mfG1049.rec
LAB: pau                h i0102 y u0202 pau
REC: pau ch u0104 u0204 d o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1050.lab vs ../HVITE-M/Rec-mau-M-24/mfG1050.rec
LAB: pau s a0104 z u0204 k a0304 r u0404 pau
REC: pau   a0104 z u0204 k a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1051.lab vs ../HVITE-M/Rec-mau-M-24/mfG1051.rec
LAB: pau sh i0103 r e0203 r u0303 pau
REC: pau z  u0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1053.lab vs ../HVITE-M/Rec-mau-M-24/mfG1053.rec
LAB: pau         h o0102 sh i0202 pau
REC: pau t o0103 j i0203 r  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1054.lab vs ../HVITE-M/Rec-mau-M-24/mfG1054.rec
LAB: pau   o0103 e0203 r u0303 pau
REC: pau k o0103 e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1055.lab vs ../HVITE-M/Rec-mau-M-24/mfG1055.rec
LAB: pau    u0103 k e0203 r u0303 pau
REC: pau sh i0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1056.lab vs ../HVITE-M/Rec-mau-M-24/mfG1056.rec
LAB: pau r o0104 k u0204 g a0304 ts u0404 pau
REC: pau   u0104 z u0204 m a0304 r  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1059.lab vs ../HVITE-M/Rec-mau-M-24/mfG1059.rec
LAB: pau         a0104 k a0204 ch a0304 N0404 pau
REC: pau a0105 t a0205 r a0305 sh i0405 i0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1060.lab vs ../HVITE-M/Rec-mau-M-24/mfG1060.rec
LAB: pau   ts    u0102 k     u0202 pau
REC: pau g e0103 j     u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1061.lab vs ../HVITE-M/Rec-mau-M-24/mfG1061.rec
LAB: pau   m     u0103 ry o0203 u0303 pau
REC: pau k o0103 u0203 ry       o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1062.lab vs ../HVITE-M/Rec-mau-M-24/mfG1062.rec
LAB: pau t o0105 r i0205 k a0305 e0405 r u0505 pau
REC: pau n o0105 r i0205 k a0305 e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1063.lab vs ../HVITE-M/Rec-mau-M-24/mfG1063.rec
LAB: pau h a0104 N0204 t a0304 i0404 pau
REC: pau g a0104 N0204 r a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1064.lab vs ../HVITE-M/Rec-mau-M-24/mfG1064.rec
LAB: pau f u0103 sh o0203 u0303 pau

REC: pau b u0103 sh o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1065.lab vs ../HVITE-M/Rec-mau-M-24/mfG1065.rec
LAB: pau m u0105 s e0205 k i0305 n     i0405 N0505 pau
REC: pau                 g e0104 N0204 i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1067.lab vs ../HVITE-M/Rec-mau-M-24/mfG1067.rec
LAB: pau    i0103 ch o0203 u0303 pau
REC: pau ch i0103 j  o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1068.lab vs ../HVITE-M/Rec-mau-M-24/mfG1068.rec
LAB: pau               h e0102 N0202 pau
REC: pau k e0104 N0204 t e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1069.lab vs ../HVITE-M/Rec-mau-M-24/mfG1069.rec
LAB: pau g u0103 q0203 t     o0303 pau
REC: pau h o0103 d     o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1071.lab vs ../HVITE-M/Rec-mau-M-24/mfG1071.rec
LAB: pau n         o0102 d o0202 pau
REC: pau n o0103 z o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1072.lab vs ../HVITE-M/Rec-mau-M-24/mfG1072.rec
LAB: pau sh o0104 u0204 r     e0304 i0404 pau
REC: pau ts u0103 t     o0203 m     e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1073.lab vs ../HVITE-M/Rec-mau-M-24/mfG1073.rec
LAB: pau sh i0103 z  u0203 k a0303 pau
REC: pau h  a0103 ts u0203 k a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1074.lab vs ../HVITE-M/Rec-mau-M-24/mfG1074.rec
LAB: pau sh i0104 r i0204 z o0304 k u0404 pau
REC: pau k  i0103 r         o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1075.lab vs ../HVITE-M/Rec-mau-M-24/mfG1075.rec
LAB: pau k u0106 s u0206 g u0306 q0406 t a0506 i0606 pau
REC: pau                 g u0104 N0204 t a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1076.lab vs ../HVITE-M/Rec-mau-M-24/mfG1076.rec
LAB: pau         m u0102 g i0202 pau
REC: pau n a0102 m         i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1077.lab vs ../HVITE-M/Rec-mau-M-24/mfG1077.rec
LAB: pau sh i0104 ts u0204 g e0304 N0404 pau
REC: pau          k  i0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1078.lab vs ../HVITE-M/Rec-mau-M-24/mfG1078.rec
LAB: pau         sh i0102 ch i0202 pau
REC: pau h i0103 n  i0203 ch i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1080.lab vs ../HVITE-M/Rec-mau-M-24/mfG1080.rec
LAB: pau sh i0103 m e0203 r u0303 pau

REC: pau k  i0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1081.lab vs ../HVITE-M/Rec-mau-M-24/mfG1081.rec
LAB: pau i0104 r e0204 m o0304 n o0404 pau
REC: pau       k e0103 m o0203 n o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1082.lab vs ../HVITE-M/Rec-mau-M-24/mfG1082.rec
LAB: pau h a0103 sh i0203 g o0303 pau
REC: pau m a0103    i0203 g o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1083.lab vs ../HVITE-M/Rec-mau-M-24/mfG1083.rec
LAB: pau s a0103 r a0203 n                  i0303 pau
REC: pau         k a0104 n a0204 sh i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1084.lab vs ../HVITE-M/Rec-mau-M-24/mfG1084.rec
LAB: pau         m     i0102 s     o0202 pau
REC: pau k o0104 u0204 z     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1086.lab vs ../HVITE-M/Rec-mau-M-24/mfG1086.rec
LAB: pau   t     o0102 k u0202 pau
REC: pau s o0103 u0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1087.lab vs ../HVITE-M/Rec-mau-M-24/mfG1087.rec
LAB: pau u0104 r     u0204   o0304 i0404 pau
REC: pau g     o0104 u0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1088.lab vs ../HVITE-M/Rec-mau-M-24/mfG1088.rec
LAB: pau           i0103 t o0203 k o0303 pau
REC: pau k i0104 n o0204 d o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1089.lab vs ../HVITE-M/Rec-mau-M-24/mfG1089.rec
LAB: pau       sh a0103 b e0203 r u0303 pau
REC: pau u0104 k  a0204 b e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1090.lab vs ../HVITE-M/Rec-mau-M-24/mfG1090.rec
LAB: pau    u0103   ch    u0203 u0303 pau
REC: pau ts u0103 k e0203 r     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1091.lab vs ../HVITE-M/Rec-mau-M-24/mfG1091.rec
LAB: pau y  a0103 k u0203 sh    a0303 pau
REC: pau ch i0103 k       a0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1092.lab vs ../HVITE-M/Rec-mau-M-24/mfG1092.rec
LAB: pau   b     u0103 N0203 r     i0303 pau
REC: pau h o0104 N0204 n     i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1093.lab vs ../HVITE-M/Rec-mau-M-24/mfG1093.rec
LAB: pau h i0103 r         a0203 k u0303 pau
REC: pau h i0104 r a0204 m e0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1094.lab vs ../HVITE-M/Rec-mau-M-24/mfG1094.rec
LAB: pau k e0104 q0204 sh       o0304 u0404 pau

REC: pau   m     u0103 sh i0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1095.lab vs ../HVITE-M/Rec-mau-M-24/mfG1095.rec
LAB: pau sh i0104 ts u0204 r e0304 N0404 pau
REC: pau          m  i0103 r e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1096.lab vs ../HVITE-M/Rec-mau-M-24/mfG1096.rec
LAB: pau h  a0103 g a0203 s u0303 pau
REC: pau ch i0103 g a0203   u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1097.lab vs ../HVITE-M/Rec-mau-M-24/mfG1097.rec
LAB: pau   o0103   u0203 b o0303 pau
REC: pau m o0103 t o0203 d e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1098.lab vs ../HVITE-M/Rec-mau-M-24/mfG1098.rec
LAB: pau   a0104 t a0204 e0304 r u0404 pau
REC: pau k a0104 m a0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1099.lab vs ../HVITE-M/Rec-mau-M-24/mfG1099.rec
LAB: pau h a0104 N0204 k     o0304 u0404 pau
REC: pau   t     a0103 N0203 b     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1100.lab vs ../HVITE-M/Rec-mau-M-24/mfG1100.rec
LAB: pau m u0103 sh i0203 r         o0303 pau
REC: pau f u0104 k  u0204 r a0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1101.lab vs ../HVITE-M/Rec-mau-M-24/mfG1101.rec
LAB: pau         s o0102 r  a0202 pau
REC: pau k u0103 g a0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1102.lab vs ../HVITE-M/Rec-mau-M-24/mfG1102.rec
LAB: pau k o0104 u0204 u0304 N0404 pau
REC: pau         m     o0102 N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1103.lab vs ../HVITE-M/Rec-mau-M-24/mfG1103.rec
LAB: pau h a0104 j i0204 m a0304 r u0404 pau
REC: pau m a0104 j i0204 w a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1104.lab vs ../HVITE-M/Rec-mau-M-24/mfG1104.rec
LAB: pau z o0102 k                   u0202 pau
REC: pau         k o0104 u0204 u0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1106.lab vs ../HVITE-M/Rec-mau-M-24/mfG1106.rec
LAB: pau s o0103 u0203 g i0303 pau
REC: pau k o0103 u0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1107.lab vs ../HVITE-M/Rec-mau-M-24/mfG1107.rec
LAB: pau                    t     o0102 o0202 pau
REC: pau k o0104 z u0204 ts u0304 m     i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1108.lab vs ../HVITE-M/Rec-mau-M-24/mfG1108.rec
LAB: pau    sh    i0104 i0204 r e0304 r u0404 pau

REC: pau ts u0104 k     i0204 d e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1109.lab vs ../HVITE-M/Rec-mau-M-24/mfG1109.rec
LAB: pau                 i0102 n u0202 pau
REC: pau k i0104 N0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1110.lab vs ../HVITE-M/Rec-mau-M-24/mfG1110.rec
LAB: pau   a0104 i0204 j o0304 u0404 pau
REC: pau k a0104 N0204 j o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1111.lab vs ../HVITE-M/Rec-mau-M-24/mfG1111.rec
LAB: pau s o0104 u0204 z o0304 u0404 pau
REC: pau k o0104 u0204 z o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1112.lab vs ../HVITE-M/Rec-mau-M-24/mfG1112.rec
LAB: pau j i0105 r e0205 q0305 t a0405 i0505 pau
REC: pau         g e0104 N0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1113.lab vs ../HVITE-M/Rec-mau-M-24/mfG1113.rec
LAB: pau   g     i0103 N0203 m          i0303 pau
REC: pau t a0104 ch    i0204 m a0304 ch i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1117.lab vs ../HVITE-M/Rec-mau-M-24/mfG1117.rec
LAB: pau s o0104 n a0204   e0304 r u0404 pau
REC: pau   u0104 m a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1118.lab vs ../HVITE-M/Rec-mau-M-24/mfG1118.rec
LAB: pau         a0103 m a0203 i0303 pau
REC: pau b a0104 i0204 k i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1119.lab vs ../HVITE-M/Rec-mau-M-24/mfG1119.rec
LAB: pau k a0103 k e0203 r u0303 pau
REC: pau s a0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1120.lab vs ../HVITE-M/Rec-mau-M-24/mfG1120.rec
LAB: pau sh i0104 N0204 g a0304 i0404 pau
REC: pau    r     i0103 g a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1121.lab vs ../HVITE-M/Rec-mau-M-24/mfG1121.rec
LAB: pau   s     o0102 r a0202 pau
REC: pau k o0103 u0203 w a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1123.lab vs ../HVITE-M/Rec-mau-M-24/mfG1123.rec
LAB: pau a0104 k i0204 r a0304 k a0404 pau
REC: pau a0104 y a0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1125.lab vs ../HVITE-M/Rec-mau-M-24/mfG1125.rec
LAB: pau h a0103 n e0203 r u0303 pau
REC: pau m a0103 n e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1126.lab vs ../HVITE-M/Rec-mau-M-24/mfG1126.rec
LAB: pau   d     a0103 N0203 sh    i0303 pau

REC: pau g a0104 i0204 j     i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1127.lab vs ../HVITE-M/Rec-mau-M-24/mfG1127.rec
LAB: pau e0104 i0204 gy    o0304 u0404 pau
REC: pau       e0103 N0203 ry    o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1129.lab vs ../HVITE-M/Rec-mau-M-24/mfG1129.rec
LAB: pau o0105 s o0205 r o0305 sh i0405 i0505 pau
REC: pau       t a0104 n o0204 sh i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1130.lab vs ../HVITE-M/Rec-mau-M-24/mfG1130.rec
LAB: pau h         a0102 r  e0202 pau
REC: pau h a0103 y a0203 sh i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1131.lab vs ../HVITE-M/Rec-mau-M-24/mfG1131.rec
LAB: pau sh o0103 k u0203 h i0303 pau
REC: pau          k i0102 k i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1132.lab vs ../HVITE-M/Rec-mau-M-24/mfG1132.rec
LAB: pau i0103 j i0203 r u0303 pau
REC: pau       h i0102 r u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1133.lab vs ../HVITE-M/Rec-mau-M-24/mfG1133.rec
LAB: pau t o0102 k                     u0202 pau
REC: pau         k o0104 u0204 n i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1134.lab vs ../HVITE-M/Rec-mau-M-24/mfG1134.rec
LAB: pau            w     a0102 N0202 pau
REC: pau sh i0103 b a0203 r     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1136.lab vs ../HVITE-M/Rec-mau-M-24/mfG1136.rec
LAB: pau g o0104 u0204 s e0304 i0404 pau
REC: pau g o0104 u0204 k e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1137.lab vs ../HVITE-M/Rec-mau-M-24/mfG1137.rec
LAB: pau k         a0102 m          u0202 pau
REC: pau k a0104 n e0204 m o0304 ch i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1138.lab vs ../HVITE-M/Rec-mau-M-24/mfG1138.rec
LAB: pau   o0104 s a0204 e0304 r u0404 pau
REC: pau m u0104 k a0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1139.lab vs ../HVITE-M/Rec-mau-M-24/mfG1139.rec
LAB: pau   o0104 k  u0204 r e0304 r u0404 pau
REC: pau m o0104 ts u0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1142.lab vs ../HVITE-M/Rec-mau-M-24/mfG1142.rec
LAB: pau sh u0102 sh i0202 pau
REC: pau ch i0102 ch i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1144.lab vs ../HVITE-M/Rec-mau-M-24/mfG1144.rec
LAB: pau         j i0103 q0203 t o0303 pau

REC: pau s u0104 t a0204 a0304 t o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1145.lab vs ../HVITE-M/Rec-mau-M-24/mfG1145.rec
LAB: pau t a0104 i0204 m e0304 N0404 pau
REC: pau s a0104 i0204 n a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1146.lab vs ../HVITE-M/Rec-mau-M-24/mfG1146.rec
LAB: pau   i0104 y a0204 sh i0304 i0404 pau
REC: pau h a0103 y a0203 sh       i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1147.lab vs ../HVITE-M/Rec-mau-M-24/mfG1147.rec
LAB: pau j o0103 u0203 k     i0303 pau
REC: pau k o0103 s     e0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1148.lab vs ../HVITE-M/Rec-mau-M-24/mfG1148.rec
LAB: pau d o0104 k u0204 r i0304 ts u0404 pau
REC: pau h o0104   u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1149.lab vs ../HVITE-M/Rec-mau-M-24/mfG1149.rec
LAB: pau ch u0104 u0204 d o0304 k u0404 pau
REC: pau    k     i0103 r o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1150.lab vs ../HVITE-M/Rec-mau-M-24/mfG1150.rec
LAB: pau   d     e0105 k  i0205 a0305 g a0405 r u0505 pau
REC: pau m o0106 u0206 sh i0306 a0406 g e0506 r u0606 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1151.lab vs ../HVITE-M/Rec-mau-M-24/mfG1151.rec
LAB: pau i0103 g                     e0203 N0303 pau
REC: pau       g e0104 N0204 m i0304 ts    u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1153.lab vs ../HVITE-M/Rec-mau-M-24/mfG1153.rec
LAB: pau y o0104 u0204 s u0304 r u0404 pau
REC: pau k o0103 o0203 s         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1155.lab vs ../HVITE-M/Rec-mau-M-24/mfG1155.rec
LAB: pau y  o0103 u0203 t o0303 pau
REC: pau ry o0103 u0203 d o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1156.lab vs ../HVITE-M/Rec-mau-M-24/mfG1156.rec
LAB: pau   k     o0101 pau
REC: pau m o0102 N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1157.lab vs ../HVITE-M/Rec-mau-M-24/mfG1157.rec
LAB: pau k a0103 k     o0203 u0303 pau
REC: pau   m     a0102 g     o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1158.lab vs ../HVITE-M/Rec-mau-M-24/mfG1158.rec
LAB: pau   d     a0103 sh i0203 N0303 pau
REC: pau b a0104 q0204 k  i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1159.lab vs ../HVITE-M/Rec-mau-M-24/mfG1159.rec
LAB: pau y o0103 r o0203 N0303 pau

REC: pau         b o0102 N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1160.lab vs ../HVITE-M/Rec-mau-M-24/mfG1160.rec
LAB: pau         h     i0102 f     u0202 pau
REC: pau r e0104 i0204 t     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1161.lab vs ../HVITE-M/Rec-mau-M-24/mfG1161.rec
LAB: pau                          ch a0101 pau
REC: pau ts u0104 r a0204 n u0304 k  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1163.lab vs ../HVITE-M/Rec-mau-M-24/mfG1163.rec
LAB: pau i0104 n     e0204 m u0304 r     i0404 pau
REC: pau r     e0104 N0204 m       e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1167.lab vs ../HVITE-M/Rec-mau-M-24/mfG1167.rec
LAB: pau u0104 r     u0204 o0304 i0404 pau
REC: pau m     o0103 y     o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1168.lab vs ../HVITE-M/Rec-mau-M-24/mfG1168.rec
LAB: pau         h o0102 r i0202 pau
REC: pau k o0103 r i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1169.lab vs ../HVITE-M/Rec-mau-M-24/mfG1169.rec
LAB: pau s a0105 sh    i0205 a0305 g e0405 r u0505 pau
REC: pau   k     a0104 k     a0204 g e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1171.lab vs ../HVITE-M/Rec-mau-M-24/mfG1171.rec
LAB: pau h o0104 u0204 m o0304 N0404 pau
REC: pau   k     o0103 m o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1173.lab vs ../HVITE-M/Rec-mau-M-24/mfG1173.rec
LAB: pau         s o0102 ch i0202 pau
REC: pau f u0103 k u0203 sh i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1174.lab vs ../HVITE-M/Rec-mau-M-24/mfG1174.rec
LAB: pau k a0104 N0204 ch    o0304 u0404 pau
REC: pau   t     a0103 N0203 b     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1175.lab vs ../HVITE-M/Rec-mau-M-24/mfG1175.rec
LAB: pau   a0103 r a0203   u0303 pau
REC: pau k a0103 g a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1177.lab vs ../HVITE-M/Rec-mau-M-24/mfG1177.rec
LAB: pau k i0103 b o0203 u0303 pau
REC: pau k i0103 h o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1178.lab vs ../HVITE-M/Rec-mau-M-24/mfG1178.rec
LAB: pau d e0104 k o0204 b o0304 k o0404 pau
REC: pau         t e0103 g o0203 r o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1179.lab vs ../HVITE-M/Rec-mau-M-24/mfG1179.rec
LAB: pau s a0104       k u0204 h i0304 N0404 pau

REC: pau s a0104 i0204 k         i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1180.lab vs ../HVITE-M/Rec-mau-M-24/mfG1180.rec
LAB: pau k a0104 N0204 k a0304 k u0404 pau
REC: pau n a0104 i0204 k a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1181.lab vs ../HVITE-M/Rec-mau-M-24/mfG1181.rec
LAB: pau h i0105 k         i0205 a0305 g e0405 r u0505 pau
REC: pau         k u0105 r i0205 a0305 g e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1182.lab vs ../HVITE-M/Rec-mau-M-24/mfG1182.rec
LAB: pau sh u0104 u0204 d a0304 N0404 pau
REC: pau ch u0104 u0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1183.lab vs ../HVITE-M/Rec-mau-M-24/mfG1183.rec
LAB: pau            u0102 t a0202 pau
REC: pau m a0103 ts u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1186.lab vs ../HVITE-M/Rec-mau-M-24/mfG1186.rec
LAB: pau u0105 s u0205 g u0305 r a0405 i0505 pau
REC: pau         i0103 g         a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1187.lab vs ../HVITE-M/Rec-mau-M-24/mfG1187.rec
LAB: pau j i0104 N0204 k a0304 k u0404 pau
REC: pau   m     i0103 g a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1188.lab vs ../HVITE-M/Rec-mau-M-24/mfG1188.rec
LAB: pau           a0102 t o0202 pau
REC: pau s a0103 b o0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1189.lab vs ../HVITE-M/Rec-mau-M-24/mfG1189.rec
LAB: pau k  i0104 t a0204 e0304 r u0404 pau
REC: pau ts u0104 t a0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1190.lab vs ../HVITE-M/Rec-mau-M-24/mfG1190.rec
LAB: pau h i0103 h a0203 N0303 pau
REC: pau g o0103 h a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1191.lab vs ../HVITE-M/Rec-mau-M-24/mfG1191.rec
LAB: pau               n e0101 pau
REC: pau s u0103 e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1192.lab vs ../HVITE-M/Rec-mau-M-24/mfG1192.rec
LAB: pau k o0104 u0204 n a0304 i0404 pau
REC: pau k o0104 u0204 g a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1193.lab vs ../HVITE-M/Rec-mau-M-24/mfG1193.rec
LAB: pau k o0104 u0204 s e0304 i0404 pau
REC: pau g o0104 u0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1194.lab vs ../HVITE-M/Rec-mau-M-24/mfG1194.rec
LAB: pau f u0102 k                          u0202 pau

REC: pau         k u0104 r u0204 sh i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1195.lab vs ../HVITE-M/Rec-mau-M-24/mfG1195.rec
LAB: pau       s a0103 s a0203 i0303 pau
REC: pau u0104 t a0204 g a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1196.lab vs ../HVITE-M/Rec-mau-M-24/mfG1196.rec
LAB: pau k a0103 s u0203 m u0303 pau
REC: pau s a0103 s u0203 g a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1199.lab vs ../HVITE-M/Rec-mau-M-24/mfG1199.rec
LAB: pau s a0104 N0204 my a0304 k u0404 pau
REC: pau   k     a0103 g  a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1200.lab vs ../HVITE-M/Rec-mau-M-24/mfG1200.rec
LAB: pau s e0103 i0203   i0303 pau
REC: pau t e0103 i0203 j i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1201.lab vs ../HVITE-M/Rec-mau-M-24/mfG1201.rec
LAB: pau               i0103 N0203 k u0303 pau
REC: pau r i0104 N0204 k     a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1202.lab vs ../HVITE-M/Rec-mau-M-24/mfG1202.rec
LAB: pau k o0104 u0204 sh i0304 k i0404 pau
REC: pau   m     o0102 sh         i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1203.lab vs ../HVITE-M/Rec-mau-M-24/mfG1203.rec
LAB: pau k o0104 u0204 s e0304 N0404 pau
REC: pau   g     o0103 z e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1204.lab vs ../HVITE-M/Rec-mau-M-24/mfG1204.rec
LAB: pau n a0104 d a0204 m e0304 r u0404 pau
REC: pau n a0104 g a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1205.lab vs ../HVITE-M/Rec-mau-M-24/mfG1205.rec
LAB: pau t e0104 q0204 p  o0304 u0404 pau
REC: pau k e0104 q0204 sh o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1206.lab vs ../HVITE-M/Rec-mau-M-24/mfG1206.rec
LAB: pau h o0102 r               i0202 pau
REC: pau k u0104 r i0204 i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1207.lab vs ../HVITE-M/Rec-mau-M-24/mfG1207.rec
LAB: pau y a0103 q0203 ts u0303 pau
REC: pau y       a0102 b  u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1208.lab vs ../HVITE-M/Rec-mau-M-24/mfG1208.rec
LAB: pau   o0102 i0202 pau
REC: pau t o0102 i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1209.lab vs ../HVITE-M/Rec-mau-M-24/mfG1209.rec
LAB: pau sh u0104 u0204 ky    o0304 u0404 pau

REC: pau f  u0103 k     u0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1211.lab vs ../HVITE-M/Rec-mau-M-24/mfG1211.rec
LAB: pau   h     o0102 sh i0202 pau
REC: pau k u0103 i0203 k  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1213.lab vs ../HVITE-M/Rec-mau-M-24/mfG1213.rec
LAB: pau         gy    a0102 k     u0202 pau
REC: pau b a0104 N0204 g     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1215.lab vs ../HVITE-M/Rec-mau-M-24/mfG1215.rec
LAB: pau sh u0104 k       u0204 d a0304 i0404 pau
REC: pau          k o0104 u0204 g a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1216.lab vs ../HVITE-M/Rec-mau-M-24/mfG1216.rec
LAB: pau n  o0104 u0204 r i0304 ts    u0404 pau
REC: pau ch o0104 u0204   h     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1217.lab vs ../HVITE-M/Rec-mau-M-24/mfG1217.rec
LAB: pau   s     e0104 k i0204 t a0304 N0404 pau
REC: pau j u0104 N0204 k         a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1218.lab vs ../HVITE-M/Rec-mau-M-24/mfG1218.rec
LAB: pau         k o0104 u0204 ts u0304 u0404 pau
REC: pau r o0103 k       o0203 ts       u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1219.lab vs ../HVITE-M/Rec-mau-M-24/mfG1219.rec
LAB: pau sh i0104 N0204 j i0304 ts u0404 pau
REC: pau    k     i0103 j i0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1220.lab vs ../HVITE-M/Rec-mau-M-24/mfG1220.rec
LAB: pau           u0102 m e0202 pau
REC: pau h i0103 g e0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1222.lab vs ../HVITE-M/Rec-mau-M-24/mfG1222.rec
LAB: pau         sh    u0102 t     o0202 pau
REC: pau k o0104 q0204 k     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1223.lab vs ../HVITE-M/Rec-mau-M-24/mfG1223.rec
LAB: pau n i0104 N0204 t a0304 i0404 pau
REC: pau b i0104 N0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1224.lab vs ../HVITE-M/Rec-mau-M-24/mfG1224.rec
LAB: pau g i0102 sh       i0202 pau
REC: pau j i0103 sh i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1226.lab vs ../HVITE-M/Rec-mau-M-24/mfG1226.rec
LAB: pau s o0104 u0204 j  u0304 u0404 pau
REC: pau   k     o0103 ky u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1227.lab vs ../HVITE-M/Rec-mau-M-24/mfG1227.rec
LAB: pau h i0104 k i0204 w a0304 k e0404 pau

REC: pau                 g a0102 k e0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1228.lab vs ../HVITE-M/Rec-mau-M-24/mfG1228.rec
LAB: pau   j     o0103 ch u0203 u0303 pau
REC: pau k u0104 u0204 ch u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1229.lab vs ../HVITE-M/Rec-mau-M-24/mfG1229.rec
LAB: pau d     e0103 N0203 sh       i0303 pau
REC: pau u0104 r     e0204 sh i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1230.lab vs ../HVITE-M/Rec-mau-M-24/mfG1230.rec
LAB: pau s a0104 N0204 k o0304 u0404 pau
REC: pau h a0104 N0204 k o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1231.lab vs ../HVITE-M/Rec-mau-M-24/mfG1231.rec
LAB: pau k       o0103 u0203 g     i0303 pau
REC: pau k o0104 u0204 n     i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1234.lab vs ../HVITE-M/Rec-mau-M-24/mfG1234.rec
LAB: pau                j       u0102 u0202 pau
REC: pau ts u0104 u0204 j i0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1235.lab vs ../HVITE-M/Rec-mau-M-24/mfG1235.rec
LAB: pau                  m       e0101 pau
REC: pau sh u0104 k u0204 m e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1236.lab vs ../HVITE-M/Rec-mau-M-24/mfG1236.rec
LAB: pau       k u0104 u0204 ch    u0304 u0404 pau
REC: pau u0103 k             e0203 r     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1237.lab vs ../HVITE-M/Rec-mau-M-24/mfG1237.rec
LAB: pau         j i0103 j i0203 ts u0303 pau
REC: pau k a0104 k u0204 j i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1238.lab vs ../HVITE-M/Rec-mau-M-24/mfG1238.rec
LAB: pau t o0103 k e0203   i0303 pau
REC: pau s o0103 r e0203 d e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1241.lab vs ../HVITE-M/Rec-mau-M-24/mfG1241.rec
LAB: pau n a0103 b     i0203 k     u0303 pau
REC: pau   k     a0104 i0204 i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1242.lab vs ../HVITE-M/Rec-mau-M-24/mfG1242.rec
LAB: pau sh i0103 k u0203 m i0303 pau
REC: pau          k u0102 m i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1243.lab vs ../HVITE-M/Rec-mau-M-24/mfG1243.rec
LAB: pau t a0104 k u0204 sh i0304 i0404 pau
REC: pau g a0103 k u0203 sh       i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1244.lab vs ../HVITE-M/Rec-mau-M-24/mfG1244.rec
LAB: pau s u0103 m a0203 s u0303 pau

REC: pau p u0103 r a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1245.lab vs ../HVITE-M/Rec-mau-M-24/mfG1245.rec
LAB: pau t       a0103 i0203 d     o0303 pau
REC: pau t a0104 i0204 b     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1246.lab vs ../HVITE-M/Rec-mau-M-24/mfG1246.rec
LAB: pau n a0103 g         e0203 k u0303 pau
REC: pau m a0104 g i0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1247.lab vs ../HVITE-M/Rec-mau-M-24/mfG1247.rec
LAB: pau t       e0103 ch o0203 u0303 pau
REC: pau t a0104 N0204 j  o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1248.lab vs ../HVITE-M/Rec-mau-M-24/mfG1248.rec
LAB: pau ts         u0102 n o0202 pau
REC: pau ts u0103 m o0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1249.lab vs ../HVITE-M/Rec-mau-M-24/mfG1249.rec
LAB: pau t a0103 i0203 h o0303 pau
REC: pau t a0103 i0203 k o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1250.lab vs ../HVITE-M/Rec-mau-M-24/mfG1250.rec
LAB: pau   ts    u0103 i0203 n     i0303 pau
REC: pau k a0104 i0204 g     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1251.lab vs ../HVITE-M/Rec-mau-M-24/mfG1251.rec
LAB: pau          s a0103 N0203 k     a0303 pau
REC: pau sh i0104 t a0204 g     a0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1252.lab vs ../HVITE-M/Rec-mau-M-24/mfG1252.rec
LAB: pau   n     i0103 q0203 k i0303 pau
REC: pau f u0104 N0204 i0304 k i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1253.lab vs ../HVITE-M/Rec-mau-M-24/mfG1253.rec
LAB: pau ch a0104 k u0204 j i0304 ts u0404 pau
REC: pau          k         u0102 z  u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1254.lab vs ../HVITE-M/Rec-mau-M-24/mfG1254.rec
LAB: pau         b a0102 k a0202 pau
REC: pau s a0103 r a0203 n i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1255.lab vs ../HVITE-M/Rec-mau-M-24/mfG1255.rec
LAB: pau sh i0103 k o0203 m u0303 pau
REC: pau          g o0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1257.lab vs ../HVITE-M/Rec-mau-M-24/mfG1257.rec
LAB: pau   e0104 N0204 ch o0304 u0404 pau
REC: pau g a0104 i0204 t  o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1258.lab vs ../HVITE-M/Rec-mau-M-24/mfG1258.rec
LAB: pau         k     u0102 s     a0202 pau

REC: pau g a0104 N0204 j     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1259.lab vs ../HVITE-M/Rec-mau-M-24/mfG1259.rec
LAB: pau f u0105 s a0205 w a0305 sh i0405 i0505 pau
REC: pau                 k a0103 k  o0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1260.lab vs ../HVITE-M/Rec-mau-M-24/mfG1260.rec
LAB: pau           b     a0103 i0203 u0303 pau
REC: pau k u0104 r i0204 i0304 m     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1262.lab vs ../HVITE-M/Rec-mau-M-24/mfG1262.rec
LAB: pau j u0103 k         e0203 N0303 pau
REC: pau         k i0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1265.lab vs ../HVITE-M/Rec-mau-M-24/mfG1265.rec
LAB: pau b u0104 q0204 sh i0304 ts u0404 pau
REC: pau   k     i0103 j  i0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1266.lab vs ../HVITE-M/Rec-mau-M-24/mfG1266.rec
LAB: pau t e0104 n u0204 g             u0304 i0404 pau
REC: pau                 g e0103 N0203 r     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1267.lab vs ../HVITE-M/Rec-mau-M-24/mfG1267.rec
LAB: pau g a0104 q0204 k       o0304 u0404 pau
REC: pau   b     a0103 k u0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1269.lab vs ../HVITE-M/Rec-mau-M-24/mfG1269.rec
LAB: pau                 d e0103 a0203 u0303 pau
REC: pau m i0104 y a0204 b u0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1271.lab vs ../HVITE-M/Rec-mau-M-24/mfG1271.rec
LAB: pau j       i0103 sh i0203 N0303 pau
REC: pau j i0104 N0204 k  o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1273.lab vs ../HVITE-M/Rec-mau-M-24/mfG1273.rec
LAB: pau           t     a0103 o0203 s u0303 pau
REC: pau s a0104 b i0204 ts    u0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1274.lab vs ../HVITE-M/Rec-mau-M-24/mfG1274.rec
LAB: pau h a0103 g e0203 m u0303 pau
REC: pau s a0103 g u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1276.lab vs ../HVITE-M/Rec-mau-M-24/mfG1276.rec
LAB: pau ky o0104 u0204 k a0304 i0404 pau
REC: pau b  o0104 u0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1277.lab vs ../HVITE-M/Rec-mau-M-24/mfG1277.rec
LAB: pau n a0104 i0204 k a0304 k u0404 pau
REC: pau z a0104 i0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1278.lab vs ../HVITE-M/Rec-mau-M-24/mfG1278.rec
LAB: pau g e0104 N0204 k i0304 N0404 pau

REC: pau g e0104 N0204   i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1279.lab vs ../HVITE-M/Rec-mau-M-24/mfG1279.rec
LAB: pau s o0103 m u0203 k u0303 pau
REC: pau g o0102 m         u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1280.lab vs ../HVITE-M/Rec-mau-M-24/mfG1280.rec
LAB: pau   o0103 d o0203 r       i0303 pau
REC: pau g a0104 s o0204 r i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1282.lab vs ../HVITE-M/Rec-mau-M-24/mfG1282.rec
LAB: pau   i0104 y a0204 sh i0304 i0404 pau
REC: pau h a0103 y a0203 sh       i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1284.lab vs ../HVITE-M/Rec-mau-M-24/mfG1284.rec
LAB: pau          d o0102 r o0202 pau
REC: pau sh i0103 g o0203 t o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1285.lab vs ../HVITE-M/Rec-mau-M-24/mfG1285.rec
LAB: pau   d     e0103 N0203 sh    i0303 pau
REC: pau m e0104 i0204 j     i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1286.lab vs ../HVITE-M/Rec-mau-M-24/mfG1286.rec
LAB: pau f u0103 u0203 f     u0303 pau
REC: pau m u0103 g     e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1287.lab vs ../HVITE-M/Rec-mau-M-24/mfG1287.rec
LAB: pau f u0104 N0204 i0304 k i0404 pau
REC: pau   b     u0103 N0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1288.lab vs ../HVITE-M/Rec-mau-M-24/mfG1288.rec
LAB: pau          k u0102 z         u0202 pau
REC: pau ts u0104 k u0204 z u0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1290.lab vs ../HVITE-M/Rec-mau-M-24/mfG1290.rec
LAB: pau         a0103 i0203 z u0303 pau
REC: pau t a0104 i0204 i0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1291.lab vs ../HVITE-M/Rec-mau-M-24/mfG1291.rec
LAB: pau j i0104 ts    u0204 z a0304 i0404 pau
REC: pau   k     o0104 u0204 k a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1292.lab vs ../HVITE-M/Rec-mau-M-24/mfG1292.rec
LAB: pau   m     a0102 n     e0202 pau
REC: pau k a0103 g     e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1293.lab vs ../HVITE-M/Rec-mau-M-24/mfG1293.rec
LAB: pau m a0104 z u0204 sh i0304 i0404 pau
REC: pau g a0103 k u0203 sh       i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1295.lab vs ../HVITE-M/Rec-mau-M-24/mfG1295.rec
LAB: pau sh i0103 j o0203 u0303 pau

REC: pau ch i0103 j o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1296.lab vs ../HVITE-M/Rec-mau-M-24/mfG1296.rec
LAB: pau   a0104 ts u0204 ry o0304 k u0404 pau
REC: pau g a0104 k  u0204 ry o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1297.lab vs ../HVITE-M/Rec-mau-M-24/mfG1297.rec
LAB: pau ts u0104 q0204 p a0304 r u0404 pau
REC: pau    k     u0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1298.lab vs ../HVITE-M/Rec-mau-M-24/mfG1298.rec
LAB: pau a0104 k i0204 r a0304 k a0404 pau
REC: pau       k a0103 r a0203 d a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1299.lab vs ../HVITE-M/Rec-mau-M-24/mfG1299.rec
LAB: pau    o0103 b i0203 r u0303 pau
REC: pau sh o0103 m i0203   N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1302.lab vs ../HVITE-M/Rec-mau-M-24/mfG1302.rec
LAB: pau         d a0103 sh i0203 N0303 pau
REC: pau k a0104 k u0204 sh i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1303.lab vs ../HVITE-M/Rec-mau-M-24/mfG1303.rec
LAB: pau n e0104 N0204 ry o0304 u0404 pau
REC: pau   e0103 N0203 ry       o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1304.lab vs ../HVITE-M/Rec-mau-M-24/mfG1304.rec
LAB: pau u0104 n a0204 z u0304 k u0404 pau
REC: pau       m a0103 z a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1305.lab vs ../HVITE-M/Rec-mau-M-24/mfG1305.rec
LAB: pau       k     i0103 o0203 k u0303 pau
REC: pau i0104 N0204 ry    o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1306.lab vs ../HVITE-M/Rec-mau-M-24/mfG1306.rec
LAB: pau sh u0104 q0204 p a0304 ts u0404 pau
REC: pau    h     i0103 t a0203 s  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1307.lab vs ../HVITE-M/Rec-mau-M-24/mfG1307.rec
LAB: pau d     a0103 k     a0203 r     a0303 pau
REC: pau u0104 r     a0204 n     a0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1308.lab vs ../HVITE-M/Rec-mau-M-24/mfG1308.rec
LAB: pau ch i0103 j o0203   u0303 pau
REC: pau b  u0103 j o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1310.lab vs ../HVITE-M/Rec-mau-M-24/mfG1310.rec
LAB: pau y o0103 n a0203 k a0303 pau
REC: pau t e0103 g a0203 r a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1311.lab vs ../HVITE-M/Rec-mau-M-24/mfG1311.rec
LAB: pau k a0103 N0203 g       o0303 pau

REC: pau k a0103       g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1312.lab vs ../HVITE-M/Rec-mau-M-24/mfG1312.rec
LAB: pau h i0104 q0204 k         o0304 sh i0404 pau
REC: pau   d     o0104 k u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1314.lab vs ../HVITE-M/Rec-mau-M-24/mfG1314.rec
LAB: pau k u0102 m u0202 pau
REC: pau k u0102 m o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1315.lab vs ../HVITE-M/Rec-mau-M-24/mfG1315.rec
LAB: pau         m     e0101 pau
REC: pau i0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1316.lab vs ../HVITE-M/Rec-mau-M-24/mfG1316.rec
LAB: pau g i0103 j         u0203 ts u0303 pau
REC: pau         j u0103 k u0203 s  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1317.lab vs ../HVITE-M/Rec-mau-M-24/mfG1317.rec
LAB: pau ts u0103 m e0203 r u0303 pau
REC: pau k  i0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1319.lab vs ../HVITE-M/Rec-mau-M-24/mfG1319.rec
LAB: pau t o0104 b i0204 ts u0304 k u0404 pau
REC: pau         r i0103 z  u0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1321.lab vs ../HVITE-M/Rec-mau-M-24/mfG1321.rec
LAB: pau h a0105 z u0205 k a0305 sh i0405 i0505 pau
REC: pau         b u0103 r a0203 sh       i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1322.lab vs ../HVITE-M/Rec-mau-M-24/mfG1322.rec
LAB: pau k     i0104 t a0204 e0304 r u0404 pau
REC: pau u0105 q0205 t a0305 e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1323.lab vs ../HVITE-M/Rec-mau-M-24/mfG1323.rec
LAB: pau         u0103 e0203 k i0303 pau
REC: pau h i0103 g     e0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1325.lab vs ../HVITE-M/Rec-mau-M-24/mfG1325.rec
LAB: pau h o0103 t e0203 r u0303 pau
REC: pau k o0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1326.lab vs ../HVITE-M/Rec-mau-M-24/mfG1326.rec
LAB: pau k a0102 t         a0202 pau
REC: pau k a0103 t a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1327.lab vs ../HVITE-M/Rec-mau-M-24/mfG1327.rec
LAB: pau k e0104 i0204 k     o0304 u0404 pau
REC: pau t e0103 g     o0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1328.lab vs ../HVITE-M/Rec-mau-M-24/mfG1328.rec
LAB: pau r i0103 g a0203   i0303 pau

REC: pau s u0103 g a0203 t a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1330.lab vs ../HVITE-M/Rec-mau-M-24/mfG1330.rec
LAB: pau    ch    o0103 ch i0203 k     u0303 pau
REC: pau sh u0104 u0204 ch       u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1331.lab vs ../HVITE-M/Rec-mau-M-24/mfG1331.rec
LAB: pau t o0103 k     o0203 r     o0303 pau
REC: pau   ch    u0104 u0204 o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1333.lab vs ../HVITE-M/Rec-mau-M-24/mfG1333.rec
LAB: pau k a0103 m o0203 k u0303 pau
REC: pau k a0103 k o0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1334.lab vs ../HVITE-M/Rec-mau-M-24/mfG1334.rec
LAB: pau y o0104 r     o0204 sh i0304 i0404 pau
REC: pau   k     o0103 u0203 sh       i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1335.lab vs ../HVITE-M/Rec-mau-M-24/mfG1335.rec
LAB: pau k o0104   u0204 g a0304 k u0404 pau
REC: pau g o0104 k u0204 r a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1336.lab vs ../HVITE-M/Rec-mau-M-24/mfG1336.rec
LAB: pau   e0103 N0203 k i0303 pau
REC: pau t e0103 N0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1338.lab vs ../HVITE-M/Rec-mau-M-24/mfG1338.rec
LAB: pau o0104 ch i0204 ts u0304 k i0404 pau
REC: pau          u0103 g  o0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1339.lab vs ../HVITE-M/Rec-mau-M-24/mfG1339.rec
LAB: pau           k     o0102 b     u0202 pau
REC: pau m a0104 z u0204 sh    i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1343.lab vs ../HVITE-M/Rec-mau-M-24/mfG1343.rec
LAB: pau             k i0101 pau
REC: pau u0103 e0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1344.lab vs ../HVITE-M/Rec-mau-M-24/mfG1344.rec
LAB: pau r i0103 k o0203 u0303 pau
REC: pau j i0102 k       o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1345.lab vs ../HVITE-M/Rec-mau-M-24/mfG1345.rec
LAB: pau k a0104 N0204 t e0304   N0404 pau
REC: pau t a0104 o0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1349.lab vs ../HVITE-M/Rec-mau-M-24/mfG1349.rec
LAB: pau s e0104 N0204 k o0304 k u0404 pau
REC: pau t a0104 N0204 d o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1351.lab vs ../HVITE-M/Rec-mau-M-24/mfG1351.rec
LAB: pau k               u0102 m i0202 pau

REC: pau k u0104 r i0204 i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1353.lab vs ../HVITE-M/Rec-mau-M-24/mfG1353.rec
LAB: pau k a0104 k e0204 a0304 sh       i0404 pau
REC: pau   a0104   y     a0204 sh i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1355.lab vs ../HVITE-M/Rec-mau-M-24/mfG1355.rec
LAB: pau k o0103 g e0203 r u0303 pau
REC: pau t o0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1356.lab vs ../HVITE-M/Rec-mau-M-24/mfG1356.rec
LAB: pau j i0105 r e0205 q0305 t a0405 i0505 pau
REC: pau j         i0104 q0204 t a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1357.lab vs ../HVITE-M/Rec-mau-M-24/mfG1357.rec
LAB: pau t a0104 k u0204 sh i0304 i0404 pau
REC: pau s a0102 k                i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1358.lab vs ../HVITE-M/Rec-mau-M-24/mfG1358.rec
LAB: pau               a0102 s       a0202 pau
REC: pau o0105 b a0205 a0305 s a0405 N0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1359.lab vs ../HVITE-M/Rec-mau-M-24/mfG1359.rec
LAB: pau h         i0103 n i0203 ch i0303 pau
REC: pau h a0104 r i0204 k i0304 r  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1360.lab vs ../HVITE-M/Rec-mau-M-24/mfG1360.rec
LAB: pau s o0103 u0203 ch i0303 pau
REC: pau s       o0102 ch i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1361.lab vs ../HVITE-M/Rec-mau-M-24/mfG1361.rec
LAB: pau         sh    u0102 m     i0202 pau
REC: pau k e0104 N0204 t     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1362.lab vs ../HVITE-M/Rec-mau-M-24/mfG1362.rec
LAB: pau a0102 k               u0202 pau
REC: pau       k a0103 r u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1364.lab vs ../HVITE-M/Rec-mau-M-24/mfG1364.rec
LAB: pau ch i0103 k o0203 k u0303 pau
REC: pau k  i0103 r o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1365.lab vs ../HVITE-M/Rec-mau-M-24/mfG1365.rec
LAB: pau s a0104 q0204 s u0304 r u0404 pau
REC: pau   k     a0102 s         u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1366.lab vs ../HVITE-M/Rec-mau-M-24/mfG1366.rec
LAB: pau k e0104 ts u0204 r o0304 N0404 pau
REC: pau   e0104 p  u0204 r o0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1368.lab vs ../HVITE-M/Rec-mau-M-24/mfG1368.rec
LAB: pau k e0104 i0204 t a0304 i0404 pau

REC: pau   g     e0103 h i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1370.lab vs ../HVITE-M/Rec-mau-M-24/mfG1370.rec
LAB: pau h a0104 i0204 k e0304 i0404 pau
REC: pau b a0104 i0204 t e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1371.lab vs ../HVITE-M/Rec-mau-M-24/mfG1371.rec
LAB: pau               g a0103 q0203 k i0303 pau
REC: pau k a0104 N0204 g       e0304 k i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1372.lab vs ../HVITE-M/Rec-mau-M-24/mfG1372.rec
LAB: pau ch       i0102 r          i0202 pau
REC: pau ch u0104 u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1375.lab vs ../HVITE-M/Rec-mau-M-24/mfG1375.rec
LAB: pau s a0104 N0204 t a0304 N0404 pau
REC: pau h a0104 N0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1376.lab vs ../HVITE-M/Rec-mau-M-24/mfG1376.rec
LAB: pau t a0104 i0204 k e0304 ts u0404 pau
REC: pau s a0104 i0204 k e0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1378.lab vs ../HVITE-M/Rec-mau-M-24/mfG1378.rec
LAB: pau k u0104 u0204 k     a0304 N0404 pau
REC: pau   u0103 r     a0203 m     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1379.lab vs ../HVITE-M/Rec-mau-M-24/mfG1379.rec
LAB: pau j i0104 ts u0204 z a0304 i0404 pau
REC: pau         sh o0103 z a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1380.lab vs ../HVITE-M/Rec-mau-M-24/mfG1380.rec
LAB: pau g o0103 u0203 k     a0303 pau
REC: pau m o0103 r     a0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1381.lab vs ../HVITE-M/Rec-mau-M-24/mfG1381.rec
LAB: pau d       o0103 ry o0203 k u0303 pau
REC: pau d o0104 u0204 ry o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1382.lab vs ../HVITE-M/Rec-mau-M-24/mfG1382.rec
LAB: pau t         a0102 t e0202 pau
REC: pau t a0103 k u0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1383.lab vs ../HVITE-M/Rec-mau-M-24/mfG1383.rec
LAB: pau         m i0102 r u0202 pau
REC: pau k o0102 m         u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1384.lab vs ../HVITE-M/Rec-mau-M-24/mfG1384.rec
LAB: pau ky    u0104 u0204 k u0304 ts    u0404 pau
REC: pau i0104 k     u0204   s     e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1385.lab vs ../HVITE-M/Rec-mau-M-24/mfG1385.rec
LAB: pau sh u0103 ry o0203 k u0303 pau

REC: pau k  i0103 ry o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1386.lab vs ../HVITE-M/Rec-mau-M-24/mfG1386.rec
LAB: pau            i0102 k u0202 pau
REC: pau sh i0103 z u0203 k a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1387.lab vs ../HVITE-M/Rec-mau-M-24/mfG1387.rec
LAB: pau         s u0103 w a0203 r u0303 pau
REC: pau k o0104 r o0204 g a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1388.lab vs ../HVITE-M/Rec-mau-M-24/mfG1388.rec
LAB: pau m  i0102 r         u0202 pau
REC: pau ts u0103 r u0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1389.lab vs ../HVITE-M/Rec-mau-M-24/mfG1389.rec
LAB: pau         s     o0102 u0202 pau
REC: pau u0103 g o0203 k     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1390.lab vs ../HVITE-M/Rec-mau-M-24/mfG1390.rec
LAB: pau y o0103 u0203 ch i0303 pau
REC: pau h o0103 u0203 sh i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1392.lab vs ../HVITE-M/Rec-mau-M-24/mfG1392.rec
LAB: pau ts u0102 m i0202 pau
REC: pau ts u0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1393.lab vs ../HVITE-M/Rec-mau-M-24/mfG1393.rec
LAB: pau m i0103 b       u0203 r i0303 pau
REC: pau         b u0103 N0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1394.lab vs ../HVITE-M/Rec-mau-M-24/mfG1394.rec
LAB: pau   i0104 q0204 sh u0304 N0404 pau
REC: pau b a0104 i0204 k  i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1395.lab vs ../HVITE-M/Rec-mau-M-24/mfG1395.rec
LAB: pau y a0103 t     a0203 r     a0303 pau
REC: pau k a0104 N0204 b     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1396.lab vs ../HVITE-M/Rec-mau-M-24/mfG1396.rec
LAB: pau f u0103 z o0203 k u0303 pau
REC: pau         m u0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1397.lab vs ../HVITE-M/Rec-mau-M-24/mfG1397.rec
LAB: pau           n     o0103 N0203 k i0303 pau
REC: pau h i0104 r o0204 g     a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1398.lab vs ../HVITE-M/Rec-mau-M-24/mfG1398.rec
LAB: pau k a0103 N0203 b u0303 pau
REC: pau d a0103 N0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1400.lab vs ../HVITE-M/Rec-mau-M-24/mfG1400.rec
LAB: pau         ky u0102 u0202 pau

REC: pau k u0103 r  u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1401.lab vs ../HVITE-M/Rec-mau-M-24/mfG1401.rec
LAB: pau k o0104 u0204 u0304 N0404 pau
REC: pau         m     o0102 N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1402.lab vs ../HVITE-M/Rec-mau-M-24/mfG1402.rec
LAB: pau    z     e0103 N0203 g o0303 pau
REC: pau ts u0103 m     o0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1405.lab vs ../HVITE-M/Rec-mau-M-24/mfG1405.rec
LAB: pau k o0104 u0204 gy    o0304 u0404 pau
REC: pau d o0103 ry    o0203 k     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1407.lab vs ../HVITE-M/Rec-mau-M-24/mfG1407.rec
LAB: pau sh o0104 u0204 s e0304 ts       u0404 pau
REC: pau k  o0104 u0204         ts u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1409.lab vs ../HVITE-M/Rec-mau-M-24/mfG1409.rec
LAB: pau k a0104 k     u0204 m     e0304 i0404 pau
REC: pau   h     a0103 z     u0203 r     e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1410.lab vs ../HVITE-M/Rec-mau-M-24/mfG1410.rec
LAB: pau   a0103 m a0203 g u0303 pau
REC: pau k a0103 m a0203   u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1411.lab vs ../HVITE-M/Rec-mau-M-24/mfG1411.rec
LAB: pau sh i0104 ts u0204 g e0304 N0404 pau
REC: pau          k  i0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1412.lab vs ../HVITE-M/Rec-mau-M-24/mfG1412.rec
LAB: pau k u0103 g a0203 ts u0303 pau
REC: pau n u0103 r a0203 s  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1413.lab vs ../HVITE-M/Rec-mau-M-24/mfG1413.rec
LAB: pau t o0102 ch                      i0202 pau
REC: pau         ch u0104 u0204 sh i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1414.lab vs ../HVITE-M/Rec-mau-M-24/mfG1414.rec
LAB: pau h               o0102 r u0202 pau
REC: pau h o0104 t o0204 N0304 d o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1415.lab vs ../HVITE-M/Rec-mau-M-24/mfG1415.rec
LAB: pau ch o0104 u0204 d a0304 i0404 pau
REC: pau t  o0104 u0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1416.lab vs ../HVITE-M/Rec-mau-M-24/mfG1416.rec
LAB: pau         k i0103 r o0203 k u0303 pau
REC: pau d o0103 k         u0203 j i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1417.lab vs ../HVITE-M/Rec-mau-M-24/mfG1417.rec
LAB: pau sh i0104 ts u0204 r e0304 N0404 pau

REC: pau             i0103 z e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1419.lab vs ../HVITE-M/Rec-mau-M-24/mfG1419.rec
LAB: pau   u0103 w a0203 s a0303 pau
REC: pau h a0103 s a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1420.lab vs ../HVITE-M/Rec-mau-M-24/mfG1420.rec
LAB: pau t o0104 u0204 s     o0304 u0404 pau
REC: pau   d     o0103 u0203 z     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1421.lab vs ../HVITE-M/Rec-mau-M-24/mfG1421.rec
LAB: pau               k e0102 ch    i0202 pau
REC: pau t e0104 q0204 k       i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1423.lab vs ../HVITE-M/Rec-mau-M-24/mfG1423.rec
LAB: pau               k a0102 sh i0202 pau
REC: pau g a0103 q0203 k          i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1424.lab vs ../HVITE-M/Rec-mau-M-24/mfG1424.rec
LAB: pau t o0106 r     i0206 a0306 ts    u0406 k     a0506 u0606 pau
REC: pau   d     e0105 k     i0205 a0305 g     a0405 r     u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1426.lab vs ../HVITE-M/Rec-mau-M-24/mfG1426.rec
LAB: pau         w a0102 k         i0202 pau
REC: pau m a0104 k i0204 k o0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1428.lab vs ../HVITE-M/Rec-mau-M-24/mfG1428.rec
LAB: pau sh i0103 k o0203 u0303 pau
REC: pau k  i0103 h o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1430.lab vs ../HVITE-M/Rec-mau-M-24/mfG1430.rec
LAB: pau   f     u0103 k a0203 i0303 pau
REC: pau r a0104 i0204 n e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1431.lab vs ../HVITE-M/Rec-mau-M-24/mfG1431.rec
LAB: pau y a0103 t a0203 r a0303 pau
REC: pau         g a0102 r a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1432.lab vs ../HVITE-M/Rec-mau-M-24/mfG1432.rec
LAB: pau t o0104 ts    u0204 z e0304 N0404 pau
REC: pau   s     a0104 i0204 z e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1434.lab vs ../HVITE-M/Rec-mau-M-24/mfG1434.rec
LAB: pau j u0104   u0204 t e0304   N0404 pau
REC: pau k u0104 z u0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1435.lab vs ../HVITE-M/Rec-mau-M-24/mfG1435.rec
LAB: pau sh i0103 t e0203 N0303 pau
REC: pau k  i0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1436.lab vs ../HVITE-M/Rec-mau-M-24/mfG1436.rec
LAB: pau z a0104 ts u0204 y o0304 u0404 pau

REC: pau d a0104    i0204 y o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1437.lab vs ../HVITE-M/Rec-mau-M-24/mfG1437.rec
LAB: pau m i0103 q0203 ts       u0303 pau
REC: pau   k     u0103 ts u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1438.lab vs ../HVITE-M/Rec-mau-M-24/mfG1438.rec
LAB: pau d a0104   N0204 d a0304 N0404 pau
REC: pau b a0104 k u0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1439.lab vs ../HVITE-M/Rec-mau-M-24/mfG1439.rec
LAB: pau   e0102 s a0202 pau
REC: pau h i0102 z a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1440.lab vs ../HVITE-M/Rec-mau-M-24/mfG1440.rec
LAB: pau ky u0104 u0204 k u0304 ts u0404 pau
REC: pau                k u0102 z  u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1444.lab vs ../HVITE-M/Rec-mau-M-24/mfG1444.rec
LAB: pau         r e0103 i0203 g     i0303 pau
REC: pau k i0103 r             e0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1445.lab vs ../HVITE-M/Rec-mau-M-24/mfG1445.rec
LAB: pau k i0103 t a0203 i0303 pau
REC: pau m i0103 r a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1447.lab vs ../HVITE-M/Rec-mau-M-24/mfG1447.rec
LAB: pau           m     o0103 u0203 k e0303 pau
REC: pau s u0104 g u0204 r     e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1448.lab vs ../HVITE-M/Rec-mau-M-24/mfG1448.rec
LAB: pau   y     o0103 y a0203 k u0303 pau
REC: pau k o0104 u0204 y a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1449.lab vs ../HVITE-M/Rec-mau-M-24/mfG1449.rec
LAB: pau t a0103   q0203 t a0303 pau
REC: pau k a0103 g a0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1450.lab vs ../HVITE-M/Rec-mau-M-24/mfG1450.rec
LAB: pau m a0103 k u0203 r a0303 pau
REC: pau         g a0102 r a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1452.lab vs ../HVITE-M/Rec-mau-M-24/mfG1452.rec
LAB: pau m o0104 ch    i0204 r     o0304 N0404 pau
REC: pau   n     i0103 k     u0203 m     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1455.lab vs ../HVITE-M/Rec-mau-M-24/mfG1455.rec
LAB: pau   k     e0104 N0204 sh i0304 k i0404 pau
REC: pau t o0104 r     i0204 h  i0304 k i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1456.lab vs ../HVITE-M/Rec-mau-M-24/mfG1456.rec
LAB: pau         j a0102 m a0202 pau

REC: pau d a0103 r a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1458.lab vs ../HVITE-M/Rec-mau-M-24/mfG1458.rec
LAB: pau f u0104 t         a0204 t a0304 b     i0404 pau
REC: pau         t a0104 d a0204   y     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1459.lab vs ../HVITE-M/Rec-mau-M-24/mfG1459.rec
LAB: pau sh i0104 h a0204 r a0304 i0404 pau
REC: pau          k a0103 r a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1460.lab vs ../HVITE-M/Rec-mau-M-24/mfG1460.rec
LAB: pau         k o0103 g a0203 t a0303 pau
REC: pau b a0102 k                 a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1461.lab vs ../HVITE-M/Rec-mau-M-24/mfG1461.rec
LAB: pau   h     a0103 i0203 ry    o0303 pau
REC: pau s a0104 i0204 b     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1463.lab vs ../HVITE-M/Rec-mau-M-24/mfG1463.rec
LAB: pau k o0104 u0204 j     i0304 ts    u0404 pau
REC: pau         b     o0103 sh    u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1464.lab vs ../HVITE-M/Rec-mau-M-24/mfG1464.rec
LAB: pau t e0104 i0204 d e0304 N0404 pau
REC: pau s e0104 N0204 g e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1465.lab vs ../HVITE-M/Rec-mau-M-24/mfG1465.rec
LAB: pau z e0104 i0204 k i0304 N0404 pau
REC: pau h e0104 i0204 k i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1466.lab vs ../HVITE-M/Rec-mau-M-24/mfG1466.rec
LAB: pau m  u0103 y o0203 u0303 pau
REC: pau ts u0103 y o0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1467.lab vs ../HVITE-M/Rec-mau-M-24/mfG1467.rec
LAB: pau         s u0103 s u0203 g u0303 pau
REC: pau t e0104 k i0204 s u0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1468.lab vs ../HVITE-M/Rec-mau-M-24/mfG1468.rec
LAB: pau         j i0102 sh    o0202 pau
REC: pau k i0103 j       u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1469.lab vs ../HVITE-M/Rec-mau-M-24/mfG1469.rec
LAB: pau               g a0102 N0202 pau
REC: pau k o0104 N0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1470.lab vs ../HVITE-M/Rec-mau-M-24/mfG1470.rec
LAB: pau         s       o0103 e0203 r u0303 pau
REC: pau w a0104 s u0204 r     e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1471.lab vs ../HVITE-M/Rec-mau-M-24/mfG1471.rec
LAB: pau sh i0103 m e0203 s  u0303 pau

REC: pau k  u0103 b e0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1472.lab vs ../HVITE-M/Rec-mau-M-24/mfG1472.rec
LAB: pau m e0104 i0204 s a0304 k u0404 pau
REC: pau k e0104 q0204 k a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1476.lab vs ../HVITE-M/Rec-mau-M-24/mfG1476.rec
LAB: pau sh i0103 t a0203 k u0303 pau
REC: pau k  i0103 k a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1477.lab vs ../HVITE-M/Rec-mau-M-24/mfG1477.rec
LAB: pau t a0104 i0204 k a0304 i0404 pau
REC: pau d a0104 i0204 t a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1478.lab vs ../HVITE-M/Rec-mau-M-24/mfG1478.rec
LAB: pau h o0102 k               a0202 pau
REC: pau         k i0103 b a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1479.lab vs ../HVITE-M/Rec-mau-M-24/mfG1479.rec
LAB: pau r o0103 u0203 h       i0303 pau
REC: pau k o0104 o0204 h i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1480.lab vs ../HVITE-M/Rec-mau-M-24/mfG1480.rec
LAB: pau sh u0103 gy o0203   u0303 pau
REC: pau n  i0103 g  o0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1483.lab vs ../HVITE-M/Rec-mau-M-24/mfG1483.rec
LAB: pau           m     e0102 i0202 pau
REC: pau t e0103 g a0203 r     a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1485.lab vs ../HVITE-M/Rec-mau-M-24/mfG1485.rec
LAB: pau s e0104 i0204 sh i0304 ts u0404 pau
REC: pau h e0104 i0204 j  i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1486.lab vs ../HVITE-M/Rec-mau-M-24/mfG1486.rec
LAB: pau   o0102 k u0202 pau
REC: pau t o0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1487.lab vs ../HVITE-M/Rec-mau-M-24/mfG1487.rec
LAB: pau k i0104 N0204 ky u0304 u0404 pau
REC: pau g u0104 N0204 sh u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1488.lab vs ../HVITE-M/Rec-mau-M-24/mfG1488.rec
LAB: pau k e0104 ts u0204 d a0304 N0404 pau
REC: pau s e0104 ts u0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1489.lab vs ../HVITE-M/Rec-mau-M-24/mfG1489.rec
LAB: pau    u0103 r a0203 m u0303 pau
REC: pau ts u0103 m a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1490.lab vs ../HVITE-M/Rec-mau-M-24/mfG1490.rec
LAB: pau h i0104 t o0204 m a0304 z u0404 pau

REC: pau m u0103 s u0203 m         e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1491.lab vs ../HVITE-M/Rec-mau-M-24/mfG1491.rec
LAB: pau k i0102 r                 i0202 pau
REC: pau         r i0103 z u0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1493.lab vs ../HVITE-M/Rec-mau-M-24/mfG1493.rec
LAB: pau                 a0102 i0202 pau
REC: pau t a0104 i0204 b o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1494.lab vs ../HVITE-M/Rec-mau-M-24/mfG1494.rec
LAB: pau   ts    u0103 t o0203 m         e0303 pau
REC: pau h i0105 q0205 k o0305 m e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1496.lab vs ../HVITE-M/Rec-mau-M-24/mfG1496.rec
LAB: pau h o0104 d o0204 k o0304 s u0404 pau
REC: pau         s u0103 g o0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1497.lab vs ../HVITE-M/Rec-mau-M-24/mfG1497.rec
LAB: pau j       o0102 sh i0202 pau
REC: pau j o0103 u0203 k  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1498.lab vs ../HVITE-M/Rec-mau-M-24/mfG1498.rec
LAB: pau j       u0103 N0203 b     i0303 pau
REC: pau j u0104 N0204 n     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1499.lab vs ../HVITE-M/Rec-mau-M-24/mfG1499.rec
LAB: pau k o0103 u0203 z a0303 pau
REC: pau   g     o0102 z a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1500.lab vs ../HVITE-M/Rec-mau-M-24/mfG1500.rec
LAB: pau s e0104 q0204 k a0304 k u0404 pau
REC: pau   h     i0103 r a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1501.lab vs ../HVITE-M/Rec-mau-M-24/mfG1501.rec
LAB: pau k i0103 ts u0203   i0303 pau
REC: pau   u0103 ts u0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1502.lab vs ../HVITE-M/Rec-mau-M-24/mfG1502.rec
LAB: pau n e0103 m a0203 k         i0303 pau
REC: pau         g a0103 k u0203 h i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1503.lab vs ../HVITE-M/Rec-mau-M-24/mfG1503.rec
LAB: pau         k     e0102 N0202 pau
REC: pau g e0103 N0203 g     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1504.lab vs ../HVITE-M/Rec-mau-M-24/mfG1504.rec
LAB: pau j i0104 q0204 t a0304 i0404 pau
REC: pau b u0104 q0204 t a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1505.lab vs ../HVITE-M/Rec-mau-M-24/mfG1505.rec
LAB: pau h i0103 r a0203 k u0303 pau

REC: pau k u0103 d a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1507.lab vs ../HVITE-M/Rec-mau-M-24/mfG1507.rec
LAB: pau b a0104 N0204 g o0304 u0404 pau
REC: pau   m     a0102 g       o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1508.lab vs ../HVITE-M/Rec-mau-M-24/mfG1508.rec
LAB: pau               j i0102 sh o0202 pau
REC: pau k a0103 i0203 j          o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1509.lab vs ../HVITE-M/Rec-mau-M-24/mfG1509.rec
LAB: pau         w a0103 b i0203 r u0303 pau
REC: pau h a0104 r i0204 k i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1512.lab vs ../HVITE-M/Rec-mau-M-24/mfG1512.rec
LAB: pau ry a0104 k u0204 s u0304 r u0404 pau
REC: pau          y u0103 z u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1514.lab vs ../HVITE-M/Rec-mau-M-24/mfG1514.rec
LAB: pau k  i0103 g e0203 N0303 pau
REC: pau sh i0103 h e0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1516.lab vs ../HVITE-M/Rec-mau-M-24/mfG1516.rec
LAB: pau   d     e0104 p     a0204 a0304 t o0404 pau
REC: pau f u0104 r     u0204 s     a0304 t o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1518.lab vs ../HVITE-M/Rec-mau-M-24/mfG1518.rec
LAB: pau n e0104 k     u0204 t a0304 i0404 pau
REC: pau   h     o0104 u0204 t a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1520.lab vs ../HVITE-M/Rec-mau-M-24/mfG1520.rec
LAB: pau k e0104 i0204 h     i0304 N0404 pau
REC: pau g e0104       N0204 i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1521.lab vs ../HVITE-M/Rec-mau-M-24/mfG1521.rec
LAB: pau       m a0103 g u0203 r e0303 pau
REC: pau a0104 b a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1522.lab vs ../HVITE-M/Rec-mau-M-24/mfG1522.rec
LAB: pau g     o0103 u0203 k a0303 pau
REC: pau a0103 b     u0203 r a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1523.lab vs ../HVITE-M/Rec-mau-M-24/mfG1523.rec
LAB: pau s u0104 u0204 g a0304 k  u0404 pau
REC: pau s a0104 N0204 g a0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1524.lab vs ../HVITE-M/Rec-mau-M-24/mfG1524.rec
LAB: pau   k     a0102 g     u0202 pau
REC: pau t a0103 b     u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1525.lab vs ../HVITE-M/Rec-mau-M-24/mfG1525.rec
LAB: pau f u0104 k     i0204 d a0304 s u0404 pau

REC: pau         o0104 N0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1526.lab vs ../HVITE-M/Rec-mau-M-24/mfG1526.rec
LAB: pau         ts    u0102 m     a0202 pau
REC: pau g a0104 q0204 k     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1527.lab vs ../HVITE-M/Rec-mau-M-24/mfG1527.rec
LAB: pau s o0104 n         o0204 u0304 ch    i0404 pau
REC: pau k o0105 n o0205 m a0305 sh    i0405 i0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1528.lab vs ../HVITE-M/Rec-mau-M-24/mfG1528.rec
LAB: pau j i0103       sh i0203 N0303 pau
REC: pau j i0103 N0203 sh       u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1530.lab vs ../HVITE-M/Rec-mau-M-24/mfG1530.rec
LAB: pau k i0103 j i0203 ts       u0303 pau
REC: pau k i0103         ts u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1532.lab vs ../HVITE-M/Rec-mau-M-24/mfG1532.rec
LAB: pau t e0104 N0204 i0304 N0404 pau
REC: pau g e0104 N0204 i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1534.lab vs ../HVITE-M/Rec-mau-M-24/mfG1534.rec
LAB: pau ry u0104 u0204 g a0304 k u0404 pau
REC: pau    n     u0103 r a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1535.lab vs ../HVITE-M/Rec-mau-M-24/mfG1535.rec
LAB: pau g e0103 j u0203 N0303 pau
REC: pau g e0103 h i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1536.lab vs ../HVITE-M/Rec-mau-M-24/mfG1536.rec
LAB: pau z a0104 ts u0204 y o0304 u0404 pau
REC: pau k a0104 ts u0204 y o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1537.lab vs ../HVITE-M/Rec-mau-M-24/mfG1537.rec
LAB: pau h  i0104 N0204 p a0304 N0404 pau
REC: pau sh i0104 N0204 p a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1540.lab vs ../HVITE-M/Rec-mau-M-24/mfG1540.rec
LAB: pau y o0103 s o0203 k u0303 pau
REC: pau r i0103 s o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1541.lab vs ../HVITE-M/Rec-mau-M-24/mfG1541.rec
LAB: pau   t     o0103 m o0203 n i0303 pau
REC: pau h o0104 N0204 m o0304 n o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1542.lab vs ../HVITE-M/Rec-mau-M-24/mfG1542.rec
LAB: pau k  i0104 N0204 k o0304 u0404 pau
REC: pau sh i0104 N0204 k o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1543.lab vs ../HVITE-M/Rec-mau-M-24/mfG1543.rec
LAB: pau h i0103 r     a0203 k u0303 pau

REC: pau   g     a0103 i0203 b u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1544.lab vs ../HVITE-M/Rec-mau-M-24/mfG1544.rec
LAB: pau    h     i0103 k a0203 r     i0303 pau
REC: pau ch u0104 u0204 k       a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1545.lab vs ../HVITE-M/Rec-mau-M-24/mfG1545.rec
LAB: pau t e0103 N0203 k i0303 pau
REC: pau d e0103 N0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1546.lab vs ../HVITE-M/Rec-mau-M-24/mfG1546.rec
LAB: pau         s a0103 sh i0203 z     u0303 pau
REC: pau t a0104 d a0204 sh       i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1547.lab vs ../HVITE-M/Rec-mau-M-24/mfG1547.rec
LAB: pau   o0104 b o0204   e0304 r u0404 pau
REC: pau k o0104 b o0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1548.lab vs ../HVITE-M/Rec-mau-M-24/mfG1548.rec
LAB: pau y u0104 u0204 s e0304 N0404 pau
REC: pau y u0104 u0204 b i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1550.lab vs ../HVITE-M/Rec-mau-M-24/mfG1550.rec
LAB: pau s u0104 j i0204 m i0304 ch i0404 pau
REC: pau                 k i0102 sh i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1551.lab vs ../HVITE-M/Rec-mau-M-24/mfG1551.rec
LAB: pau n a0103   i0203 f u0303 pau
REC: pau k a0103 r i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1552.lab vs ../HVITE-M/Rec-mau-M-24/mfG1552.rec
LAB: pau   o0104 k u0204 r e0304 r u0404 pau
REC: pau s u0104 g u0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1553.lab vs ../HVITE-M/Rec-mau-M-24/mfG1553.rec
LAB: pau b a0102 k a0202 pau
REC: pau k a0102   N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1554.lab vs ../HVITE-M/Rec-mau-M-24/mfG1554.rec
LAB: pau         e0104 k i0204 t a0304 i0404 pau
REC: pau y u0104 u0204 k         a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1555.lab vs ../HVITE-M/Rec-mau-M-24/mfG1555.rec
LAB: pau h i0103 r a0203 k u0303 pau
REC: pau k u0103 d a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1556.lab vs ../HVITE-M/Rec-mau-M-24/mfG1556.rec
LAB: pau sh i0102 k         i0202 pau
REC: pau          k u0102 g i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1558.lab vs ../HVITE-M/Rec-mau-M-24/mfG1558.rec
LAB: pau k a0106 g a0206 y a0306 k a0406 sh i0506 i0606 pau

REC: pau   u0106 r a0206 y a0306 m a0406 sh i0506 i0606 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1559.lab vs ../HVITE-M/Rec-mau-M-24/mfG1559.rec
LAB: pau i0105 i0205 k a0305 g e0405 N0505 pau
REC: pau             k i0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1560.lab vs ../HVITE-M/Rec-mau-M-24/mfG1560.rec
LAB: pau y o0103         u0203 i0303 pau
REC: pau h o0103 k o0203 r     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1561.lab vs ../HVITE-M/Rec-mau-M-24/mfG1561.rec
LAB: pau   e0104 N0204 s o0304 u0404 pau
REC: pau k e0104 N0204 t o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1562.lab vs ../HVITE-M/Rec-mau-M-24/mfG1562.rec
LAB: pau y o0103         u0203 i0303 pau
REC: pau h o0103 k o0203 r     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1564.lab vs ../HVITE-M/Rec-mau-M-24/mfG1564.rec
LAB: pau t a0103 k a0203 s a0303 pau
REC: pau d a0103 m a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1565.lab vs ../HVITE-M/Rec-mau-M-24/mfG1565.rec
LAB: pau ch u0104 u0204   o0304 u0404 pau
REC: pau k  u0104 u0204 k o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1566.lab vs ../HVITE-M/Rec-mau-M-24/mfG1566.rec
LAB: pau t a0104 ch i0204 k         i0304 r u0404 pau
REC: pau                  k a0103 r i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1567.lab vs ../HVITE-M/Rec-mau-M-24/mfG1567.rec
LAB: pau h a0103 k     a0203 i0303 pau
REC: pau k a0103 i0203 g     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1568.lab vs ../HVITE-M/Rec-mau-M-24/mfG1568.rec
LAB: pau   j     i0103 k a0203 N0303 pau
REC: pau k u0104 u0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1569.lab vs ../HVITE-M/Rec-mau-M-24/mfG1569.rec
LAB: pau                s o0102 u0202 pau
REC: pau sh i0104 N0204 z o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1571.lab vs ../HVITE-M/Rec-mau-M-24/mfG1571.rec
LAB: pau s e0104 i0204 t o0304 u0404 pau
REC: pau h e0104 i0204 k o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1572.lab vs ../HVITE-M/Rec-mau-M-24/mfG1572.rec
LAB: pau sh o0103 k u0203 j i0303 pau
REC: pau d  o0103 k u0203 j i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1573.lab vs ../HVITE-M/Rec-mau-M-24/mfG1573.rec
LAB: pau ch a0104 k u0204 j i0304 ts u0404 pau

REC: pau k  a0104 k u0204 j i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1574.lab vs ../HVITE-M/Rec-mau-M-24/mfG1574.rec
LAB: pau t e0103 b     i0203 k     i0303 pau
REC: pau m e0104 N0204 d     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1578.lab vs ../HVITE-M/Rec-mau-M-24/mfG1578.rec
LAB: pau         t     a0103 i0203 h o0303 pau
REC: pau g a0104 i0204 k     o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1579.lab vs ../HVITE-M/Rec-mau-M-24/mfG1579.rec
LAB: pau   a0103 r a0203 s u0303 pau
REC: pau d a0103 r a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1580.lab vs ../HVITE-M/Rec-mau-M-24/mfG1580.rec
LAB: pau t a0104 q0204 s         u0304 r u0404 pau
REC: pau t a0104       s u0204 k a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1581.lab vs ../HVITE-M/Rec-mau-M-24/mfG1581.rec
LAB: pau s a0103 s     o0203 u0303 pau
REC: pau s       a0102 k     u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1582.lab vs ../HVITE-M/Rec-mau-M-24/mfG1582.rec
LAB: pau i0103 k                  a0203 g a0303 pau
REC: pau       k a0104 n a0204 sh i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1585.lab vs ../HVITE-M/Rec-mau-M-24/mfG1585.rec
LAB: pau g e0104 N0204 sh o0304 u0404 pau
REC: pau b e0104 N0204 sh o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1588.lab vs ../HVITE-M/Rec-mau-M-24/mfG1588.rec
LAB: pau sh u0104 u0204 t o0304 k u0404 pau
REC: pau k  u0104 u0204 k o0304   u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1590.lab vs ../HVITE-M/Rec-mau-M-24/mfG1590.rec
LAB: pau s o0104 r o0204 b a0304 N0404 pau
REC: pau k o0104   N0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1591.lab vs ../HVITE-M/Rec-mau-M-24/mfG1591.rec
LAB: pau y u0103 k i0203 k i0303 pau
REC: pau         b u0102 k i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1593.lab vs ../HVITE-M/Rec-mau-M-24/mfG1593.rec
LAB: pau n  i0102 j i0202 pau
REC: pau sh i0102 j i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1596.lab vs ../HVITE-M/Rec-mau-M-24/mfG1596.rec
LAB: pau sh i0103 r a0203 g a0303 pau
REC: pau ch i0103 k a0203 r a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1597.lab vs ../HVITE-M/Rec-mau-M-24/mfG1597.rec
LAB: pau         k a0104 N0204 sh o0304 u0404 pau

REC: pau h a0102 k                      o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1598.lab vs ../HVITE-M/Rec-mau-M-24/mfG1598.rec
LAB: pau           u0102 t a0202 pau
REC: pau k u0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1600.lab vs ../HVITE-M/Rec-mau-M-24/mfG1600.rec
LAB: pau         k             e0103 e0203 s u0303 pau
REC: pau d e0105 k i0205 a0305 g     a0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1601.lab vs ../HVITE-M/Rec-mau-M-24/mfG1601.rec
LAB: pau         u0103 e0203 r u0303 pau
REC: pau m e0103 g     u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1603.lab vs ../HVITE-M/Rec-mau-M-24/mfG1603.rec
LAB: pau   a0104 k a0204 ch a0304 N0404 pau
REC: pau d a0104   N0204 d  a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1605.lab vs ../HVITE-M/Rec-mau-M-24/mfG1605.rec
LAB: pau y o0103 u0203 sh i0303 pau
REC: pau f u0103 u0203 sh i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1606.lab vs ../HVITE-M/Rec-mau-M-24/mfG1606.rec
LAB: pau   b     o0103 k i0203 N0303 pau
REC: pau t o0104 u0204 g o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1608.lab vs ../HVITE-M/Rec-mau-M-24/mfG1608.rec
LAB: pau   m     o0102 sh    i0202 pau
REC: pau b o0103 k     i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1610.lab vs ../HVITE-M/Rec-mau-M-24/mfG1610.rec
LAB: pau ky o0104   u0204 s e0304 i0404 pau
REC: pau t  o0104 k u0204 t e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1611.lab vs ../HVITE-M/Rec-mau-M-24/mfG1611.rec
LAB: pau          k a0103 e0203 r u0303 pau
REC: pau ts u0104 t a0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1613.lab vs ../HVITE-M/Rec-mau-M-24/mfG1613.rec
LAB: pau         h i0103 r o0203 b a0303 pau
REC: pau m i0104 n o0204 g a0304 s u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1614.lab vs ../HVITE-M/Rec-mau-M-24/mfG1614.rec
LAB: pau o0103 m         a0203 k e0303 pau
REC: pau u0104 m a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1616.lab vs ../HVITE-M/Rec-mau-M-24/mfG1616.rec
LAB: pau i0103 k         u0203 r a0303 pau
REC: pau       k a0103 g a0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1619.lab vs ../HVITE-M/Rec-mau-M-24/mfG1619.rec
LAB: pau z e0103   N0203 b u0303 pau

REC: pau n e0103 z u0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1621.lab vs ../HVITE-M/Rec-mau-M-24/mfG1621.rec
LAB: pau s a0103 N0203 j       i0303 pau
REC: pau g a0104 N0204 j o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1622.lab vs ../HVITE-M/Rec-mau-M-24/mfG1622.rec
LAB: pau k o0104 u0204 j i0304 ts u0404 pau
REC: pau k o0104 u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1623.lab vs ../HVITE-M/Rec-mau-M-24/mfG1623.rec
LAB: pau k i0104 N0204 ch    o0304 u0404 pau
REC: pau   g     e0103 N0203 d     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1624.lab vs ../HVITE-M/Rec-mau-M-24/mfG1624.rec
LAB: pau   h     e0102 r i0202 pau
REC: pau k a0103 e0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1625.lab vs ../HVITE-M/Rec-mau-M-24/mfG1625.rec
LAB: pau f       u0102 r     i0202 pau
REC: pau f u0103 k     i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1626.lab vs ../HVITE-M/Rec-mau-M-24/mfG1626.rec
LAB: pau         i0103 z     u0203 m i0303 pau
REC: pau k u0104 r     i0204 i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1628.lab vs ../HVITE-M/Rec-mau-M-24/mfG1628.rec
LAB: pau   a0103 m a0203 i0303 pau
REC: pau t a0103 g a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1629.lab vs ../HVITE-M/Rec-mau-M-24/mfG1629.rec
LAB: pau            u0102 m i0202 pau
REC: pau ch i0103 g i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1631.lab vs ../HVITE-M/Rec-mau-M-24/mfG1631.rec
LAB: pau         i0103 z e0203 N0303 pau
REC: pau g u0104 u0204 z e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1634.lab vs ../HVITE-M/Rec-mau-M-24/mfG1634.rec
LAB: pau               i0102 t     a0202 pau
REC: pau h i0104 N0204 p     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1635.lab vs ../HVITE-M/Rec-mau-M-24/mfG1635.rec
LAB: pau m  i0104 w a0204 t a0304 s u0404 pau
REC: pau ch i0104 r a0204 b a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1636.lab vs ../HVITE-M/Rec-mau-M-24/mfG1636.rec
LAB: pau   e0104 N0204 s o0304 u0404 pau
REC: pau g e0104 N0204 s o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1637.lab vs ../HVITE-M/Rec-mau-M-24/mfG1637.rec
LAB: pau ts u0103 y     o0203 i0303 pau

REC: pau ts u0103 i0203 n     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1638.lab vs ../HVITE-M/Rec-mau-M-24/mfG1638.rec
LAB: pau         k a0102 k u0202 pau
REC: pau s a0103 k u0203 r a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1639.lab vs ../HVITE-M/Rec-mau-M-24/mfG1639.rec
LAB: pau s o0104 u0204 z o0304 u0404 pau
REC: pau k o0104 u0204 z o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1640.lab vs ../HVITE-M/Rec-mau-M-24/mfG1640.rec
LAB: pau s e0104 i0204 t o0304 N0404 pau
REC: pau h e0104 i0204 b o0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1641.lab vs ../HVITE-M/Rec-mau-M-24/mfG1641.rec
LAB: pau f u0103 ry o0203 u0303 pau
REC: pau m u0103 ry o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1642.lab vs ../HVITE-M/Rec-mau-M-24/mfG1642.rec
LAB: pau   a0104 s a0204 b a0304 N0404 pau
REC: pau m a0104 k a0204 n a0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1643.lab vs ../HVITE-M/Rec-mau-M-24/mfG1643.rec
LAB: pau n a0103 o0203 s u0303 pau
REC: pau t a0103 o0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1644.lab vs ../HVITE-M/Rec-mau-M-24/mfG1644.rec
LAB: pau sh u0104 u0204 y     o0304 u0404 pau
REC: pau    k     o0103 u0203 ry    o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1646.lab vs ../HVITE-M/Rec-mau-M-24/mfG1646.rec
LAB: pau                 g o0104 k u0204 r a0304 k u0404 pau
REC: pau p u0105 r o0205 g         u0305 r a0405 m u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1647.lab vs ../HVITE-M/Rec-mau-M-24/mfG1647.rec
LAB: pau s o0103 e0203 r u0303 pau
REC: pau   u0103 e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1648.lab vs ../HVITE-M/Rec-mau-M-24/mfG1648.rec
LAB: pau o0103 w a0203 r u0303 pau
REC: pau       m a0102 r u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1649.lab vs ../HVITE-M/Rec-mau-M-24/mfG1649.rec
LAB: pau m i0103 m     o0203 t         o0303 pau
REC: pau m       o0104 q0204 t o0304 m o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1650.lab vs ../HVITE-M/Rec-mau-M-24/mfG1650.rec
LAB: pau m                o0103 u0203 s u0303 pau
REC: pau m o0104 ts u0204 r     e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1651.lab vs ../HVITE-M/Rec-mau-M-24/mfG1651.rec
LAB: pau                  j a0102 m a0202 pau

REC: pau ch i0104 r a0204 b a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1653.lab vs ../HVITE-M/Rec-mau-M-24/mfG1653.rec
LAB: pau h a0104 n a0204 r e0304 r u0404 pau
REC: pau n a0104 g a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1654.lab vs ../HVITE-M/Rec-mau-M-24/mfG1654.rec
LAB: pau         m u0102 n       e0202 pau
REC: pau s a0104 k u0204 n e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1655.lab vs ../HVITE-M/Rec-mau-M-24/mfG1655.rec
LAB: pau k u0104 u0204 k     o0304 u0404 pau
REC: pau m u0103 sh    i0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1656.lab vs ../HVITE-M/Rec-mau-M-24/mfG1656.rec
LAB: pau y o0103 r o0203 N0303 pau
REC: pau m u0103 r o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1657.lab vs ../HVITE-M/Rec-mau-M-24/mfG1657.rec
LAB: pau ch o0104 u0204 h     o0304 u0404 pau
REC: pau    d     o0103 u0203 m     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1658.lab vs ../HVITE-M/Rec-mau-M-24/mfG1658.rec
LAB: pau k a0103 ch    i0203 k     u0303 pau
REC: pau   i0104 q0204 sh    u0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1659.lab vs ../HVITE-M/Rec-mau-M-24/mfG1659.rec
LAB: pau       sh i0103 b     o0203 u0303 pau
REC: pau u0103 sh       i0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1660.lab vs ../HVITE-M/Rec-mau-M-24/mfG1660.rec
LAB: pau            r     e0102 i0202 pau
REC: pau sh i0103 t e0203 k     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1661.lab vs ../HVITE-M/Rec-mau-M-24/mfG1661.rec
LAB: pau               k a0103 i0203 j o0303 pau
REC: pau g a0104 i0204 k       o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1662.lab vs ../HVITE-M/Rec-mau-M-24/mfG1662.rec
LAB: pau r o0104 q0204 k     a0304 a0404 pau
REC: pau   d     o0103 u0203 s     a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1663.lab vs ../HVITE-M/Rec-mau-M-24/mfG1663.rec
LAB: pau k i0102 j i0202 pau
REC: pau r i0102 k a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1664.lab vs ../HVITE-M/Rec-mau-M-24/mfG1664.rec
LAB: pau k a0104 m i0204 k u0304 z u0404 pau
REC: pau k         a0103 k u0203 d o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1665.lab vs ../HVITE-M/Rec-mau-M-24/mfG1665.rec
LAB: pau k a0102 g       e0202 pau

REC: pau k a0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1667.lab vs ../HVITE-M/Rec-mau-M-24/mfG1667.rec
LAB: pau   a0103 m a0203 g u0303 pau
REC: pau n a0103 r a0203 b u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1669.lab vs ../HVITE-M/Rec-mau-M-24/mfG1669.rec
LAB: pau i0104 k     u0204 s e0304 i0404 pau
REC: pau g     o0104 u0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1671.lab vs ../HVITE-M/Rec-mau-M-24/mfG1671.rec
LAB: pau   a0102 n       i0202 pau
REC: pau t a0103 n i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1672.lab vs ../HVITE-M/Rec-mau-M-24/mfG1672.rec
LAB: pau ky o0104 u0204 d a0304 i0404 pau
REC: pau    g     o0103 r a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1673.lab vs ../HVITE-M/Rec-mau-M-24/mfG1673.rec
LAB: pau h a0104 i0204 t a0304 ts u0404 pau
REC: pau k a0104 i0204 k a0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1674.lab vs ../HVITE-M/Rec-mau-M-24/mfG1674.rec
LAB: pau   ts    u0103 k e0203 r     u0303 pau
REC: pau h e0104 N0204 k       o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1677.lab vs ../HVITE-M/Rec-mau-M-24/mfG1677.rec
LAB: pau   a0104 ts u0204 ry o0304 k u0404 pau
REC: pau g a0104 k  u0204 ry o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1678.lab vs ../HVITE-M/Rec-mau-M-24/mfG1678.rec
LAB: pau a0104 j i0204 w a0304 u0404 pau
REC: pau u0104 k a0204 g a0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1679.lab vs ../HVITE-M/Rec-mau-M-24/mfG1679.rec
LAB: pau k a0104 N0204 sh i0304 N0404 pau
REC: pau t a0104 N0204 k  e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1681.lab vs ../HVITE-M/Rec-mau-M-24/mfG1681.rec
LAB: pau k o0104 k u0204 s a0304 N0404 pau
REC: pau s e0104 k i0204 t a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1682.lab vs ../HVITE-M/Rec-mau-M-24/mfG1682.rec
LAB: pau    i0103 k u0203 r         a0303 pau
REC: pau sh i0103         r a0203 g a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1683.lab vs ../HVITE-M/Rec-mau-M-24/mfG1683.rec
LAB: pau         m     i0104 o0204 r o0304 s u0404 pau
REC: pau y o0104 u0204 s     u0304 r         u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1685.lab vs ../HVITE-M/Rec-mau-M-24/mfG1685.rec
LAB: pau r o0104 N0204 s o0304 u0404 pau

REC: pau k o0104 u0204 s o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1687.lab vs ../HVITE-M/Rec-mau-M-24/mfG1687.rec
LAB: pau       y o0102 b i0202 pau
REC: pau o0103 y o0203 b i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1688.lab vs ../HVITE-M/Rec-mau-M-24/mfG1688.rec
LAB: pau k     a0104 i0204 h a0304 ts u0404 pau
REC: pau u0104 r     i0204 d a0304 s  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1689.lab vs ../HVITE-M/Rec-mau-M-24/mfG1689.rec
LAB: pau   sh    i0102 n     a0202 pau
REC: pau k i0103 g     a0203 e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1690.lab vs ../HVITE-M/Rec-mau-M-24/mfG1690.rec
LAB: pau                  sh i0102 t a0202 pau
REC: pau ch i0104 r a0204 b  a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1691.lab vs ../HVITE-M/Rec-mau-M-24/mfG1691.rec
LAB: pau ky o0103 z e0203 ts u0303 pau
REC: pau ts u0103 r u0203 s  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1692.lab vs ../HVITE-M/Rec-mau-M-24/mfG1692.rec
LAB: pau s u0103 g a0203 t         a0303 pau
REC: pau h e0104 d a0204 t a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1693.lab vs ../HVITE-M/Rec-mau-M-24/mfG1693.rec
LAB: pau   a0104 N0204 z a0304 N0404 pau
REC: pau d a0104 N0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1694.lab vs ../HVITE-M/Rec-mau-M-24/mfG1694.rec
LAB: pau ch o0103 u0203 s a0303 pau
REC: pau k  o0103 u0203 z a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1698.lab vs ../HVITE-M/Rec-mau-M-24/mfG1698.rec
LAB: pau   o0102 b                 i0202 pau
REC: pau t o0104 b i0204 k o0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1699.lab vs ../HVITE-M/Rec-mau-M-24/mfG1699.rec
LAB: pau ch u0104 u0204 j u0304 N0404 pau
REC: pau    k     i0103 j u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1700.lab vs ../HVITE-M/Rec-mau-M-24/mfG1700.rec
LAB: pau ky u0104 u0204 k u0304 ts u0404 pau
REC: pau    g     i0103 j u0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1701.lab vs ../HVITE-M/Rec-mau-M-24/mfG1701.rec
LAB: pau           o0102 n i0202 pau
REC: pau t o0103 m o0203 n i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1703.lab vs ../HVITE-M/Rec-mau-M-24/mfG1703.rec
LAB: pau o0105 ch i0205 b u0305 r e0405 r u0505 pau

REC: pau u0105 r  i0205 k i0305 r e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1706.lab vs ../HVITE-M/Rec-mau-M-24/mfG1706.rec
LAB: pau f u0103         k u0203 r o0303 pau
REC: pau f u0103 r o0203 k         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1707.lab vs ../HVITE-M/Rec-mau-M-24/mfG1707.rec
LAB: pau ts u0102 k u0202 pau
REC: pau    u0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1710.lab vs ../HVITE-M/Rec-mau-M-24/mfG1710.rec
LAB: pau sh i0104 n a0204 b i0304 r u0404 pau
REC: pau d  o0103 n         a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1711.lab vs ../HVITE-M/Rec-mau-M-24/mfG1711.rec
LAB: pau k e0104 N0204 s e0304 ts    u0404 pau
REC: pau   t     e0103 s       u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1714.lab vs ../HVITE-M/Rec-mau-M-24/mfG1714.rec
LAB: pau k a0103 e0203 r u0303 pau
REC: pau h a0103 e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1716.lab vs ../HVITE-M/Rec-mau-M-24/mfG1716.rec
LAB: pau r a0103 j i0203   o0303 pau
REC: pau k a0103 j i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1717.lab vs ../HVITE-M/Rec-mau-M-24/mfG1717.rec
LAB: pau u0104 q0204 t o0304 r     i0404 pau
REC: pau f     u0103 t       o0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1718.lab vs ../HVITE-M/Rec-mau-M-24/mfG1718.rec
LAB: pau n a0103 r e0203 r u0303 pau
REC: pau t a0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1719.lab vs ../HVITE-M/Rec-mau-M-24/mfG1719.rec
LAB: pau s o0103 n o0203 t a0303 pau
REC: pau         g o0102 z a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1722.lab vs ../HVITE-M/Rec-mau-M-24/mfG1722.rec
LAB: pau k     e0104 q0204 k o0304 u0404 pau
REC: pau u0103 g     o0203 k       i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1724.lab vs ../HVITE-M/Rec-mau-M-24/mfG1724.rec
LAB: pau u0105 s u0205 g u0305 r a0405 i0505 pau
REC: pau       m i0103 g a0203 r       u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1725.lab vs ../HVITE-M/Rec-mau-M-24/mfG1725.rec
LAB: pau           f     u0103 k e0203 r u0303 pau
REC: pau h i0105 k i0205 u0305 k e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1726.lab vs ../HVITE-M/Rec-mau-M-24/mfG1726.rec
LAB: pau m i0104 w a0204 t a0304 s          u0404 pau

REC: pau m                 a0103 s a0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1727.lab vs ../HVITE-M/Rec-mau-M-24/mfG1727.rec
LAB: pau ry o0103 k                       a0203 N0303 pau
REC: pau          k o0104 r o0204 g a0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1729.lab vs ../HVITE-M/Rec-mau-M-24/mfG1729.rec
LAB: pau         i0104 q0204 s a0304   i0404 pau
REC: pau b u0104 ts    u0204 k a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1730.lab vs ../HVITE-M/Rec-mau-M-24/mfG1730.rec
LAB: pau                       ts u0102 ch i0202 pau
REC: pau z u0104 i0204 h i0304 ts          u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1732.lab vs ../HVITE-M/Rec-mau-M-24/mfG1732.rec
LAB: pau                  j a0101 pau
REC: pau sh i0103 k a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1733.lab vs ../HVITE-M/Rec-mau-M-24/mfG1733.rec
LAB: pau    t     o0103 k u0203 sh    u0303 pau
REC: pau ch u0104 u0204 k       e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1736.lab vs ../HVITE-M/Rec-mau-M-24/mfG1736.rec
LAB: pau m o0105 u0205 sh i0305 k o0405 m u0505 pau
REC: pau         o0104 sh i0204 k o0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1737.lab vs ../HVITE-M/Rec-mau-M-24/mfG1737.rec
LAB: pau d a0102 N0202 pau
REC: pau g a0102 N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1738.lab vs ../HVITE-M/Rec-mau-M-24/mfG1738.rec
LAB: pau ts u0104 k i0204 s a0304 s u0404 pau
REC: pau          h i0103 t a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1740.lab vs ../HVITE-M/Rec-mau-M-24/mfG1740.rec
LAB: pau   e0103 n o0203 g u0303 pau
REC: pau n e0103 m u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1741.lab vs ../HVITE-M/Rec-mau-M-24/mfG1741.rec
LAB: pau               f u0102 d a0202 pau
REC: pau k o0104 u0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1742.lab vs ../HVITE-M/Rec-mau-M-24/mfG1742.rec
LAB: pau   u0103 w a0203 k i0303 pau
REC: pau k a0103 w a0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1743.lab vs ../HVITE-M/Rec-mau-M-24/mfG1743.rec
LAB: pau f u0104 k i0204 d a0304 s u0404 pau
REC: pau         k u0103 r a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1744.lab vs ../HVITE-M/Rec-mau-M-24/mfG1744.rec
LAB: pau j o0104 u0204 ky    u0304 u0404 pau

REC: pau d o0103 ry    o0203 k     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1745.lab vs ../HVITE-M/Rec-mau-M-24/mfG1745.rec
LAB: pau         s  a0103 s u0203 g a0303 pau
REC: pau d a0104 ts u0204 r a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1746.lab vs ../HVITE-M/Rec-mau-M-24/mfG1746.rec
LAB: pau   m     a0102 k     u0202 pau
REC: pau t a0103 b     u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1747.lab vs ../HVITE-M/Rec-mau-M-24/mfG1747.rec
LAB: pau k e0103 m u0203 r          i0303 pau
REC: pau                 r i0102 ts u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1748.lab vs ../HVITE-M/Rec-mau-M-24/mfG1748.rec
LAB: pau d e0103 N0203 k i0303 pau
REC: pau t e0103 N0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1750.lab vs ../HVITE-M/Rec-mau-M-24/mfG1750.rec
LAB: pau ch u0104 u0204 sh i0304 N0404 pau
REC: pau k  e0104 i0204 h  i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1751.lab vs ../HVITE-M/Rec-mau-M-24/mfG1751.rec
LAB: pau           w     a0103 i0203 r o0303 pau
REC: pau k o0104 m a0204 g     o0304 m a0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1752.lab vs ../HVITE-M/Rec-mau-M-24/mfG1752.rec
LAB: pau sh i0103 t     a0203 i0303 pau
REC: pau    b     a0103 i0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1753.lab vs ../HVITE-M/Rec-mau-M-24/mfG1753.rec
LAB: pau s u0104 q0204 p     a0304 i0404 pau
REC: pau   k     a0103 i0203 g     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1754.lab vs ../HVITE-M/Rec-mau-M-24/mfG1754.rec
LAB: pau h a0103 r a0203   u0303 pau
REC: pau d a0103 r a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1755.lab vs ../HVITE-M/Rec-mau-M-24/mfG1755.rec
LAB: pau k i0103 b o0203 u0303 pau
REC: pau s e0103 r o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1756.lab vs ../HVITE-M/Rec-mau-M-24/mfG1756.rec
LAB: pau n i0104 N0204 m e0304 i0404 pau
REC: pau k i0104 N0204 b e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1757.lab vs ../HVITE-M/Rec-mau-M-24/mfG1757.rec
LAB: pau t a0104   i0204 t e0304 i0404 pau
REC: pau k a0104 k u0204 t e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1758.lab vs ../HVITE-M/Rec-mau-M-24/mfG1758.rec
LAB: pau         s o0103 r e0203 r u0303 pau

REC: pau k u0104 z u0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1760.lab vs ../HVITE-M/Rec-mau-M-24/mfG1760.rec
LAB: pau                 a0103 o0203 i0303 pau
REC: pau u0104 r a0204 g i0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1761.lab vs ../HVITE-M/Rec-mau-M-24/mfG1761.rec
LAB: pau   y     o0103 u0203 t     o0303 pau
REC: pau t o0104 u0204 s     e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1763.lab vs ../HVITE-M/Rec-mau-M-24/mfG1763.rec
LAB: pau         z u0103 r e0203 r u0303 pau
REC: pau n a0104 z u0204 k e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1765.lab vs ../HVITE-M/Rec-mau-M-24/mfG1765.rec
LAB: pau h o0102 sh i0202 pau
REC: pau t o0102 k  i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1766.lab vs ../HVITE-M/Rec-mau-M-24/mfG1766.rec
LAB: pau         h i0103 m i0203 ts u0303 pau
REC: pau k a0104 k u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1767.lab vs ../HVITE-M/Rec-mau-M-24/mfG1767.rec
LAB: pau         h a0101 pau
REC: pau g a0102 r a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1768.lab vs ../HVITE-M/Rec-mau-M-24/mfG1768.rec
LAB: pau           o0102 k u0202 pau
REC: pau s o0103 s o0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1769.lab vs ../HVITE-M/Rec-mau-M-24/mfG1769.rec
LAB: pau k i0103 n     o0203 u0303 pau
REC: pau k       u0102 m     o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1771.lab vs ../HVITE-M/Rec-mau-M-24/mfG1771.rec
LAB: pau t a0103 t e0203 r u0303 pau
REC: pau h a0103 g e0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1773.lab vs ../HVITE-M/Rec-mau-M-24/mfG1773.rec
LAB: pau k i0103 k a0203 i0303 pau
REC: pau r i0103 g a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1774.lab vs ../HVITE-M/Rec-mau-M-24/mfG1774.rec
LAB: pau         n       a0101 pau
REC: pau h i0103 n a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1775.lab vs ../HVITE-M/Rec-mau-M-24/mfG1775.rec
LAB: pau t a0103 m e0203 r u0303 pau
REC: pau s a0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1776.lab vs ../HVITE-M/Rec-mau-M-24/mfG1776.rec
LAB: pau         t a0102 n a0202 pau

REC: pau d a0103 r a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1777.lab vs ../HVITE-M/Rec-mau-M-24/mfG1777.rec
LAB: pau f u0103 sh o0203 u0303 pau
REC: pau m i0103 h  o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1778.lab vs ../HVITE-M/Rec-mau-M-24/mfG1778.rec
LAB: pau ts u0103 m e0203 r u0303 pau
REC: pau    u0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1779.lab vs ../HVITE-M/Rec-mau-M-24/mfG1779.rec
LAB: pau         y o0102 u0202 pau
REC: pau m u0103 r o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1780.lab vs ../HVITE-M/Rec-mau-M-24/mfG1780.rec
LAB: pau k a0103 k e0203 r u0303 pau
REC: pau b a0103 k e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1781.lab vs ../HVITE-M/Rec-mau-M-24/mfG1781.rec
LAB: pau a0104 k u0204 m a0304 d e0404 pau
REC: pau         a0103 m a0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1784.lab vs ../HVITE-M/Rec-mau-M-24/mfG1784.rec
LAB: pau                    g     a0102 N0202 pau
REC: pau ch i0104 k a0204 m i0304 ch    i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1785.lab vs ../HVITE-M/Rec-mau-M-24/mfG1785.rec
LAB: pau y a0104 q0204 k a0304 i0404 pau
REC: pau t a0104 i0204 g a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1786.lab vs ../HVITE-M/Rec-mau-M-24/mfG1786.rec
LAB: pau sh i0103       j o0203 u0303 pau
REC: pau k  i0103 N0203 j       o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1787.lab vs ../HVITE-M/Rec-mau-M-24/mfG1787.rec
LAB: pau         k a0104 ts u0204 d     o0304 u0404 pau
REC: pau b a0103 k                u0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1789.lab vs ../HVITE-M/Rec-mau-M-24/mfG1789.rec
LAB: pau sh u0103 ry o0203 k u0303 pau
REC: pau k  i0103 ry o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1790.lab vs ../HVITE-M/Rec-mau-M-24/mfG1790.rec
LAB: pau t a0102 n a0202 pau
REC: pau g a0102 r a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1791.lab vs ../HVITE-M/Rec-mau-M-24/mfG1791.rec
LAB: pau h i0103 h a0203 N0303 pau
REC: pau k i0103 b a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1792.lab vs ../HVITE-M/Rec-mau-M-24/mfG1792.rec
LAB: pau         u0103 w     a0203 k     i0303 pau

REC: pau t a0104 d     a0204 sh    i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1793.lab vs ../HVITE-M/Rec-mau-M-24/mfG1793.rec
LAB: pau k e0104 q0204 k a0304 k u0404 pau
REC: pau   t     a0103 m a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1794.lab vs ../HVITE-M/Rec-mau-M-24/mfG1794.rec
LAB: pau g e0104 i0204 j u0304 ts u0404 pau
REC: pau g e0104 N0204 j i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1795.lab vs ../HVITE-M/Rec-mau-M-24/mfG1795.rec
LAB: pau z a0104 r a0204 z a0304 r a0404 pau
REC: pau         n a0103 g a0203 s a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1796.lab vs ../HVITE-M/Rec-mau-M-24/mfG1796.rec
LAB: pau t a0104 q0204 s  u0304 r  u0404 pau
REC: pau               sh a0102 ts u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1797.lab vs ../HVITE-M/Rec-mau-M-24/mfG1797.rec
LAB: pau   a0102 k u0202 pau
REC: pau k a0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1798.lab vs ../HVITE-M/Rec-mau-M-24/mfG1798.rec
LAB: pau s a0103 k u0203 sh a0303 pau
REC: pau k a0103 k a0203 r  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1799.lab vs ../HVITE-M/Rec-mau-M-24/mfG1799.rec
LAB: pau   o0104 b o0204 e0304 r u0404 pau
REC: pau k o0104 g o0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1801.lab vs ../HVITE-M/Rec-mau-M-24/mfG1801.rec
LAB: pau   h     i0103 k a0203 N0303 pau
REC: pau k u0104 u0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1804.lab vs ../HVITE-M/Rec-mau-M-24/mfG1804.rec
LAB: pau   sh    u0102 sh    i0202 pau
REC: pau k i0103 h     o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1806.lab vs ../HVITE-M/Rec-mau-M-24/mfG1806.rec
LAB: pau         b a0102 s u0202 pau
REC: pau p u0103 r a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1808.lab vs ../HVITE-M/Rec-mau-M-24/mfG1808.rec
LAB: pau               h e0102 r  i0202 pau
REC: pau m e0104 i0204 b u0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1809.lab vs ../HVITE-M/Rec-mau-M-24/mfG1809.rec
LAB: pau k o0104 u0204 b a0304 i0404 pau
REC: pau b o0104 u0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1810.lab vs ../HVITE-M/Rec-mau-M-24/mfG1810.rec
LAB: pau h a0104 k       i0204 d a0304 s u0404 pau

REC: pau         k i0104 N0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1811.lab vs ../HVITE-M/Rec-mau-M-24/mfG1811.rec
LAB: pau                y o0103 ky o0203 u0303 pau
REC: pau ry o0104 u0204 y          o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1812.lab vs ../HVITE-M/Rec-mau-M-24/mfG1812.rec
LAB: pau k e0103 i0203 sh a0303 pau
REC: pau t e0103 i0203 k  a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1813.lab vs ../HVITE-M/Rec-mau-M-24/mfG1813.rec
LAB: pau f u0103 u0203 f     u0303 pau
REC: pau m u0103 k     o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1814.lab vs ../HVITE-M/Rec-mau-M-24/mfG1814.rec
LAB: pau n  o0103 k o0203 r u0303 pau
REC: pau ts u0103 m o0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1816.lab vs ../HVITE-M/Rec-mau-M-24/mfG1816.rec
LAB: pau    o0103 k i0203 r u0303 pau
REC: pau ch i0103 g i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1817.lab vs ../HVITE-M/Rec-mau-M-24/mfG1817.rec
LAB: pau   a0104 N0204 z e0304 N0404 pau
REC: pau t a0104 N0204 k e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1818.lab vs ../HVITE-M/Rec-mau-M-24/mfG1818.rec
LAB: pau         ts u0103 m     a0203 r     i0303 pau
REC: pau j i0104 ts       u0204 z     a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1821.lab vs ../HVITE-M/Rec-mau-M-24/mfG1821.rec
LAB: pau f u0104 m i0204 k i0304 r i0404 pau
REC: pau         m         i0102 g i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1822.lab vs ../HVITE-M/Rec-mau-M-24/mfG1822.rec
LAB: pau                  m a0102 t a0202 pau
REC: pau sh a0103 g a0203 m         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1823.lab vs ../HVITE-M/Rec-mau-M-24/mfG1823.rec
LAB: pau k u0103 r o0203 u0303 pau
REC: pau d o0102 r       o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1824.lab vs ../HVITE-M/Rec-mau-M-24/mfG1824.rec
LAB: pau   h     o0103 u0203 sh    i0303 pau
REC: pau z o0104 u0204 k     i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1826.lab vs ../HVITE-M/Rec-mau-M-24/mfG1826.rec
LAB: pau s a0105 k a0205 n o0305 b o0405 r u0505 pau
REC: pau t o0105 r i0205 m o0305 d o0405 s u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1827.lab vs ../HVITE-M/Rec-mau-M-24/mfG1827.rec
LAB: pau s  u0104 t a0204   N0304 p u0404 pau

REC: pau ts u0104 r a0204 n u0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1828.lab vs ../HVITE-M/Rec-mau-M-24/mfG1828.rec
LAB: pau sh o0106 u0206 g a0306 q0406 k o0506 u0606 pau
REC: pau                k o0104 q0204 k o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1830.lab vs ../HVITE-M/Rec-mau-M-24/mfG1830.rec
LAB: pau         s o0103 r e0203 r u0303 pau
REC: pau k u0104 z u0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1832.lab vs ../HVITE-M/Rec-mau-M-24/mfG1832.rec
LAB: pau                 s  a0102 k u0202 pau
REC: pau g a0104 k u0204 ry o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1833.lab vs ../HVITE-M/Rec-mau-M-24/mfG1833.rec
LAB: pau n i0103 m o0203 ts u0303 pau
REC: pau m i0103 n u0203 k  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1836.lab vs ../HVITE-M/Rec-mau-M-24/mfG1836.rec
LAB: pau h  a0104 q0204 t e0304 N0404 pau
REC: pau sh i0104 N0204 r i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1837.lab vs ../HVITE-M/Rec-mau-M-24/mfG1837.rec
LAB: pau k a0104 ts u0204 y  o0304   u0404 pau
REC: pau g a0104 k  u0204 ry o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1838.lab vs ../HVITE-M/Rec-mau-M-24/mfG1838.rec
LAB: pau               w a0102 N0202 pau
REC: pau k a0104 N0204 j i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1839.lab vs ../HVITE-M/Rec-mau-M-24/mfG1839.rec
LAB: pau o0104 d o0204 k a0304 s u0404 pau
REC: pau       t o0103 k a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1840.lab vs ../HVITE-M/Rec-mau-M-24/mfG1840.rec
LAB: pau y o0103 s o0203 k u0303 pau
REC: pau y o0103 k o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1841.lab vs ../HVITE-M/Rec-mau-M-24/mfG1841.rec
LAB: pau           k     o0103 q0203 k a0303 pau
REC: pau t a0104 n o0204 sh    i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1843.lab vs ../HVITE-M/Rec-mau-M-24/mfG1843.rec
LAB: pau n a0105 g u0205 s a0305 m e0405 r u0505 pau
REC: pau         t a0104 k a0204 m e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1844.lab vs ../HVITE-M/Rec-mau-M-24/mfG1844.rec
LAB: pau   n     i0103 j u0203 u0303 pau
REC: pau s u0104 i0204 j u0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1845.lab vs ../HVITE-M/Rec-mau-M-24/mfG1845.rec
LAB: pau sh o0104 u0204 d a0304 k u0404 pau

REC: pau k  o0104 u0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1847.lab vs ../HVITE-M/Rec-mau-M-24/mfG1847.rec
LAB: pau n a0103 k a0203 m i0303 pau
REC: pau k a0103 g a0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1849.lab vs ../HVITE-M/Rec-mau-M-24/mfG1849.rec
LAB: pau h  i0103 r a0203 k u0303 pau
REC: pau sh i0103 b a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1852.lab vs ../HVITE-M/Rec-mau-M-24/mfG1852.rec
LAB: pau h a0103   i0203 ry o0303 pau
REC: pau m a0103 r u0203 d  e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1854.lab vs ../HVITE-M/Rec-mau-M-24/mfG1854.rec
LAB: pau w a0104 s u0204 r e0304 r u0404 pau
REC: pau         k a0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1855.lab vs ../HVITE-M/Rec-mau-M-24/mfG1855.rec
LAB: pau k e0104 N0204 t e0304 i0404 pau
REC: pau s e0104 N0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1856.lab vs ../HVITE-M/Rec-mau-M-24/mfG1856.rec
LAB: pau k e0104 ts u0204 r o0304 N0404 pau
REC: pau         m  u0103 r o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1858.lab vs ../HVITE-M/Rec-mau-M-24/mfG1858.rec
LAB: pau ch i0104 j i0204 r e0304 r u0404 pau
REC: pau k  o0104 j i0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1859.lab vs ../HVITE-M/Rec-mau-M-24/mfG1859.rec
LAB: pau             k i0102 r i0202 pau
REC: pau i0103 N0203 k         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1861.lab vs ../HVITE-M/Rec-mau-M-24/mfG1861.rec
LAB: pau   a0104 k a0204 r u0304 i0404 pau
REC: pau k a0103 k a0203 r       u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1863.lab vs ../HVITE-M/Rec-mau-M-24/mfG1863.rec
LAB: pau t e0104 q0204 k a0304 k u0404 pau
REC: pau r i0104 N0204 k a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1865.lab vs ../HVITE-M/Rec-mau-M-24/mfG1865.rec
LAB: pau n a0103 k     a0203 m     i0303 pau
REC: pau g a0104 N0204 r     a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1866.lab vs ../HVITE-M/Rec-mau-M-24/mfG1866.rec
LAB: pau       s e0103 m a0203 r u0303 pau
REC: pau u0104 z u0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1867.lab vs ../HVITE-M/Rec-mau-M-24/mfG1867.rec
LAB: pau   y     u0103 t     o0203 r i0303 pau

REC: pau f u0104 g     o0204 u0304 r i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1868.lab vs ../HVITE-M/Rec-mau-M-24/mfG1868.rec
LAB: pau b e0104 N0204 sh    o0304 u0404 pau
REC: pau   ch    o0103 q0203 t     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1870.lab vs ../HVITE-M/Rec-mau-M-24/mfG1870.rec
LAB: pau ch i0102   i0202 pau
REC: pau ch i0102 r u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1871.lab vs ../HVITE-M/Rec-mau-M-24/mfG1871.rec
LAB: pau h i0103 n i0203 ch i0303 pau
REC: pau h i0103 n a0203 t  a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1872.lab vs ../HVITE-M/Rec-mau-M-24/mfG1872.rec
LAB: pau k i0103 k     a0203 i0303 pau
REC: pau k a0104 i0204 i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1873.lab vs ../HVITE-M/Rec-mau-M-24/mfG1873.rec
LAB: pau t e0104 i0204 n e0304 i0404 pau
REC: pau r e0104 N0204 m e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1876.lab vs ../HVITE-M/Rec-mau-M-24/mfG1876.rec
LAB: pau m o0103 u0203 k     e0303 pau
REC: pau b o0104 u0204 e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1877.lab vs ../HVITE-M/Rec-mau-M-24/mfG1877.rec
LAB: pau u0105 ts u0205 k u0305 sh i0405 i0505 pau
REC: pau                k u0102 sh       i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1878.lab vs ../HVITE-M/Rec-mau-M-24/mfG1878.rec
LAB: pau ts         u0103 ts u0203 k u0303 pau
REC: pau ts u0104 k u0204 z  u0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1879.lab vs ../HVITE-M/Rec-mau-M-24/mfG1879.rec
LAB: pau ch o0104 q0204 k a0304 N0404 pau
REC: pau b  u0104 N0204 t a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1882.lab vs ../HVITE-M/Rec-mau-M-24/mfG1882.rec
LAB: pau h i0105 k i0205 ts u0305 k e0405 r u0505 pau
REC: pau h i0105 k i0205    u0305 k e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1883.lab vs ../HVITE-M/Rec-mau-M-24/mfG1883.rec
LAB: pau                  n         a0101 pau
REC: pau ts u0104 r a0204 n u0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1884.lab vs ../HVITE-M/Rec-mau-M-24/mfG1884.rec
LAB: pau         y o0103 s a0203 N0303 pau
REC: pau k o0104 k u0204 s a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1885.lab vs ../HVITE-M/Rec-mau-M-24/mfG1885.rec
LAB: pau sh i0103 k a0203 k e0303 pau

REC: pau          k a0102 g e0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1888.lab vs ../HVITE-M/Rec-mau-M-24/mfG1888.rec
LAB: pau y u0103 t a0203 k a0303 pau
REC: pau s a0103 k a0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1889.lab vs ../HVITE-M/Rec-mau-M-24/mfG1889.rec
LAB: pau         n     o0103 N0203 k i0303 pau
REC: pau r o0104 N0204 j     i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1890.lab vs ../HVITE-M/Rec-mau-M-24/mfG1890.rec
LAB: pau f u0102 k                u0202 pau
REC: pau         k u0103 ts u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1891.lab vs ../HVITE-M/Rec-mau-M-24/mfG1891.rec
LAB: pau                 s a0102 t e0202 pau
REC: pau t a0103 r a0203 s         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1892.lab vs ../HVITE-M/Rec-mau-M-24/mfG1892.rec
LAB: pau h o0103 t e0203 r u0303 pau
REC: pau k o0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1893.lab vs ../HVITE-M/Rec-mau-M-24/mfG1893.rec
LAB: pau   a0104 y a0204 m a0304 r u0404 pau
REC: pau k a0104 t a0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1895.lab vs ../HVITE-M/Rec-mau-M-24/mfG1895.rec
LAB: pau k o0104 u0204 s o0304 u0404 pau
REC: pau h o0104 u0204 h o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1896.lab vs ../HVITE-M/Rec-mau-M-24/mfG1896.rec
LAB: pau         ch i0102 i0202 pau
REC: pau n i0103 j  u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1897.lab vs ../HVITE-M/Rec-mau-M-24/mfG1897.rec
LAB: pau t a0104 m a0204 sh i0304   i0404 pau
REC: pau h a0104   q0204 k  i0304 r i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1898.lab vs ../HVITE-M/Rec-mau-M-24/mfG1898.rec
LAB: pau h i0105 k         i0205 a0305 g e0405 r u0505 pau
REC: pau         k u0105 r i0205 a0305 g e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1899.lab vs ../HVITE-M/Rec-mau-M-24/mfG1899.rec
LAB: pau s e0103 r               o0203 N0303 pau
REC: pau         r i0103 z u0203 m     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1900.lab vs ../HVITE-M/Rec-mau-M-24/mfG1900.rec
LAB: pau         k a0103 ch i0203 k u0303 pau
REC: pau m a0104 g i0204 r  e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1901.lab vs ../HVITE-M/Rec-mau-M-24/mfG1901.rec
LAB: pau s e0104 ts u0204 m e0304 i0404 pau

REC: pau            u0102 m       e0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1902.lab vs ../HVITE-M/Rec-mau-M-24/mfG1902.rec
LAB: pau w a0102 k i0202 pau
REC: pau t a0102 k i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1903.lab vs ../HVITE-M/Rec-mau-M-24/mfG1903.rec
LAB: pau t e0104 i0204 sh u0304 ts u0404 pau
REC: pau g e0104 N0204 j  i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1904.lab vs ../HVITE-M/Rec-mau-M-24/mfG1904.rec
LAB: pau n a0104 i0204 sh o0304 k u0404 pau
REC: pau t a0104 i0204 sh o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1905.lab vs ../HVITE-M/Rec-mau-M-24/mfG1905.rec
LAB: pau           o0102 k u0202 pau
REC: pau t o0103 k a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1906.lab vs ../HVITE-M/Rec-mau-M-24/mfG1906.rec
LAB: pau n u0102 g u0202 pau
REC: pau k u0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1907.lab vs ../HVITE-M/Rec-mau-M-24/mfG1907.rec
LAB: pau s e0104 i0204 j i0304 N0404 pau
REC: pau   g     e0103 j u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1912.lab vs ../HVITE-M/Rec-mau-M-24/mfG1912.rec
LAB: pau s u0102 m i0202 pau
REC: pau b u0102   N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1913.lab vs ../HVITE-M/Rec-mau-M-24/mfG1913.rec
LAB: pau j i0104 q0204 k e0304 N0404 pau
REC: pau k o0104 u0204 k e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1914.lab vs ../HVITE-M/Rec-mau-M-24/mfG1914.rec
LAB: pau j u0103 my    o0203 u0303 pau
REC: pau k o0103 u0203 ry    o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1915.lab vs ../HVITE-M/Rec-mau-M-24/mfG1915.rec
LAB: pau o0103 j i0203 g i0303 pau
REC: pau       k i0102 r i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1916.lab vs ../HVITE-M/Rec-mau-M-24/mfG1916.rec
LAB: pau         m a0102 ts u0202 pau
REC: pau b a0103 k u0203 r  o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1917.lab vs ../HVITE-M/Rec-mau-M-24/mfG1917.rec
LAB: pau k o0104 z u0204 ts u0304 m i0404 pau
REC: pau y u0103 z          u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1918.lab vs ../HVITE-M/Rec-mau-M-24/mfG1918.rec
LAB: pau y a0103 s u0203 m i0303 pau

REC: pau t a0103 k u0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1919.lab vs ../HVITE-M/Rec-mau-M-24/mfG1919.rec
LAB: pau           o0102 b i0202 pau
REC: pau k o0103 r i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1920.lab vs ../HVITE-M/Rec-mau-M-24/mfG1920.rec
LAB: pau         a0102 j     i0202 pau
REC: pau k a0103 k     o0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1921.lab vs ../HVITE-M/Rec-mau-M-24/mfG1921.rec
LAB: pau k u0102 m i0202 pau
REC: pau d e0102 m o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1922.lab vs ../HVITE-M/Rec-mau-M-24/mfG1922.rec
LAB: pau ky u0104 u0204 g e0304 k i0404 pau
REC: pau r  e0103 i0203 g         i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1923.lab vs ../HVITE-M/Rec-mau-M-24/mfG1923.rec
LAB: pau          f u0102 n e0202 pau
REC: pau ts u0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1924.lab vs ../HVITE-M/Rec-mau-M-24/mfG1924.rec
LAB: pau i0104 ch i0204 n i0304 ch i0404 pau
REC: pau       h  i0103 n i0203 ch i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1926.lab vs ../HVITE-M/Rec-mau-M-24/mfG1926.rec
LAB: pau j o0104 u0204 j u0304 N0404 pau
REC: pau r o0104 u0204 j i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1927.lab vs ../HVITE-M/Rec-mau-M-24/mfG1927.rec
LAB: pau o0105 s     o0205 r     o0305 sh    i0405 i0505 pau
REC: pau z     o0104 k     u0204 z     o0304 k     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1929.lab vs ../HVITE-M/Rec-mau-M-24/mfG1929.rec
LAB: pau ch o0104 u0204 sh u0304 u0404 pau
REC: pau k  o0104 u0204 sh u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1930.lab vs ../HVITE-M/Rec-mau-M-24/mfG1930.rec
LAB: pau k i0103 m o0203 n o0303 pau
REC: pau   u0103 g o0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1931.lab vs ../HVITE-M/Rec-mau-M-24/mfG1931.rec
LAB: pau               j i0102 k o0202 pau
REC: pau h o0104 u0204 k o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1932.lab vs ../HVITE-M/Rec-mau-M-24/mfG1932.rec
LAB: pau s o0104 r o0204 e0304 r u0404 pau
REC: pau t a0104 t o0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1933.lab vs ../HVITE-M/Rec-mau-M-24/mfG1933.rec
LAB: pau sh i0104 ts u0204 gy o0304 u0404 pau

REC: pau n  i0104    N0204 gy o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1934.lab vs ../HVITE-M/Rec-mau-M-24/mfG1934.rec
LAB: pau         h i0101 pau
REC: pau b i0102 r u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1935.lab vs ../HVITE-M/Rec-mau-M-24/mfG1935.rec
LAB: pau o0104 o0204 k               i0304 i0404 pau
REC: pau             k e0103 m u0203 r     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1936.lab vs ../HVITE-M/Rec-mau-M-24/mfG1936.rec
LAB: pau k a0104 i0204 j     o0304 u0404 pau
REC: pau   t     a0103 i0203 h     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1938.lab vs ../HVITE-M/Rec-mau-M-24/mfG1938.rec
LAB: pau k e0104 i0204 z a0304 i0404 pau
REC: pau g e0104 N0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1939.lab vs ../HVITE-M/Rec-mau-M-24/mfG1939.rec
LAB: pau sh u0104 u0204 ry o0304 u0404 pau
REC: pau k  o0103 u0203 ry       o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1943.lab vs ../HVITE-M/Rec-mau-M-24/mfG1943.rec
LAB: pau m u0103 k                        o0203 u0303 pau
REC: pau         k u0104 ts u0204 r o0304 g     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1944.lab vs ../HVITE-M/Rec-mau-M-24/mfG1944.rec
LAB: pau sh o0104 u0204 j i0304 r  u0404 pau
REC: pau    h     o0103 k e0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1946.lab vs ../HVITE-M/Rec-mau-M-24/mfG1946.rec
LAB: pau m i0103 e0203 r u0303 pau
REC: pau h i0103 e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1947.lab vs ../HVITE-M/Rec-mau-M-24/mfG1947.rec
LAB: pau       m a0102 ch i0202 pau
REC: pau a0103 k i0203 r  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1948.lab vs ../HVITE-M/Rec-mau-M-24/mfG1948.rec
LAB: pau k         i0102 ch i0202 pau
REC: pau k o0103 t o0203 b  a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1949.lab vs ../HVITE-M/Rec-mau-M-24/mfG1949.rec
LAB: pau t a0104 i0204 s a0304 k u0404 pau
REC: pau n a0104 i0204 k a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1950.lab vs ../HVITE-M/Rec-mau-M-24/mfG1950.rec
LAB: pau y u0103 s u0203 g u0303 pau
REC: pau r i0103 z u0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1951.lab vs ../HVITE-M/Rec-mau-M-24/mfG1951.rec
LAB: pau k       u0105 i0205 ch i0305 g a0405   u0505 pau

REC: pau k i0105 r     i0205 h  i0305 r a0405 k u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1952.lab vs ../HVITE-M/Rec-mau-M-24/mfG1952.rec
LAB: pau sh u0102 t o0202 pau
REC: pau k  a0102 d o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1953.lab vs ../HVITE-M/Rec-mau-M-24/mfG1953.rec
LAB: pau   k     o0104 u0204 g a0304 k u0404 pau
REC: pau h i0104 r     o0204 g a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1954.lab vs ../HVITE-M/Rec-mau-M-24/mfG1954.rec
LAB: pau k o0104 u0204 s     a0304 i0404 pau
REC: pau   r     o0103 u0203 h     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1955.lab vs ../HVITE-M/Rec-mau-M-24/mfG1955.rec
LAB: pau ch i0104 r a0204 b     a0304 r     u0404 pau
REC: pau ch               a0103 w     a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1956.lab vs ../HVITE-M/Rec-mau-M-24/mfG1956.rec
LAB: pau ch i0103 k a0203 r a0303 pau
REC: pau          k a0102 r a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1958.lab vs ../HVITE-M/Rec-mau-M-24/mfG1958.rec
LAB: pau j u0104 N0204 sh i0304 N0404 pau
REC: pau   k     u0103 sh i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1959.lab vs ../HVITE-M/Rec-mau-M-24/mfG1959.rec
LAB: pau s a0103 g u0203 r u0303 pau
REC: pau k a0103 g i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1960.lab vs ../HVITE-M/Rec-mau-M-24/mfG1960.rec
LAB: pau d       a0103 sh i0203 N0303 pau
REC: pau d a0104 i0204 k  i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1962.lab vs ../HVITE-M/Rec-mau-M-24/mfG1962.rec
LAB: pau ts         u0102 b u0202 pau
REC: pau ts u0103 z u0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1964.lab vs ../HVITE-M/Rec-mau-M-24/mfG1964.rec
LAB: pau         s     a0102 y     a0202 pau
REC: pau g a0104 i0204 t     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1965.lab vs ../HVITE-M/Rec-mau-M-24/mfG1965.rec
LAB: pau m e0104 i0204 s a0304 k u0404 pau
REC: pau   n     e0103 s a0203 g e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1966.lab vs ../HVITE-M/Rec-mau-M-24/mfG1966.rec
LAB: pau h e0104 N0204 k     o0304 u0404 pau
REC: pau   g     e0103 N0203 g     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1968.lab vs ../HVITE-M/Rec-mau-M-24/mfG1968.rec
LAB: pau         ts    u0102 m     a0202 pau

REC: pau k o0104 N0204 n     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1969.lab vs ../HVITE-M/Rec-mau-M-24/mfG1969.rec
LAB: pau j i0104 q0204 s a0304 i0404 pau
REC: pau k a0104 i0204 s a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1970.lab vs ../HVITE-M/Rec-mau-M-24/mfG1970.rec
LAB: pau   y     a0102 k     u0202 pau
REC: pau t a0103 b     u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1971.lab vs ../HVITE-M/Rec-mau-M-24/mfG1971.rec
LAB: pau t e0103 h a0203 i0303 pau
REC: pau k e0103 h a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1972.lab vs ../HVITE-M/Rec-mau-M-24/mfG1972.rec
LAB: pau k       i0102 j     i0202 pau
REC: pau k i0103 k     i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1973.lab vs ../HVITE-M/Rec-mau-M-24/mfG1973.rec
LAB: pau r a0104 N0204 z a0304 ts u0404 pau
REC: pau s a0104 N0204 g a0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1976.lab vs ../HVITE-M/Rec-mau-M-24/mfG1976.rec
LAB: pau           o0103 sh i0203 m u0303 pau
REC: pau k u0104 r u0204 sh i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1979.lab vs ../HVITE-M/Rec-mau-M-24/mfG1979.rec
LAB: pau n a0103 m             a0203 e0303 pau
REC: pau k a0104 m a0204 e0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1983.lab vs ../HVITE-M/Rec-mau-M-24/mfG1983.rec
LAB: pau   m     a0102 ch i0202 pau
REC: pau s a0103 N0203 j  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1985.lab vs ../HVITE-M/Rec-mau-M-24/mfG1985.rec
LAB: pau          ts u0102 b o0202 pau
REC: pau sh i0103 n  o0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1987.lab vs ../HVITE-M/Rec-mau-M-24/mfG1987.rec
LAB: pau u0102 m         a0202 pau
REC: pau       m a0102 g o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1988.lab vs ../HVITE-M/Rec-mau-M-24/mfG1988.rec
LAB: pau b e0103 q0203 d o0303 pau
REC: pau g e0103 N0203 d o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1991.lab vs ../HVITE-M/Rec-mau-M-24/mfG1991.rec
LAB: pau h o0104 u0204 k o0304 k u0404 pau
REC: pau               d o0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1992.lab vs ../HVITE-M/Rec-mau-M-24/mfG1992.rec
LAB: pau h o0103 k o0203 r i0303 pau

REC: pau f u0103 t o0203   i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1994.lab vs ../HVITE-M/Rec-mau-M-24/mfG1994.rec
LAB: pau k e0104 N0204 t e0304 i0404 pau
REC: pau s e0104 N0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1995.lab vs ../HVITE-M/Rec-mau-M-24/mfG1995.rec
LAB: pau f u0104 m i0204 k i0304 r i0404 pau
REC: pau         m o0103 m i0203 j i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1997.lab vs ../HVITE-M/Rec-mau-M-24/mfG1997.rec
LAB: pau y u0103 r u0203 m u0303 pau
REC: pau j u0103   N0203 b i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG1998.lab vs ../HVITE-M/Rec-mau-M-24/mfG1998.rec
LAB: pau f u0104 q0204 k o0304 u0404 pau
REC: pau b a0104 N0204 g o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2000.lab vs ../HVITE-M/Rec-mau-M-24/mfG2000.rec
LAB: pau h e0104   i0204 b o0304 N0404 pau
REC: pau t e0104 n u0204 g u0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2001.lab vs ../HVITE-M/Rec-mau-M-24/mfG2001.rec
LAB: pau t  o0104 k u0204 t e0304 i0404 pau
REC: pau ts u0103 k         u0203 e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2003.lab vs ../HVITE-M/Rec-mau-M-24/mfG2003.rec
LAB: pau sh u0104 u0204 d a0304 N0404 pau
REC: pau ch u0104 u0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2004.lab vs ../HVITE-M/Rec-mau-M-24/mfG2004.rec
LAB: pau   i0104 q0204 t a0304 i0404 pau
REC: pau g u0104 N0204 t a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2005.lab vs ../HVITE-M/Rec-mau-M-24/mfG2005.rec
LAB: pau       k  i0103 b a0203 N0303 pau
REC: pau i0104 ch i0204 b a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2007.lab vs ../HVITE-M/Rec-mau-M-24/mfG2007.rec
LAB: pau t a0103 i0203 r         a0303 pau
REC: pau k a0103       r a0203 d a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2009.lab vs ../HVITE-M/Rec-mau-M-24/mfG2009.rec
LAB: pau       t a0103 b e0203 r u0303 pau
REC: pau u0104 k a0204 b e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2010.lab vs ../HVITE-M/Rec-mau-M-24/mfG2010.rec
LAB: pau   o0103 o0203 u0303 pau
REC: pau s o0102 t     o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2011.lab vs ../HVITE-M/Rec-mau-M-24/mfG2011.rec
LAB: pau         y o0102 s o0202 pau

REC: pau f u0103 z o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2013.lab vs ../HVITE-M/Rec-mau-M-24/mfG2013.rec
LAB: pau   h     a0102 k u0202 pau
REC: pau p a0103 N0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2014.lab vs ../HVITE-M/Rec-mau-M-24/mfG2014.rec
LAB: pau sh               o0102 u0202 pau
REC: pau sh i0103 n o0203 b     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2015.lab vs ../HVITE-M/Rec-mau-M-24/mfG2015.rec
LAB: pau t e0104 k i0204 s u0304 r u0404 pau
REC: pau         k e0103 z u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2017.lab vs ../HVITE-M/Rec-mau-M-24/mfG2017.rec
LAB: pau s a0104 k     u0204 h     i0304 N0404 pau
REC: pau   h     a0103 g     e0203 m     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2019.lab vs ../HVITE-M/Rec-mau-M-24/mfG2019.rec
LAB: pau k o0104 b o0204 r e0304 r u0404 pau
REC: pau k o0104 g o0204   e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2020.lab vs ../HVITE-M/Rec-mau-M-24/mfG2020.rec
LAB: pau               f u0102 i0202 pau
REC: pau k o0104 u0204 n i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2021.lab vs ../HVITE-M/Rec-mau-M-24/mfG2021.rec
LAB: pau                       n a0101 pau
REC: pau g a0104 i0204 k o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2022.lab vs ../HVITE-M/Rec-mau-M-24/mfG2022.rec
LAB: pau y u0103 r u0203 m u0303 pau
REC: pau t a0103 k u0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2023.lab vs ../HVITE-M/Rec-mau-M-24/mfG2023.rec
LAB: pau   s     a0102 k         i0202 pau
REC: pau h a0104 q0204 k i0304 r i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2025.lab vs ../HVITE-M/Rec-mau-M-24/mfG2025.rec
LAB: pau   a0103 N0203 i0303 pau
REC: pau k a0102 g     u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2026.lab vs ../HVITE-M/Rec-mau-M-24/mfG2026.rec
LAB: pau sh i0103 m e0203 s  u0303 pau
REC: pau ky o0103 z e0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2027.lab vs ../HVITE-M/Rec-mau-M-24/mfG2027.rec
LAB: pau k a0103 y o0203 u0303 pau
REC: pau s a0103 y o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2028.lab vs ../HVITE-M/Rec-mau-M-24/mfG2028.rec
LAB: pau ts u0104 t a0204 e0304 r u0404 pau

REC: pau k  i0104 t a0204 e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2029.lab vs ../HVITE-M/Rec-mau-M-24/mfG2029.rec
LAB: pau sh i0103         m e0203 i0303 pau
REC: pau m  i0103 j i0203 m       e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2030.lab vs ../HVITE-M/Rec-mau-M-24/mfG2030.rec
LAB: pau               b a0103 i0203 u0303 pau
REC: pau u0104 r i0204 d a0304 s     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2031.lab vs ../HVITE-M/Rec-mau-M-24/mfG2031.rec
LAB: pau   a0103 k i0203 r u0303 pau
REC: pau s a0103 k u0203 r a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2034.lab vs ../HVITE-M/Rec-mau-M-24/mfG2034.rec
LAB: pau         h e0102 y  a0202 pau
REC: pau k u0103 g a0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2035.lab vs ../HVITE-M/Rec-mau-M-24/mfG2035.rec
LAB: pau k a0104 k u0204 j i0304 ts u0404 pau
REC: pau         h a0103 j i0203 k  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2036.lab vs ../HVITE-M/Rec-mau-M-24/mfG2036.rec
LAB: pau k a0104 N0204 s o0304 k u0404 pau
REC: pau k       a0103 z o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2037.lab vs ../HVITE-M/Rec-mau-M-24/mfG2037.rec
LAB: pau       ts u0102 m         u0202 pau
REC: pau u0104 ts u0204 m u0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2038.lab vs ../HVITE-M/Rec-mau-M-24/mfG2038.rec
LAB: pau   i0104 k u0204 s e0304 i0404 pau
REC: pau k i0104   N0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2039.lab vs ../HVITE-M/Rec-mau-M-24/mfG2039.rec
LAB: pau         s a0104 b i0204 r e0304 r u0404 pau
REC: pau k u0105 t a0205 b i0305 r e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2040.lab vs ../HVITE-M/Rec-mau-M-24/mfG2040.rec
LAB: pau          k a0103 n a0203 g u0303 pau
REC: pau ts u0104 k a0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2041.lab vs ../HVITE-M/Rec-mau-M-24/mfG2041.rec
LAB: pau          m o0102 r       u0202 pau
REC: pau ts u0104 k u0204 r o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2043.lab vs ../HVITE-M/Rec-mau-M-24/mfG2043.rec
LAB: pau z u0106 u0206 z u0306 u0406 sh    i0506 i0606 pau
REC: pau                 d     o0103 u0203 k     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2045.lab vs ../HVITE-M/Rec-mau-M-24/mfG2045.rec
LAB: pau          s a0103 sh i0203 z u0303 pau

REC: pau sh i0104 t a0204 sh i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2046.lab vs ../HVITE-M/Rec-mau-M-24/mfG2046.rec
LAB: pau         ry o0102 u0202 pau
REC: pau m u0103 r  o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2048.lab vs ../HVITE-M/Rec-mau-M-24/mfG2048.rec
LAB: pau y  o0103 ky    o0203 u0303 pau
REC: pau ts u0103 u0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2049.lab vs ../HVITE-M/Rec-mau-M-24/mfG2049.rec
LAB: pau         k u0102 sh i0202 pau
REC: pau g i0103 j u0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2050.lab vs ../HVITE-M/Rec-mau-M-24/mfG2050.rec
LAB: pau m e0104 i0204 ry    o0304 u0404 pau
REC: pau   g     e0103 N0203 g     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2053.lab vs ../HVITE-M/Rec-mau-M-24/mfG2053.rec
LAB: pau k i0103 t e0203 i0303 pau
REC: pau         s e0102 i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2054.lab vs ../HVITE-M/Rec-mau-M-24/mfG2054.rec
LAB: pau           n     a0103 g e0203 r u0303 pau
REC: pau k i0105 r i0205 a0305 g e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2055.lab vs ../HVITE-M/Rec-mau-M-24/mfG2055.rec
LAB: pau k  i0103 k a0203 k u0303 pau
REC: pau sh i0103   a0203 g e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2056.lab vs ../HVITE-M/Rec-mau-M-24/mfG2056.rec
LAB: pau h e0103 N0203 j i0303 pau
REC: pau h o0103 N0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2057.lab vs ../HVITE-M/Rec-mau-M-24/mfG2057.rec
LAB: pau f u0103 sh i0203 N0303 pau
REC: pau k u0103 sh i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2058.lab vs ../HVITE-M/Rec-mau-M-24/mfG2058.rec
LAB: pau         a0102 k i0202 pau
REC: pau g a0103 q0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2059.lab vs ../HVITE-M/Rec-mau-M-24/mfG2059.rec
LAB: pau s o0104 u0204 z     o0304 u0404 pau
REC: pau   i0103 t     o0203 k     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2060.lab vs ../HVITE-M/Rec-mau-M-24/mfG2060.rec
LAB: pau         i0102 r o0202 pau
REC: pau t a0103 i0203 d o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2062.lab vs ../HVITE-M/Rec-mau-M-24/mfG2062.rec
LAB: pau e0104 N0204 g       e0304 k i0404 pau

REC: pau             g e0103 N0203 g o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2063.lab vs ../HVITE-M/Rec-mau-M-24/mfG2063.rec
LAB: pau       k  u0102 m         o0202 pau
REC: pau u0104 ts u0204 m u0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2064.lab vs ../HVITE-M/Rec-mau-M-24/mfG2064.rec
LAB: pau k i0103 s e0203 ts u0303 pau
REC: pau h i0103 k e0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2065.lab vs ../HVITE-M/Rec-mau-M-24/mfG2065.rec
LAB: pau    d     a0104 N0204 b o0304   u0404 pau
REC: pau ch i0104 k     a0204 g o0304 r o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2066.lab vs ../HVITE-M/Rec-mau-M-24/mfG2066.rec
LAB: pau b u0104 N0204 sh    o0304 u0404 pau
REC: pau   k     o0103 N0203 ky    o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2067.lab vs ../HVITE-M/Rec-mau-M-24/mfG2067.rec
LAB: pau h o0102 j o0202 pau
REC: pau s o0102 k o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2069.lab vs ../HVITE-M/Rec-mau-M-24/mfG2069.rec
LAB: pau m e0103 sh i0203 t a0303 pau
REC: pau   i0103 k  u0203 r a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2070.lab vs ../HVITE-M/Rec-mau-M-24/mfG2070.rec
LAB: pau   y     o0102 k         u0202 pau
REC: pau t e0104 o0204 k u0304 r e0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2071.lab vs ../HVITE-M/Rec-mau-M-24/mfG2071.rec
LAB: pau d a0102 k u0202 pau
REC: pau m a0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2072.lab vs ../HVITE-M/Rec-mau-M-24/mfG2072.rec
LAB: pau g o0103 u0203 k a0303 pau
REC: pau t o0103 o0203 k a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2073.lab vs ../HVITE-M/Rec-mau-M-24/mfG2073.rec
LAB: pau         k u0104 N0204 sh o0304 u0404 pau
REC: pau m u0103 k                o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2074.lab vs ../HVITE-M/Rec-mau-M-24/mfG2074.rec
LAB: pau n a0103 o0203 s u0303 pau
REC: pau t a0103 o0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2075.lab vs ../HVITE-M/Rec-mau-M-24/mfG2075.rec
LAB: pau h e0103 k o0203 m u0303 pau
REC: pau         g o0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2076.lab vs ../HVITE-M/Rec-mau-M-24/mfG2076.rec
LAB: pau   a0103 k e0203 r u0303 pau

REC: pau t a0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2077.lab vs ../HVITE-M/Rec-mau-M-24/mfG2077.rec
LAB: pau h i0103 s o0203 m u0303 pau
REC: pau h e0103 k o0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2078.lab vs ../HVITE-M/Rec-mau-M-24/mfG2078.rec
LAB: pau m i0103 m a0203   u0303 pau
REC: pau m i0103 g a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2081.lab vs ../HVITE-M/Rec-mau-M-24/mfG2081.rec
LAB: pau sh o0104 k u0204 n i0304 N0404 pau
REC: pau sh o0104 k u0204   i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2082.lab vs ../HVITE-M/Rec-mau-M-24/mfG2082.rec
LAB: pau   m     o0102 j       i0202 pau
REC: pau r o0104 u0204 j i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2083.lab vs ../HVITE-M/Rec-mau-M-24/mfG2083.rec
LAB: pau r o0104 q0204 k a0304   a0404 pau
REC: pau m o0104 q0204 p a0304 r a0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2084.lab vs ../HVITE-M/Rec-mau-M-24/mfG2084.rec
LAB: pau h a0102 k         o0202 pau
REC: pau b a0103 k u0203 r o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2087.lab vs ../HVITE-M/Rec-mau-M-24/mfG2087.rec
LAB: pau k  a0104 k u0204 j i0304 ts u0404 pau
REC: pau ch a0104 k u0204 j i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2089.lab vs ../HVITE-M/Rec-mau-M-24/mfG2089.rec
LAB: pau ky o0104 u0204 d o0304 u0404 pau
REC: pau h  o0104 u0204 d o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2090.lab vs ../HVITE-M/Rec-mau-M-24/mfG2090.rec
LAB: pau               h o0102 k     a0202 pau
REC: pau k o0104 u0204 h       a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2092.lab vs ../HVITE-M/Rec-mau-M-24/mfG2092.rec
LAB: pau   o0103 n a0203 k a0303 pau
REC: pau m a0103 t a0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2093.lab vs ../HVITE-M/Rec-mau-M-24/mfG2093.rec
LAB: pau h o0104 h     o0204 e0304 m u0404 pau
REC: pau   m     o0103 r     e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2094.lab vs ../HVITE-M/Rec-mau-M-24/mfG2094.rec
LAB: pau   n     i0103 j u0203 u0303 pau
REC: pau s u0104 i0204 j u0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2095.lab vs ../HVITE-M/Rec-mau-M-24/mfG2095.rec
LAB: pau k o0104 u0204 u0304 N0404 pau

REC: pau k i0103 h     o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2096.lab vs ../HVITE-M/Rec-mau-M-24/mfG2096.rec
LAB: pau   i0103 b a0203 r u0303 pau
REC: pau k i0103 m a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2097.lab vs ../HVITE-M/Rec-mau-M-24/mfG2097.rec
LAB: pau h i0105 q0205 k     a0305 k e0405 r u0505 pau
REC: pau k i0105 r     i0205 a0305 g e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2099.lab vs ../HVITE-M/Rec-mau-M-24/mfG2099.rec
LAB: pau t o0104 q0204 ky    u0304 u0404 pau
REC: pau   b     o0103 o0203 r     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2100.lab vs ../HVITE-M/Rec-mau-M-24/mfG2100.rec
LAB: pau h o0104 u0204 h     o0304 u0404 pau
REC: pau   i0103 t     o0203 k     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2101.lab vs ../HVITE-M/Rec-mau-M-24/mfG2101.rec
LAB: pau s o0103 u0203 g o0303 pau
REC: pau   k     o0102 r o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2102.lab vs ../HVITE-M/Rec-mau-M-24/mfG2102.rec
LAB: pau y  u0103 r u0203 m u0303 pau
REC: pau sh u0103   u0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2103.lab vs ../HVITE-M/Rec-mau-M-24/mfG2103.rec
LAB: pau                  t o0102 sh i0202 pau
REC: pau m o0104 ch i0204 d a0304 s  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2105.lab vs ../HVITE-M/Rec-mau-M-24/mfG2105.rec
LAB: pau   e0103 N0203 k i0303 pau
REC: pau t e0103 N0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2106.lab vs ../HVITE-M/Rec-mau-M-24/mfG2106.rec
LAB: pau       j     i0102 m     i0202 pau
REC: pau o0104 i0204 sh    i0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2107.lab vs ../HVITE-M/Rec-mau-M-24/mfG2107.rec
LAB: pau   s     e0103 N0203 r     o0303 pau
REC: pau d e0104 N0204 t     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2108.lab vs ../HVITE-M/Rec-mau-M-24/mfG2108.rec
LAB: pau f u0103 s e0203 i0303 pau
REC: pau f         u0102 e0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2110.lab vs ../HVITE-M/Rec-mau-M-24/mfG2110.rec
LAB: pau               m o0102 r u0202 pau
REC: pau i0104 t a0204 z u0304 r a0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2111.lab vs ../HVITE-M/Rec-mau-M-24/mfG2111.rec
LAB: pau n i0103 k     u0203 i0303 pau

REC: pau g i0103 N0203 m     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2112.lab vs ../HVITE-M/Rec-mau-M-24/mfG2112.rec
LAB: pau         k a0102 g e0202 pau
REC: pau t a0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2113.lab vs ../HVITE-M/Rec-mau-M-24/mfG2113.rec
LAB: pau k e0104 i0204 r e0304 k     i0404 pau
REC: pau t a0104 i0204         o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2115.lab vs ../HVITE-M/Rec-mau-M-24/mfG2115.rec
LAB: pau         k a0103 y o0203 u0303 pau
REC: pau t a0104 d a0204 y o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2116.lab vs ../HVITE-M/Rec-mau-M-24/mfG2116.rec
LAB: pau j u0106 u0206 i0306 ch i0406 g a0506 ts u0606 pau
REC: pau                     k  u0103 g a0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2117.lab vs ../HVITE-M/Rec-mau-M-24/mfG2117.rec
LAB: pau   h     i0102 b i0202 pau
REC: pau k i0103 i0203 r o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2119.lab vs ../HVITE-M/Rec-mau-M-24/mfG2119.rec
LAB: pau    s     u0103 g a0203 t a0303 pau
REC: pau ts u0104 u0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2120.lab vs ../HVITE-M/Rec-mau-M-24/mfG2120.rec
LAB: pau   a0104 s a0204   q0304 t e0404 pau
REC: pau m i0104 k a0204 k e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2121.lab vs ../HVITE-M/Rec-mau-M-24/mfG2121.rec
LAB: pau   a0102 k a0202 pau
REC: pau g a0102 r a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2127.lab vs ../HVITE-M/Rec-mau-M-24/mfG2127.rec
LAB: pau          t a0102 r u0202 pau
REC: pau sh i0103 b a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2128.lab vs ../HVITE-M/Rec-mau-M-24/mfG2128.rec
LAB: pau ry o0103 k               a0203 N0303 pau
REC: pau          k a0104 i0204 r a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2129.lab vs ../HVITE-M/Rec-mau-M-24/mfG2129.rec
LAB: pau                 u0102 t a0202 pau
REC: pau u0104 r a0204 g i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2130.lab vs ../HVITE-M/Rec-mau-M-24/mfG2130.rec
LAB: pau sh i0104 ts u0204 gy    o0304 u0404 pau
REC: pau          k  o0103 u0203 ry    o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2131.lab vs ../HVITE-M/Rec-mau-M-24/mfG2131.rec
LAB: pau h  a0104 r i0204 a0304 i0404 pau

REC: pau ts u0104 r i0204 a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2133.lab vs ../HVITE-M/Rec-mau-M-24/mfG2133.rec
LAB: pau t o0103 n a0203 r i0303 pau
REC: pau s u0103 m a0203   i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2134.lab vs ../HVITE-M/Rec-mau-M-24/mfG2134.rec
LAB: pau m i0106 s e0206 b i0306 r     a0406 k a0506 s  u0606 pau
REC: pau         s               e0104 i0204 k a0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2135.lab vs ../HVITE-M/Rec-mau-M-24/mfG2135.rec
LAB: pau s  a0104 b i0204 ts u0304 k u0404 pau
REC: pau ts u0104 r a0204 n  u0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2136.lab vs ../HVITE-M/Rec-mau-M-24/mfG2136.rec
LAB: pau                 m a0102 d a0202 pau
REC: pau t a0103 t a0203 m         i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2137.lab vs ../HVITE-M/Rec-mau-M-24/mfG2137.rec
LAB: pau m a0103 s u0203   i0303 pau
REC: pau k a0103 s u0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2139.lab vs ../HVITE-M/Rec-mau-M-24/mfG2139.rec
LAB: pau m a0103 z a0203 r u0303 pau
REC: pau k a0103 z a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2142.lab vs ../HVITE-M/Rec-mau-M-24/mfG2142.rec
LAB: pau y  o0102 k o0202 pau
REC: pau sh i0102 r o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2143.lab vs ../HVITE-M/Rec-mau-M-24/mfG2143.rec
LAB: pau         n a0103 m e0203 r u0303 pau
REC: pau k u0104 r a0204 b e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2145.lab vs ../HVITE-M/Rec-mau-M-24/mfG2145.rec
LAB: pau       j i0103 k i0203 n i0303 pau
REC: pau o0103 j         i0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2146.lab vs ../HVITE-M/Rec-mau-M-24/mfG2146.rec
LAB: pau m  i0103 k a0203 N0303 pau
REC: pau sh i0103   a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2147.lab vs ../HVITE-M/Rec-mau-M-24/mfG2147.rec
LAB: pau s a0103 k a0203 r i0303 pau
REC: pau k a0103 r i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2148.lab vs ../HVITE-M/Rec-mau-M-24/mfG2148.rec
LAB: pau y o0103 k o0203 k u0303 pau
REC: pau         r o0102 k u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2149.lab vs ../HVITE-M/Rec-mau-M-24/mfG2149.rec
LAB: pau s e0104 i0204 j i0304 ts u0404 pau

REC: pau h e0104 i0204 j i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2150.lab vs ../HVITE-M/Rec-mau-M-24/mfG2150.rec
LAB: pau                   o0102 b i0202 pau
REC: pau t o0104 g a0204 m e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2151.lab vs ../HVITE-M/Rec-mau-M-24/mfG2151.rec
LAB: pau            sh    i0102 N0202 pau
REC: pau ch i0103 j i0203 m     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2152.lab vs ../HVITE-M/Rec-mau-M-24/mfG2152.rec
LAB: pau s a0104 i0204 g a0304 i0404 pau
REC: pau t a0104 i0204 g a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2153.lab vs ../HVITE-M/Rec-mau-M-24/mfG2153.rec
LAB: pau t o0102 k               u0202 pau
REC: pau         k u0103 r u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2156.lab vs ../HVITE-M/Rec-mau-M-24/mfG2156.rec
LAB: pau    i0104 k i0204 n         a0304 r i0404 pau
REC: pau sh i0104         n a0204 b i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2157.lab vs ../HVITE-M/Rec-mau-M-24/mfG2157.rec
LAB: pau         k     i0103 sh    o0203 u0303 pau
REC: pau s u0104 i0204 ch    o0304 k     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2158.lab vs ../HVITE-M/Rec-mau-M-24/mfG2158.rec
LAB: pau b o0102 ch i0202 pau
REC: pau t o0102 k  i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2159.lab vs ../HVITE-M/Rec-mau-M-24/mfG2159.rec
LAB: pau    o0103 k i0203 r u0303 pau
REC: pau ch i0103 g i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2160.lab vs ../HVITE-M/Rec-mau-M-24/mfG2160.rec
LAB: pau   a0104 q0204 p a0304 k u0404 pau
REC: pau h a0104 N0204 k a0304   N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2161.lab vs ../HVITE-M/Rec-mau-M-24/mfG2161.rec
LAB: pau k a0103 ch i0203 k u0303 pau
REC: pau s a0103 g  u0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2163.lab vs ../HVITE-M/Rec-mau-M-24/mfG2163.rec
LAB: pau s a0103 k a0203 r i0303 pau
REC: pau k a0103 g a0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2165.lab vs ../HVITE-M/Rec-mau-M-24/mfG2165.rec
LAB: pau sh i0103 ts         u0203 d o0303 pau
REC: pau k  u0104 ts u0204 r o0304 g u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2166.lab vs ../HVITE-M/Rec-mau-M-24/mfG2166.rec
LAB: pau          s u0101 pau

REC: pau ts u0102 g u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2167.lab vs ../HVITE-M/Rec-mau-M-24/mfG2167.rec
LAB: pau   h     a0102 ch    i0202 pau
REC: pau k a0103 r     u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2168.lab vs ../HVITE-M/Rec-mau-M-24/mfG2168.rec
LAB: pau sh u0104 k     u0204 d a0304 i0404 pau
REC: pau    h     e0104 i0204 g a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2169.lab vs ../HVITE-M/Rec-mau-M-24/mfG2169.rec
LAB: pau j u0103 k e0203 N0303 pau
REC: pau k i0103 k e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2170.lab vs ../HVITE-M/Rec-mau-M-24/mfG2170.rec
LAB: pau s o0104 k       u0204 t a0304 ts u0404 pau
REC: pau         k a0104 i0204 k a0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2171.lab vs ../HVITE-M/Rec-mau-M-24/mfG2171.rec
LAB: pau h a0104 k i0204 d a0304 s u0404 pau
REC: pau         n a0103 r a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2172.lab vs ../HVITE-M/Rec-mau-M-24/mfG2172.rec
LAB: pau   o0103 d o0203 r i0303 pau
REC: pau m i0103 b u0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2174.lab vs ../HVITE-M/Rec-mau-M-24/mfG2174.rec
LAB: pau k o0102 u0202 pau
REC: pau m o0102 u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2175.lab vs ../HVITE-M/Rec-mau-M-24/mfG2175.rec
LAB: pau m a0103 t o0203 m o0303 pau
REC: pau k a0103 k o0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2176.lab vs ../HVITE-M/Rec-mau-M-24/mfG2176.rec
LAB: pau   i0104 q0204 t a0304 i0404 pau
REC: pau g u0104 N0204 t a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2177.lab vs ../HVITE-M/Rec-mau-M-24/mfG2177.rec
LAB: pau    sh    o0103 r u0203 i0303 pau
REC: pau ky u0103 u0203 r       i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2178.lab vs ../HVITE-M/Rec-mau-M-24/mfG2178.rec
LAB: pau         y a0103 t a0203 r a0303 pau
REC: pau k u0104 s a0204 b a0304 n a0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2179.lab vs ../HVITE-M/Rec-mau-M-24/mfG2179.rec
LAB: pau   k     i0103 g e0203 N0303 pau
REC: pau h o0104 u0204 g e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2180.lab vs ../HVITE-M/Rec-mau-M-24/mfG2180.rec
LAB: pau t e0104 q0204 k     i0304 N0404 pau

REC: pau t e0104       N0204 i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2181.lab vs ../HVITE-M/Rec-mau-M-24/mfG2181.rec
LAB: pau sh i0103 z u0203 k                 a0303 pau
REC: pau          f u0104 k u0204 r a0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2182.lab vs ../HVITE-M/Rec-mau-M-24/mfG2182.rec
LAB: pau m e0104 N0204 s     e0304 k     i0404 pau
REC: pau         sh    i0103 b     u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2183.lab vs ../HVITE-M/Rec-mau-M-24/mfG2183.rec
LAB: pau h a0104 r i0204 a0304 i0404 pau
REC: pau k a0103 r       a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2184.lab vs ../HVITE-M/Rec-mau-M-24/mfG2184.rec
LAB: pau k e0104 q0204 t e0304 N0404 pau
REC: pau   m     u0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2185.lab vs ../HVITE-M/Rec-mau-M-24/mfG2185.rec
LAB: pau d o0104 k u0204 r i0304 ts u0404 pau
REC: pau k o0104 k u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2186.lab vs ../HVITE-M/Rec-mau-M-24/mfG2186.rec
LAB: pau   sh    i0103 ts u0203 d o0303 pau
REC: pau t a0104 q0204 s  u0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2187.lab vs ../HVITE-M/Rec-mau-M-24/mfG2187.rec
LAB: pau m a0104 d o0204 g         u0304 ch i0404 pau
REC: pau                 g a0103 k u0203 h  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2188.lab vs ../HVITE-M/Rec-mau-M-24/mfG2188.rec
LAB: pau       t o0104 d o0204 k e0304 r u0404 pau
REC: pau u0105 n u0205 b o0305 r e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2189.lab vs ../HVITE-M/Rec-mau-M-24/mfG2189.rec
LAB: pau t o0104       r i0204 k e0304 s  u0404 pau
REC: pau n o0104 u0204 r         i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2190.lab vs ../HVITE-M/Rec-mau-M-24/mfG2190.rec
LAB: pau         m a0102 m e0202 pau
REC: pau t a0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2192.lab vs ../HVITE-M/Rec-mau-M-24/mfG2192.rec
LAB: pau h i0105 k         i0205 u0305 k e0405 r u0505 pau
REC: pau         k u0105 r i0205 a0305 g e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2194.lab vs ../HVITE-M/Rec-mau-M-24/mfG2194.rec
LAB: pau          b     u0103 sh    o0203 u0303 pau
REC: pau u0104 ch i0204 k     o0304 m     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2195.lab vs ../HVITE-M/Rec-mau-M-24/mfG2195.rec
LAB: pau         u0103 d o0203 N0303 pau

REC: pau d o0104 N0204 d o0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2196.lab vs ../HVITE-M/Rec-mau-M-24/mfG2196.rec
LAB: pau y a0103 k o0203 u0303 pau
REC: pau h a0102 k       o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2198.lab vs ../HVITE-M/Rec-mau-M-24/mfG2198.rec
LAB: pau t a0104 i0204 t e0304   i0404 pau
REC: pau t a0104 o0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2200.lab vs ../HVITE-M/Rec-mau-M-24/mfG2200.rec
LAB: pau h a0104 ch i0204 g a0304 ts u0404 pau
REC: pau h a0104 k  i0204 d a0304 s  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2202.lab vs ../HVITE-M/Rec-mau-M-24/mfG2202.rec
LAB: pau          r i0103 k o0203 u0303 pau
REC: pau sh i0102 r               o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2203.lab vs ../HVITE-M/Rec-mau-M-24/mfG2203.rec
LAB: pau f u0102 d         a0202 pau
REC: pau k u0103 d a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2205.lab vs ../HVITE-M/Rec-mau-M-24/mfG2205.rec
LAB: pau   w     a0102 r     u0202 pau
REC: pau t a0103 b     u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2208.lab vs ../HVITE-M/Rec-mau-M-24/mfG2208.rec
LAB: pau       m a0103 w a0203 r i0303 pau
REC: pau i0104 t a0204 w a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2209.lab vs ../HVITE-M/Rec-mau-M-24/mfG2209.rec
LAB: pau         i0103 t e0203 N0303 pau
REC: pau b o0104 u0204 k e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2210.lab vs ../HVITE-M/Rec-mau-M-24/mfG2210.rec
LAB: pau n         i0102 r u0202 pau
REC: pau n i0103 g a0203 t e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2211.lab vs ../HVITE-M/Rec-mau-M-24/mfG2211.rec
LAB: pau k u0103 s a0203 r i0303 pau
REC: pau s u0103 p a0203   i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2212.lab vs ../HVITE-M/Rec-mau-M-24/mfG2212.rec
LAB: pau k o0103 n     o0203 m     u0303 pau
REC: pau r o0104 N0204 b     u0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2213.lab vs ../HVITE-M/Rec-mau-M-24/mfG2213.rec
LAB: pau h o0105 k o0205 r o0305 b i0405 r u0505 pau
REC: pau k o0105 k o0205 r o0305 m i0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2214.lab vs ../HVITE-M/Rec-mau-M-24/mfG2214.rec
LAB: pau s e0104 q0204 k e0304 N0404 pau

REC: pau   m     u0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2215.lab vs ../HVITE-M/Rec-mau-M-24/mfG2215.rec
LAB: pau            ch    o0103 q0203 t o0303 pau
REC: pau k u0104 ts u0204 r     o0304 g u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2217.lab vs ../HVITE-M/Rec-mau-M-24/mfG2217.rec
LAB: pau r e0102 ts u0202 pau
REC: pau k u0102 z  u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2218.lab vs ../HVITE-M/Rec-mau-M-24/mfG2218.rec
LAB: pau m u0103 k             o0203 u0303 pau
REC: pau         k o0103 u0203 h     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2220.lab vs ../HVITE-M/Rec-mau-M-24/mfG2220.rec
LAB: pau    t     o0102 sh    i0202 pau
REC: pau ch o0103 k     i0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2223.lab vs ../HVITE-M/Rec-mau-M-24/mfG2223.rec
LAB: pau               k i0102 s o0202 pau
REC: pau a0104 z u0204 k a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2224.lab vs ../HVITE-M/Rec-mau-M-24/mfG2224.rec
LAB: pau k o0104 k u0204 h     a0304 k     u0404 pau
REC: pau         b a0104 N0204 g     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2226.lab vs ../HVITE-M/Rec-mau-M-24/mfG2226.rec
LAB: pau         sh u0102 sh i0202 pau
REC: pau k e0103 sh i0203 k  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2227.lab vs ../HVITE-M/Rec-mau-M-24/mfG2227.rec
LAB: pau ts u0102   i0202 pau
REC: pau ts u0102 g u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2228.lab vs ../HVITE-M/Rec-mau-M-24/mfG2228.rec
LAB: pau   b     o0103 o0203 r u0303 pau
REC: pau k u0103 r     o0203   u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2229.lab vs ../HVITE-M/Rec-mau-M-24/mfG2229.rec
LAB: pau y u0104 u0204 k a0304 N0404 pau
REC: pau k u0104 u0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2231.lab vs ../HVITE-M/Rec-mau-M-24/mfG2231.rec
LAB: pau         k         e0101 pau
REC: pau h e0103 k o0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2232.lab vs ../HVITE-M/Rec-mau-M-24/mfG2232.rec
LAB: pau h     o0103 s     o0203 i0303 pau
REC: pau u0104 r     u0204 o0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2233.lab vs ../HVITE-M/Rec-mau-M-24/mfG2233.rec
LAB: pau k e0103 sh    i0203 k       i0303 pau

REC: pau k       e0104 i0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2234.lab vs ../HVITE-M/Rec-mau-M-24/mfG2234.rec
LAB: pau z a0104   i0204 g a0304 k u0404 pau
REC: pau k a0104 t a0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2235.lab vs ../HVITE-M/Rec-mau-M-24/mfG2235.rec
LAB: pau         m u0102 sh i0202 pau
REC: pau k u0103 j i0203 k  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2237.lab vs ../HVITE-M/Rec-mau-M-24/mfG2237.rec
LAB: pau k a0103 t e0203   i0303 pau
REC: pau s a0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2238.lab vs ../HVITE-M/Rec-mau-M-24/mfG2238.rec
LAB: pau                 n a0103 g o0203 r i0303 pau
REC: pau t a0104 z u0204 n         e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2239.lab vs ../HVITE-M/Rec-mau-M-24/mfG2239.rec
LAB: pau h a0103 r a0203   u0303 pau
REC: pau d a0103 r a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2240.lab vs ../HVITE-M/Rec-mau-M-24/mfG2240.rec
LAB: pau               k a0102 j i0202 pau
REC: pau g a0103 q0203 k         i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2241.lab vs ../HVITE-M/Rec-mau-M-24/mfG2241.rec
LAB: pau sh i0103 b o0203 m u0303 pau
REC: pau sh i0103 k o0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2242.lab vs ../HVITE-M/Rec-mau-M-24/mfG2242.rec
LAB: pau s u0103 u0203 p u0303 pau
REC: pau   ts    u0102 r u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2243.lab vs ../HVITE-M/Rec-mau-M-24/mfG2243.rec
LAB: pau    w     a0103 k a0203 i0303 pau
REC: pau ch u0104 u0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2244.lab vs ../HVITE-M/Rec-mau-M-24/mfG2244.rec
LAB: pau sh i0104 i0204 r e0304 r u0404 pau
REC: pau    k     i0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2245.lab vs ../HVITE-M/Rec-mau-M-24/mfG2245.rec
LAB: pau f u0104 q0204 k             o0304 u0404 pau
REC: pau               k a0103 N0203 g     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2246.lab vs ../HVITE-M/Rec-mau-M-24/mfG2246.rec
LAB: pau           f     u0102 sh    i0202 pau
REC: pau k o0104 k u0204 s     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2247.lab vs ../HVITE-M/Rec-mau-M-24/mfG2247.rec
LAB: pau d o0104 u0204 d o0304 u0404 pau

REC: pau b o0104 u0204 d o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2248.lab vs ../HVITE-M/Rec-mau-M-24/mfG2248.rec
LAB: pau k     a0103 i0203 g         i0303 pau
REC: pau u0104 r     a0204 g i0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2249.lab vs ../HVITE-M/Rec-mau-M-24/mfG2249.rec
LAB: pau         t a0102 n a0202 pau
REC: pau d a0103 r a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2255.lab vs ../HVITE-M/Rec-mau-M-24/mfG2255.rec
LAB: pau j o0104 u0204 n e0304 ts u0404 pau
REC: pau k o0104 N0204 g e0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2256.lab vs ../HVITE-M/Rec-mau-M-24/mfG2256.rec
LAB: pau k i0104 N0204 k o0304 u0404 pau
REC: pau r e0104 N0204 g o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2257.lab vs ../HVITE-M/Rec-mau-M-24/mfG2257.rec
LAB: pau sh u0104 u0204 s a0304 i0404 pau
REC: pau sh u0104 u0204 k a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2258.lab vs ../HVITE-M/Rec-mau-M-24/mfG2258.rec
LAB: pau h a0104 s a0204 m a0304 r u0404 pau
REC: pau k a0104 t a0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2259.lab vs ../HVITE-M/Rec-mau-M-24/mfG2259.rec
LAB: pau j i0103 d     o0203 u0303 pau
REC: pau k i0103 q0203 t     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2260.lab vs ../HVITE-M/Rec-mau-M-24/mfG2260.rec
LAB: pau f u0103 m e0203   i0303 pau
REC: pau h a0103 g e0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2262.lab vs ../HVITE-M/Rec-mau-M-24/mfG2262.rec
LAB: pau   i0103   q0203 k a0303 pau
REC: pau h i0103 r u0203 m a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2263.lab vs ../HVITE-M/Rec-mau-M-24/mfG2263.rec
LAB: pau   ch    u0104 u0204 r i0304 ts u0404 pau
REC: pau d o0104 k     u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2268.lab vs ../HVITE-M/Rec-mau-M-24/mfG2268.rec
LAB: pau                         o0102 i0202 pau
REC: pau t o0104 b i0204 k o0304 m     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2269.lab vs ../HVITE-M/Rec-mau-M-24/mfG2269.rec
LAB: pau   o0103 s     o0203 u0303 pau
REC: pau k o0104 u0204 u0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2270.lab vs ../HVITE-M/Rec-mau-M-24/mfG2270.rec
LAB: pau n a0104 i0204 y o0304 u0404 pau

REC: pau d a0104 i0204 y o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2272.lab vs ../HVITE-M/Rec-mau-M-24/mfG2272.rec
LAB: pau ry a0104 k     u0204 s u0304 r  u0404 pau
REC: pau    m     e0104 i0204 b u0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2274.lab vs ../HVITE-M/Rec-mau-M-24/mfG2274.rec
LAB: pau t a0103   e0203 r u0303 pau
REC: pau k a0103 g e0203   N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2275.lab vs ../HVITE-M/Rec-mau-M-24/mfG2275.rec
LAB: pau          j       u0102 u0202 pau
REC: pau ch i0103 j i0203 m     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2276.lab vs ../HVITE-M/Rec-mau-M-24/mfG2276.rec
LAB: pau k a0104 k u0204 sh i0304 N0404 pau
REC: pau k a0104   N0204 sh i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2277.lab vs ../HVITE-M/Rec-mau-M-24/mfG2277.rec
LAB: pau k e0104 N0204 k     a0304 i0404 pau
REC: pau         g     u0103 a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2279.lab vs ../HVITE-M/Rec-mau-M-24/mfG2279.rec
LAB: pau t a0104 d a0204 sh    i0304 i0404 pau
REC: pau         g a0103 q0203 k     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2280.lab vs ../HVITE-M/Rec-mau-M-24/mfG2280.rec
LAB: pau   a0104 b a0204 r e0304 r u0404 pau
REC: pau n a0104 g a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2282.lab vs ../HVITE-M/Rec-mau-M-24/mfG2282.rec
LAB: pau   a0104 i0204 ky    o0304 u0404 pau
REC: pau h a0103 m     o0203 n     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2283.lab vs ../HVITE-M/Rec-mau-M-24/mfG2283.rec
LAB: pau a0103 m     a0203 r       i0303 pau
REC: pau g     a0104 N0204 r a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2285.lab vs ../HVITE-M/Rec-mau-M-24/mfG2285.rec
LAB: pau e0105 ch i0205 k e0305 q0405 t o0505 pau
REC: pau       r  o0104 k e0204 q0304 t o0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2286.lab vs ../HVITE-M/Rec-mau-M-24/mfG2286.rec
LAB: pau sh o0106 u0206 g     a0306 q0406 k o0506 u0606 pau
REC: pau sh       o0104 u0204 d     a0304 k       u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2288.lab vs ../HVITE-M/Rec-mau-M-24/mfG2288.rec
LAB: pau   o0103 e0203 r u0303 pau
REC: pau s o0103 e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2289.lab vs ../HVITE-M/Rec-mau-M-24/mfG2289.rec
LAB: pau k u0104 r     u0204 sh i0304 m u0404 pau

REC: pau k       o0104 u0204 sh i0304   N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2290.lab vs ../HVITE-M/Rec-mau-M-24/mfG2290.rec
LAB: pau               f u0103 ts u0203 u0303 pau
REC: pau d o0104 u0204 b u0304 ts       u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2291.lab vs ../HVITE-M/Rec-mau-M-24/mfG2291.rec
LAB: pau   i0102 r o0202 pau
REC: pau m i0102 z o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2292.lab vs ../HVITE-M/Rec-mau-M-24/mfG2292.rec
LAB: pau sh o0103 y u0203 u0303 pau
REC: pau ts u0102 y       u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2293.lab vs ../HVITE-M/Rec-mau-M-24/mfG2293.rec
LAB: pau k u0103 k  u0203 r u0303 pau
REC: pau         ts u0102 r u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2294.lab vs ../HVITE-M/Rec-mau-M-24/mfG2294.rec
LAB: pau k i0103 k a0203 i0303 pau
REC: pau j i0103 k a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2295.lab vs ../HVITE-M/Rec-mau-M-24/mfG2295.rec
LAB: pau b e0104 N0204 sh o0304 u0404 pau
REC: pau k e0104 N0204 t  o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2296.lab vs ../HVITE-M/Rec-mau-M-24/mfG2296.rec
LAB: pau sh i0104 i0204 z u0304 N0404 pau
REC: pau    g     e0103 j u0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2297.lab vs ../HVITE-M/Rec-mau-M-24/mfG2297.rec
LAB: pau s o0103 m     u0203 k     u0303 pau
REC: pau r o0104 N0204 b     u0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2298.lab vs ../HVITE-M/Rec-mau-M-24/mfG2298.rec
LAB: pau d o0104 k u0204 r i0304 ts u0404 pau
REC: pau n o0104   u0204 r i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2299.lab vs ../HVITE-M/Rec-mau-M-24/mfG2299.rec
LAB: pau sh i0103 q0203 k i0303 pau
REC: pau k  o0103 u0203 g i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2300.lab vs ../HVITE-M/Rec-mau-M-24/mfG2300.rec
LAB: pau s a0103 k u0203 sh         a0303 pau
REC: pau h a0103         sh i0203 g o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2302.lab vs ../HVITE-M/Rec-mau-M-24/mfG2302.rec
LAB: pau       k a0102 t a0202 pau
REC: pau u0103 w a0203 s a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2303.lab vs ../HVITE-M/Rec-mau-M-24/mfG2303.rec
LAB: pau y o0103 u0203 k  i0303 pau

REC: pau y o0103 u0203 ch i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2304.lab vs ../HVITE-M/Rec-mau-M-24/mfG2304.rec
LAB: pau s e0103 N0203 r o0303 pau
REC: pau z e0103 N0203 g o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2305.lab vs ../HVITE-M/Rec-mau-M-24/mfG2305.rec
LAB: pau   i0102 r  o0202 pau
REC: pau b o0102 ch i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2306.lab vs ../HVITE-M/Rec-mau-M-24/mfG2306.rec
LAB: pau t o0103   u0203 sh o0303 pau
REC: pau s o0103 k u0203 d  o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2307.lab vs ../HVITE-M/Rec-mau-M-24/mfG2307.rec
LAB: pau m a0103 t     a0203 g     u0303 pau
REC: pau d a0104 N0204 d     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2309.lab vs ../HVITE-M/Rec-mau-M-24/mfG2309.rec
LAB: pau   i0103 n o0203 r u0303 pau
REC: pau m i0103 m o0203 t o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2310.lab vs ../HVITE-M/Rec-mau-M-24/mfG2310.rec
LAB: pau m i0103 r e0203   N0303 pau
REC: pau k i0103 m e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2311.lab vs ../HVITE-M/Rec-mau-M-24/mfG2311.rec
LAB: pau s a0104 N0204 s e0304 i0404 pau
REC: pau d a0104 N0204 t e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2312.lab vs ../HVITE-M/Rec-mau-M-24/mfG2312.rec
LAB: pau f u0103 s e0203 r u0303 pau
REC: pau k a0103 s e0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2313.lab vs ../HVITE-M/Rec-mau-M-24/mfG2313.rec
LAB: pau g e0104 N0204 j i0304 ts u0404 pau
REC: pau   e0104 N0204 p i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2314.lab vs ../HVITE-M/Rec-mau-M-24/mfG2314.rec
LAB: pau   i0104 q0204 sh    o0304 u0404 pau
REC: pau m i0103 k     o0203 m     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2315.lab vs ../HVITE-M/Rec-mau-M-24/mfG2315.rec
LAB: pau f u0104 N0204 sh    i0304 ts u0404 pau
REC: pau         m     o0103 u0203 s  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2316.lab vs ../HVITE-M/Rec-mau-M-24/mfG2316.rec
LAB: pau         r       e0103 j     a0203 a0303 pau
REC: pau y u0104 r u0204 y     a0304 k     a0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2317.lab vs ../HVITE-M/Rec-mau-M-24/mfG2317.rec
LAB: pau t a0103 m e0203 r u0303 pau

REC: pau m a0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2318.lab vs ../HVITE-M/Rec-mau-M-24/mfG2318.rec
LAB: pau k o0103 t e0203 N0303 pau
REC: pau k a0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2319.lab vs ../HVITE-M/Rec-mau-M-24/mfG2319.rec
LAB: pau k e0104 q0204 k a0304 N0404 pau
REC: pau g e0104 N0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2320.lab vs ../HVITE-M/Rec-mau-M-24/mfG2320.rec
LAB: pau i0104 y o0204 i0304 y       o0404 pau
REC: pau       k o0104 u0204 y o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2321.lab vs ../HVITE-M/Rec-mau-M-24/mfG2321.rec
LAB: pau ny o0104 u0204 b     o0304 u0404 pau
REC: pau          k     u0102 m     o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2326.lab vs ../HVITE-M/Rec-mau-M-24/mfG2326.rec
LAB: pau s e0104 e0204 t       a0304 a0404 pau
REC: pau               t e0102 r     a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2327.lab vs ../HVITE-M/Rec-mau-M-24/mfG2327.rec
LAB: pau               k o0103 t a0203 i0303 pau
REC: pau b u0104 N0204 k         a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2328.lab vs ../HVITE-M/Rec-mau-M-24/mfG2328.rec
LAB: pau k e0103 e0203 s u0303 pau
REC: pau k a0103 e0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2330.lab vs ../HVITE-M/Rec-mau-M-24/mfG2330.rec
LAB: pau y o0103 r o0203 N0303 pau
REC: pau m u0103 r o0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2331.lab vs ../HVITE-M/Rec-mau-M-24/mfG2331.rec
LAB: pau                       sh i0103 a0203 i0303 pau
REC: pau h i0105 k i0205 a0305 g  e0405 r     u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2332.lab vs ../HVITE-M/Rec-mau-M-24/mfG2332.rec
LAB: pau s o0103 u0203 i0303 pau
REC: pau d o0103 u0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2333.lab vs ../HVITE-M/Rec-mau-M-24/mfG2333.rec
LAB: pau t e0103 b i0203 k i0303 pau
REC: pau d e0103   N0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2334.lab vs ../HVITE-M/Rec-mau-M-24/mfG2334.rec
LAB: pau h a0103 n                 a0203 s u0303 pau
REC: pau         n a0104 z u0204 k e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2335.lab vs ../HVITE-M/Rec-mau-M-24/mfG2335.rec
LAB: pau ry u0104 u0204 g a0304 k u0404 pau

REC: pau k  i0104 N0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2336.lab vs ../HVITE-M/Rec-mau-M-24/mfG2336.rec
LAB: pau s a0103 m a0203 s u0303 pau
REC: pau d a0103 m a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2337.lab vs ../HVITE-M/Rec-mau-M-24/mfG2337.rec
LAB: pau s a0104 N0204 t a0304 N0404 pau
REC: pau k a0104 N0204 b a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2338.lab vs ../HVITE-M/Rec-mau-M-24/mfG2338.rec
LAB: pau         n i0103 k u0203 m u0303 pau
REC: pau m u0102 n                 e0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2339.lab vs ../HVITE-M/Rec-mau-M-24/mfG2339.rec
LAB: pau t a0103 N0203 sh o0303 pau
REC: pau g a0103 N0203 sh o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2340.lab vs ../HVITE-M/Rec-mau-M-24/mfG2340.rec
LAB: pau               k       e0101 pau
REC: pau k e0104 q0204 k o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2341.lab vs ../HVITE-M/Rec-mau-M-24/mfG2341.rec
LAB: pau       m a0102 y u0202 pau
REC: pau a0103 g a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2342.lab vs ../HVITE-M/Rec-mau-M-24/mfG2342.rec
LAB: pau sh o0104 k u0204 n i0304 N0404 pau
REC: pau k  o0104   u0204 n i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2343.lab vs ../HVITE-M/Rec-mau-M-24/mfG2343.rec
LAB: pau d a0105 i0205 j     o0305 u0405 b     u0505 pau
REC: pau         g     a0104 i0204 t     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2344.lab vs ../HVITE-M/Rec-mau-M-24/mfG2344.rec
LAB: pau t e0103 s     u0203 u0303 pau
REC: pau t       o0102 k     u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2345.lab vs ../HVITE-M/Rec-mau-M-24/mfG2345.rec
LAB: pau h o0104 u0204 d     o0304 u0404 pau
REC: pau         m     o0102 n     o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2346.lab vs ../HVITE-M/Rec-mau-M-24/mfG2346.rec
LAB: pau   m     o0102 t o0202 pau
REC: pau n o0103 N0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2347.lab vs ../HVITE-M/Rec-mau-M-24/mfG2347.rec
LAB: pau         h i0102 y u0202 pau
REC: pau s u0103 r u0203 t o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2349.lab vs ../HVITE-M/Rec-mau-M-24/mfG2349.rec
LAB: pau k i0103 s a0203 k         u0303 pau

REC: pau k         a0103 k e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2350.lab vs ../HVITE-M/Rec-mau-M-24/mfG2350.rec
LAB: pau m e0104       N0204 e0304 k i0404 pau
REC: pau k e0104 i0204 r     e0304 k i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2351.lab vs ../HVITE-M/Rec-mau-M-24/mfG2351.rec
LAB: pau          h i0101 pau
REC: pau ts u0102 g u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2353.lab vs ../HVITE-M/Rec-mau-M-24/mfG2353.rec
LAB: pau s a0104 i0204 sh u0304 u0404 pau
REC: pau k a0104 N0204 sh u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2354.lab vs ../HVITE-M/Rec-mau-M-24/mfG2354.rec
LAB: pau s a0103 k u0203 r a0303 pau
REC: pau h a0103 s a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2358.lab vs ../HVITE-M/Rec-mau-M-24/mfG2358.rec
LAB: pau d a0103 r a0203 k u0303 pau
REC: pau m a0103   i0203 g o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2361.lab vs ../HVITE-M/Rec-mau-M-24/mfG2361.rec
LAB: pau s o0104 u0204 sh i0304 k         i0404 pau
REC: pau               d  o0103 k u0203 j i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2362.lab vs ../HVITE-M/Rec-mau-M-24/mfG2362.rec
LAB: pau y o0103 y a0203 k u0303 pau
REC: pau k u0103 d a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2363.lab vs ../HVITE-M/Rec-mau-M-24/mfG2363.rec
LAB: pau   u0103 w a0203 s a0303 pau
REC: pau h a0103 s a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2364.lab vs ../HVITE-M/Rec-mau-M-24/mfG2364.rec
LAB: pau   o0104 r i0204 m o0304 n o0404 pau
REC: pau t o0104 b i0204 k o0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2365.lab vs ../HVITE-M/Rec-mau-M-24/mfG2365.rec
LAB: pau n i0104 N0204 sh i0304 k i0404 pau
REC: pau k e0104 N0204 sh i0304 k i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2369.lab vs ../HVITE-M/Rec-mau-M-24/mfG2369.rec
LAB: pau m e0104 N0204 m i0304 ts u0404 pau
REC: pau g e0104 N0204 m i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2370.lab vs ../HVITE-M/Rec-mau-M-24/mfG2370.rec
LAB: pau b a0104 i0204 t e0304 N0404 pau
REC: pau d a0104 i0204 b e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2371.lab vs ../HVITE-M/Rec-mau-M-24/mfG2371.rec
LAB: pau s a0104 b i0204 sh    i0304 i0404 pau

REC: pau s         a0103 q0203 k     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2373.lab vs ../HVITE-M/Rec-mau-M-24/mfG2373.rec
LAB: pau ts u0102 n o0202 pau
REC: pau k  u0102 m o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2374.lab vs ../HVITE-M/Rec-mau-M-24/mfG2374.rec
LAB: pau         o0104 o0204 b a0304 a0404 pau
REC: pau k o0103 t     o0203 b       a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2375.lab vs ../HVITE-M/Rec-mau-M-24/mfG2375.rec
LAB: pau ky o0104 u0204 s e0304 i0404 pau
REC: pau g  o0104 u0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2378.lab vs ../HVITE-M/Rec-mau-M-24/mfG2378.rec
LAB: pau       sh i0103 m a0203 r u0303 pau
REC: pau u0104 z  u0204 m a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2380.lab vs ../HVITE-M/Rec-mau-M-24/mfG2380.rec
LAB: pau y u0103 s u0203 g u0303 pau
REC: pau y         u0102 g e0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2382.lab vs ../HVITE-M/Rec-mau-M-24/mfG2382.rec
LAB: pau m i0103 ts u0203 d o0303 pau
REC: pau m          i0102 z o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2383.lab vs ../HVITE-M/Rec-mau-M-24/mfG2383.rec
LAB: pau n i0104 ch i0204 j o0304 u0404 pau
REC: pau g e0104 k  i0204 j o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2386.lab vs ../HVITE-M/Rec-mau-M-24/mfG2386.rec
LAB: pau ky o0104 u0204 k a0304 i0404 pau
REC: pau ky o0104 u0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2387.lab vs ../HVITE-M/Rec-mau-M-24/mfG2387.rec
LAB: pau h  i0103 g  a0203 sh i0303 pau
REC: pau ch o0103 ch i0203 k  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2388.lab vs ../HVITE-M/Rec-mau-M-24/mfG2388.rec
LAB: pau h o0102 sh         i0202 pau
REC: pau   o0103 sh i0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2389.lab vs ../HVITE-M/Rec-mau-M-24/mfG2389.rec
LAB: pau   t     a0104 i0204 s         a0304 k u0404 pau
REC: pau n a0105 g     u0205 s a0305 m e0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2390.lab vs ../HVITE-M/Rec-mau-M-24/mfG2390.rec
LAB: pau h i0103 k       u0203 i0303 pau
REC: pau         k u0102 g     i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2392.lab vs ../HVITE-M/Rec-mau-M-24/mfG2392.rec
LAB: pau k a0104 m i0204 n     a0304 r     i0404 pau

REC: pau           t     a0103 g     a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2393.lab vs ../HVITE-M/Rec-mau-M-24/mfG2393.rec
LAB: pau g a0104 q0204 k       o0304 u0404 pau
REC: pau   b     a0103 k u0203 r     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2395.lab vs ../HVITE-M/Rec-mau-M-24/mfG2395.rec
LAB: pau         y a0102 b         u0202 pau
REC: pau k u0104 r a0204 b e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2396.lab vs ../HVITE-M/Rec-mau-M-24/mfG2396.rec
LAB: pau s e0104 i0204 s a0304 k u0404 pau
REC: pau k e0104 q0204 s a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2397.lab vs ../HVITE-M/Rec-mau-M-24/mfG2397.rec
LAB: pau h         a0103 y a0203 sh i0303 pau
REC: pau h a0104 r i0204 k i0304 r  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2398.lab vs ../HVITE-M/Rec-mau-M-24/mfG2398.rec
LAB: pau sh i0104 ts u0204 g e0304 N0404 pau
REC: pau j  i0104 ts u0204   e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2399.lab vs ../HVITE-M/Rec-mau-M-24/mfG2399.rec
LAB: pau                 t     a0101 pau
REC: pau g a0104 q0204 k o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2400.lab vs ../HVITE-M/Rec-mau-M-24/mfG2400.rec
LAB: pau   k     a0103 e0203 r i0303 pau
REC: pau h i0103 d     a0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2402.lab vs ../HVITE-M/Rec-mau-M-24/mfG2402.rec
LAB: pau ky u0104 u0204 k u0304 ts u0404 pau
REC: pau    n     u0103 k e0203 r  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2403.lab vs ../HVITE-M/Rec-mau-M-24/mfG2403.rec
LAB: pau f u0102 k         u0202 pau
REC: pau f u0103 k u0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2404.lab vs ../HVITE-M/Rec-mau-M-24/mfG2404.rec
LAB: pau   o0103 N0203 n       a0303 pau
REC: pau k o0104 N0204 n a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2405.lab vs ../HVITE-M/Rec-mau-M-24/mfG2405.rec
LAB: pau       m o0102 r u0202 pau
REC: pau i0103 k u0203 r a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2406.lab vs ../HVITE-M/Rec-mau-M-24/mfG2406.rec
LAB: pau h a0103 g a0203 s u0303 pau
REC: pau n a0103 r a0203 s u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2407.lab vs ../HVITE-M/Rec-mau-M-24/mfG2407.rec
LAB: pau k         a0102 t a0202 pau

REC: pau k a0103 r a0203 d a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2408.lab vs ../HVITE-M/Rec-mau-M-24/mfG2408.rec
LAB: pau sh i0104 N0204 s a0304 ts u0404 pau
REC: pau j  i0104 N0204 k a0304 k  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2409.lab vs ../HVITE-M/Rec-mau-M-24/mfG2409.rec
LAB: pau h a0104 ts u0204 b a0304 i0404 pau
REC: pau k a0104 k  u0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2410.lab vs ../HVITE-M/Rec-mau-M-24/mfG2410.rec
LAB: pau         ry o0103 k     a0203 N0303 pau
REC: pau s u0104 t  a0204 N0304 p     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2411.lab vs ../HVITE-M/Rec-mau-M-24/mfG2411.rec
LAB: pau sh i0103 g a0203 i0303 pau
REC: pau k  u0103 r a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2412.lab vs ../HVITE-M/Rec-mau-M-24/mfG2412.rec
LAB: pau k e0104 i0204 k e0304 N0404 pau
REC: pau t e0104 i0204 d e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2414.lab vs ../HVITE-M/Rec-mau-M-24/mfG2414.rec
LAB: pau y  a0104 r i0204 k a0304 t a0404 pau
REC: pau sh i0103 r         a0203 g a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2415.lab vs ../HVITE-M/Rec-mau-M-24/mfG2415.rec
LAB: pau          h     i0103 j     o0203 u0303 pau
REC: pau u0104 ch i0204 k     o0304 m     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2416.lab vs ../HVITE-M/Rec-mau-M-24/mfG2416.rec
LAB: pau ts u0102 g i0202 pau
REC: pau k  u0102 g i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2417.lab vs ../HVITE-M/Rec-mau-M-24/mfG2417.rec
LAB: pau sh i0103 t e0203 N0303 pau
REC: pau k  i0103 g e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2419.lab vs ../HVITE-M/Rec-mau-M-24/mfG2419.rec
LAB: pau          n a0102 r u0202 pau
REC: pau ts u0103 m a0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2420.lab vs ../HVITE-M/Rec-mau-M-24/mfG2420.rec
LAB: pau sh       i0103 m i0203 N0303 pau
REC: pau sh i0104 N0204 n e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2422.lab vs ../HVITE-M/Rec-mau-M-24/mfG2422.rec
LAB: pau g     a0103 q0203 k i0303 pau
REC: pau i0103 t     a0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2423.lab vs ../HVITE-M/Rec-mau-M-24/mfG2423.rec
LAB: pau h a0104 q0204 t e0304 N0404 pau

REC: pau   m     i0103 r e0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2424.lab vs ../HVITE-M/Rec-mau-M-24/mfG2424.rec
LAB: pau k a0102 s         u0202 pau
REC: pau k a0103 s u0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2426.lab vs ../HVITE-M/Rec-mau-M-24/mfG2426.rec
LAB: pau           a0103 w a0203 r e0303 pau
REC: pau n a0104 g a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2429.lab vs ../HVITE-M/Rec-mau-M-24/mfG2429.rec
LAB: pau          s     e0102 N0202 pau
REC: pau sh i0103 N0203 g     i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2430.lab vs ../HVITE-M/Rec-mau-M-24/mfG2430.rec
LAB: pau t a0104       k u0204 sh i0304   i0404 pau
REC: pau h a0104 q0204 k          i0304 r i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2432.lab vs ../HVITE-M/Rec-mau-M-24/mfG2432.rec
LAB: pau   h     a0103 i0203 ry    o0303 pau
REC: pau d a0104 i0204 y     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2433.lab vs ../HVITE-M/Rec-mau-M-24/mfG2433.rec
LAB: pau   sh    i0103 b     a0203 r     u0303 pau
REC: pau d o0104 k     u0204 d     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2434.lab vs ../HVITE-M/Rec-mau-M-24/mfG2434.rec
LAB: pau                 k i0103 r a0203 u0303 pau
REC: pau m i0104 y a0204 b u0304 r       u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2436.lab vs ../HVITE-M/Rec-mau-M-24/mfG2436.rec
LAB: pau h a0102 k         u0202 pau
REC: pau g a0103 k u0203 h i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2437.lab vs ../HVITE-M/Rec-mau-M-24/mfG2437.rec
LAB: pau          m       e0103 i0203 sh i0303 pau
REC: pau ts u0104 m e0204 k     o0304 m  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2438.lab vs ../HVITE-M/Rec-mau-M-24/mfG2438.rec
LAB: pau k o0104 u0204 k o0304 u0404 pau
REC: pau   s     o0102 k       o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2439.lab vs ../HVITE-M/Rec-mau-M-24/mfG2439.rec
LAB: pau h i0103 k                     u0203 i0303 pau
REC: pau         k u0104 r i0204 i0304 m     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2442.lab vs ../HVITE-M/Rec-mau-M-24/mfG2442.rec
LAB: pau ts u0104 i0204 ky    u0304 u0404 pau
REC: pau          h     i0102 y     u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2443.lab vs ../HVITE-M/Rec-mau-M-24/mfG2443.rec
LAB: pau k e0104 i0204 r     e0304 k i0404 pau

REC: pau               u0103 e0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2444.lab vs ../HVITE-M/Rec-mau-M-24/mfG2444.rec
LAB: pau s u0102 n     a0202 pau
REC: pau   g     a0102 N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2447.lab vs ../HVITE-M/Rec-mau-M-24/mfG2447.rec
LAB: pau          e0104 k i0204 t a0304 i0404 pau
REC: pau ky o0104 u0204 k         a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2448.lab vs ../HVITE-M/Rec-mau-M-24/mfG2448.rec
LAB: pau h  a0103 n a0203 s         u0303 pau
REC: pau sh i0104 r a0204 s e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2449.lab vs ../HVITE-M/Rec-mau-M-24/mfG2449.rec
LAB: pau         t a0104 d a0204 sh    i0304 i0404 pau
REC: pau k a0103 t               a0203 ch    i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2450.lab vs ../HVITE-M/Rec-mau-M-24/mfG2450.rec
LAB: pau j i0103 t a0203 i0303 pau
REC: pau g i0103 k a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2452.lab vs ../HVITE-M/Rec-mau-M-24/mfG2452.rec
LAB: pau         e0101 pau
REC: pau k e0102 N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2453.lab vs ../HVITE-M/Rec-mau-M-24/mfG2453.rec
LAB: pau s o0103 sh    i0203 k     i0303 pau
REC: pau         u0103 ch    u0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2454.lab vs ../HVITE-M/Rec-mau-M-24/mfG2454.rec
LAB: pau y a0103 k o0203 u0303 pau
REC: pau           g     o0101 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2455.lab vs ../HVITE-M/Rec-mau-M-24/mfG2455.rec
LAB: pau sh         i0103 b a0203 i0303 pau
REC: pau sh u0104 k u0204 d a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2457.lab vs ../HVITE-M/Rec-mau-M-24/mfG2457.rec
LAB: pau   i0102 r                  u0202 pau
REC: pau k u0104 r u0204 sh i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2458.lab vs ../HVITE-M/Rec-mau-M-24/mfG2458.rec
LAB: pau g e0102 sh         a0202 pau
REC: pau         sh i0102 t a0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2460.lab vs ../HVITE-M/Rec-mau-M-24/mfG2460.rec
LAB: pau m i0103 u0203 ch i0303 pau
REC: pau y u0103 u0203 r  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2461.lab vs ../HVITE-M/Rec-mau-M-24/mfG2461.rec
LAB: pau   o0103 t o0203 s         u0303 pau

REC: pau n o0104 b o0204 s e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2462.lab vs ../HVITE-M/Rec-mau-M-24/mfG2462.rec
LAB: pau               r a0103 q0203 p     a0303 pau
REC: pau g a0104 N0204 r             a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2464.lab vs ../HVITE-M/Rec-mau-M-24/mfG2464.rec
LAB: pau k i0103 k a0203 k u0303 pau
REC: pau d o0103 n a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2467.lab vs ../HVITE-M/Rec-mau-M-24/mfG2467.rec
LAB: pau f u0104 g     o0204 u0304 r i0404 pau
REC: pau   k     a0103 m     o0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2468.lab vs ../HVITE-M/Rec-mau-M-24/mfG2468.rec
LAB: pau       m o0102 r     i0202 pau
REC: pau o0103 m       o0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2470.lab vs ../HVITE-M/Rec-mau-M-24/mfG2470.rec
LAB: pau h o0104 N0204 r a0304 i0404 pau
REC: pau g e0104 N0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2474.lab vs ../HVITE-M/Rec-mau-M-24/mfG2474.rec
LAB: pau b a0104 k u0204 z e0304 N0404 pau
REC: pau d a0104   N0204 z e0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2475.lab vs ../HVITE-M/Rec-mau-M-24/mfG2475.rec
LAB: pau k a0103 k o0203   i0303 pau
REC: pau n a0103 g o0203 r i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2476.lab vs ../HVITE-M/Rec-mau-M-24/mfG2476.rec
LAB: pau               w a0101 pau
REC: pau i0103 k a0203 g a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2477.lab vs ../HVITE-M/Rec-mau-M-24/mfG2477.rec
LAB: pau s a0104 ts u0204 e0304 i0404 pau
REC: pau h a0103 z  u0203 r     e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2479.lab vs ../HVITE-M/Rec-mau-M-24/mfG2479.rec
LAB: pau h o0104 u0204 d o0304 u0404 pau
REC: pau r o0104 u0204 d o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2481.lab vs ../HVITE-M/Rec-mau-M-24/mfG2481.rec
LAB: pau t o0102   o0202 pau
REC: pau g o0102 m u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2482.lab vs ../HVITE-M/Rec-mau-M-24/mfG2482.rec
LAB: pau sh i0103 m a0203 u0303 pau
REC: pau ch i0103 g a0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2483.lab vs ../HVITE-M/Rec-mau-M-24/mfG2483.rec
LAB: pau sh i0103 N0203 w a0303 pau

REC: pau m  i0103 N0203 n a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2486.lab vs ../HVITE-M/Rec-mau-M-24/mfG2486.rec
LAB: pau t a0104 y a0204 s       u0304 i0404 pau
REC: pau m u0104 r a0204 s a0304 k     i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2487.lab vs ../HVITE-M/Rec-mau-M-24/mfG2487.rec
LAB: pau m         u0102 n e0202 pau
REC: pau m o0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2488.lab vs ../HVITE-M/Rec-mau-M-24/mfG2488.rec
LAB: pau b e0103   q0203 d o0303 pau
REC: pau k e0103 r e0203 d o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2490.lab vs ../HVITE-M/Rec-mau-M-24/mfG2490.rec
LAB: pau d       e0104 sh    a0204 b a0304 r u0404 pau
REC: pau d e0105 k     i0205 a0305 g a0405 r u0505 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2491.lab vs ../HVITE-M/Rec-mau-M-24/mfG2491.rec
LAB: pau         a0103 i0203 d     a0303 pau
REC: pau k a0104 i0204 g     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2492.lab vs ../HVITE-M/Rec-mau-M-24/mfG2492.rec
LAB: pau sh o0104 u0204 s e0304 ts u0404 pau
REC: pau    ky    o0103 z e0203 ts u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2493.lab vs ../HVITE-M/Rec-mau-M-24/mfG2493.rec
LAB: pau s e0104 i0204 k e0304 ts u0404 pau
REC: pau h e0104 i0204 j i0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2494.lab vs ../HVITE-M/Rec-mau-M-24/mfG2494.rec
LAB: pau         ts u0103 m a0203 m u0303 pau
REC: pau t a0104 s  u0204 k a0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2495.lab vs ../HVITE-M/Rec-mau-M-24/mfG2495.rec
LAB: pau k o0104 u0204 k     e0304   N0404 pau
REC: pau h o0104 h     o0204 e0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2496.lab vs ../HVITE-M/Rec-mau-M-24/mfG2496.rec
LAB: pau sh i0103 b a0203 i0303 pau
REC: pau k  u0103 r a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2497.lab vs ../HVITE-M/Rec-mau-M-24/mfG2497.rec
LAB: pau g i0102 sh               i0202 pau
REC: pau         sh i0103 q0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2499.lab vs ../HVITE-M/Rec-mau-M-24/mfG2499.rec
LAB: pau j i0103 k               a0203 k u0303 pau
REC: pau         k i0104 N0204 g a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2500.lab vs ../HVITE-M/Rec-mau-M-24/mfG2500.rec
LAB: pau g e0104 N0204 k a0304 N0404 pau

REC: pau g e0104 N0204 k i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2501.lab vs ../HVITE-M/Rec-mau-M-24/mfG2501.rec
LAB: pau s u0103 k u0203 u0303 pau
REC: pau   u0102 k       u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2502.lab vs ../HVITE-M/Rec-mau-M-24/mfG2502.rec
LAB: pau       d o0102 r o0202 pau
REC: pau u0103 g o0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2503.lab vs ../HVITE-M/Rec-mau-M-24/mfG2503.rec
LAB: pau t a0104 d a0204 sh    i0304 i0404 pau
REC: pau t               a0102 k     i0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2504.lab vs ../HVITE-M/Rec-mau-M-24/mfG2504.rec
LAB: pau   s     o0103 u0203 g     o0303 pau
REC: pau d o0104 N0204 d     o0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2508.lab vs ../HVITE-M/Rec-mau-M-24/mfG2508.rec
LAB: pau   t     o0102 o0202 pau
REC: pau k u0102 r     o0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2509.lab vs ../HVITE-M/Rec-mau-M-24/mfG2509.rec
LAB: pau k e0104 q0204 sh i0304 N0404 pau
REC: pau d a0104 i0204 k  o0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2510.lab vs ../HVITE-M/Rec-mau-M-24/mfG2510.rec
LAB: pau    y     u0104 i0204 g o0304 N0404 pau
REC: pau sh i0103 n     o0203 g       u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2511.lab vs ../HVITE-M/Rec-mau-M-24/mfG2511.rec
LAB: pau   k     e0102 ch i0202 pau
REC: pau g i0103 N0203 m  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2512.lab vs ../HVITE-M/Rec-mau-M-24/mfG2512.rec
LAB: pau   z     a0102 ts u0202 pau
REC: pau b u0103 N0203 r  i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2513.lab vs ../HVITE-M/Rec-mau-M-24/mfG2513.rec
LAB: pau         d o0103 k a0203 s u0303 pau
REC: pau h i0104 t o0204 m a0304 z u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2514.lab vs ../HVITE-M/Rec-mau-M-24/mfG2514.rec
LAB: pau n o0103 b     o0203 r     u0303 pau
REC: pau g a0104 q0204 k     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2515.lab vs ../HVITE-M/Rec-mau-M-24/mfG2515.rec
LAB: pau n a0104 i0204 sh o0304 k u0404 pau
REC: pau t a0104 i0204 sh o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2516.lab vs ../HVITE-M/Rec-mau-M-24/mfG2516.rec
LAB: pau         s     o0102 b     o0202 pau

REC: pau g a0104 q0204 k     o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2517.lab vs ../HVITE-M/Rec-mau-M-24/mfG2517.rec
LAB: pau               m a0102 n e0202 pau
REC: pau a0104 b a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2519.lab vs ../HVITE-M/Rec-mau-M-24/mfG2519.rec
LAB: pau i0104 N0204 sh o0304 u0404 pau
REC: pau r     i0103 k  o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2520.lab vs ../HVITE-M/Rec-mau-M-24/mfG2520.rec
LAB: pau d o0103 n a0203 r u0303 pau
REC: pau t o0103 m a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2521.lab vs ../HVITE-M/Rec-mau-M-24/mfG2521.rec
LAB: pau h a0104 ts u0204 r a0304 ts u0404 pau
REC: pau         n  a0103 r a0203 s  u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2522.lab vs ../HVITE-M/Rec-mau-M-24/mfG2522.rec
LAB: pau k u0104 u0204 k a0304 N0404 pau
REC: pau b i0104 N0204 k a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2523.lab vs ../HVITE-M/Rec-mau-M-24/mfG2523.rec
LAB: pau f u0104 k a0204 n o0304 u0404 pau
REC: pau         k a0103 n o0203 u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2524.lab vs ../HVITE-M/Rec-mau-M-24/mfG2524.rec
LAB: pau         f u0102 k u0202 pau
REC: pau m u0103 s u0203 m e0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2525.lab vs ../HVITE-M/Rec-mau-M-24/mfG2525.rec
LAB: pau ky o0104 u0204 s e0304 i0404 pau
REC: pau gy o0104 u0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2526.lab vs ../HVITE-M/Rec-mau-M-24/mfG2526.rec
LAB: pau h i0104 k i0204 w     a0304 k     e0404 pau
REC: pau         d a0104 N0204 t     e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2527.lab vs ../HVITE-M/Rec-mau-M-24/mfG2527.rec
LAB: pau sh o0103 u0203 g o0303 pau
REC: pau d  o0103 u0203 m o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2528.lab vs ../HVITE-M/Rec-mau-M-24/mfG2528.rec
LAB: pau s a0104   i0204 g a0304 i0404 pau
REC: pau h a0104 r i0204   a0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2529.lab vs ../HVITE-M/Rec-mau-M-24/mfG2529.rec
LAB: pau         s o0103 sh i0203 k i0303 pau
REC: pau k u0104 r u0204 sh i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2530.lab vs ../HVITE-M/Rec-mau-M-24/mfG2530.rec
LAB: pau f u0104 k u0204 r a0304 m u0404 pau

REC: pau           o0103 g a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2531.lab vs ../HVITE-M/Rec-mau-M-24/mfG2531.rec
LAB: pau d a0103          r a0203 k u0303 pau
REC: pau m a0103 ts u0203 r         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2532.lab vs ../HVITE-M/Rec-mau-M-24/mfG2532.rec
LAB: pau t e0104 q0204 p     o0304 u0404 pau
REC: pau   b     e0103 q0203 d     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2534.lab vs ../HVITE-M/Rec-mau-M-24/mfG2534.rec
LAB: pau ts u0103 k e0203 r u0303 pau
REC: pau s  u0103 b e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2536.lab vs ../HVITE-M/Rec-mau-M-24/mfG2536.rec
LAB: pau         k     i0103 k     a0203 i0303 pau
REC: pau u0104 r a0204 g     i0304 r     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2537.lab vs ../HVITE-M/Rec-mau-M-24/mfG2537.rec
LAB: pau k o0104 u0204 hy o0304 u0404 pau
REC: pau d o0104 u0204 y  o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2538.lab vs ../HVITE-M/Rec-mau-M-24/mfG2538.rec
LAB: pau    h     i0103 y a0203 k u0303 pau
REC: pau ts u0104 u0204 y a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2539.lab vs ../HVITE-M/Rec-mau-M-24/mfG2539.rec
LAB: pau h a0104 N0204 e0304 i0404 pau
REC: pau   k     a0102 g     e0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2540.lab vs ../HVITE-M/Rec-mau-M-24/mfG2540.rec
LAB: pau n i0103 g e0203 r u0303 pau
REC: pau k u0103 r e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2542.lab vs ../HVITE-M/Rec-mau-M-24/mfG2542.rec
LAB: pau   k     i0103 sh o0203 u0303 pau
REC: pau t o0103 u0203 sh       o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2543.lab vs ../HVITE-M/Rec-mau-M-24/mfG2543.rec
LAB: pau f u0103 k         e0203 r u0303 pau
REC: pau         k o0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2544.lab vs ../HVITE-M/Rec-mau-M-24/mfG2544.rec
LAB: pau s a0105 m a0205 t a0305 g e0405 r u0505 pau
REC: pau         h a0104 n a0204 r e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2545.lab vs ../HVITE-M/Rec-mau-M-24/mfG2545.rec
LAB: pau sh o0104 u0204 sh       o0304 u0404 pau
REC: pau ky o0104 u0204 sh u0304 k     u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2546.lab vs ../HVITE-M/Rec-mau-M-24/mfG2546.rec
LAB: pau a0103 r     a0203 s         u0303 pau

REC: pau m     a0104 q0204 s u0304 g u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2547.lab vs ../HVITE-M/Rec-mau-M-24/mfG2547.rec
LAB: pau k         a0102 z e0202 pau
REC: pau k a0103 s e0203 g u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2548.lab vs ../HVITE-M/Rec-mau-M-24/mfG2548.rec
LAB: pau sh i0103 m e0203 r u0303 pau
REC: pau sh i0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2549.lab vs ../HVITE-M/Rec-mau-M-24/mfG2549.rec
LAB: pau ch o0104 q0204 k a0304 N0404 pau
REC: pau k  o0104 N0204 d a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2550.lab vs ../HVITE-M/Rec-mau-M-24/mfG2550.rec
LAB: pau         r i0103 y     u0203 u0303 pau
REC: pau k o0103 r       i0203 r     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2551.lab vs ../HVITE-M/Rec-mau-M-24/mfG2551.rec
LAB: pau sh i0103 r a0203 g a0303 pau
REC: pau    i0103 k a0203 g a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2552.lab vs ../HVITE-M/Rec-mau-M-24/mfG2552.rec
LAB: pau ny u0104 u0204 g a0304 k u0404 pau
REC: pau    m     i0103 g a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2553.lab vs ../HVITE-M/Rec-mau-M-24/mfG2553.rec
LAB: pau   ry    o0104 u0204 g a0304 e0404 pau
REC: pau m e0104 n     o0204 m a0304 e0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2557.lab vs ../HVITE-M/Rec-mau-M-24/mfG2557.rec
LAB: pau    s     a0102 g     i0202 pau
REC: pau ts u0103 k     a0203 i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2559.lab vs ../HVITE-M/Rec-mau-M-24/mfG2559.rec
LAB: pau m e0102 s               u0202 pau
REC: pau         s u0103 e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2561.lab vs ../HVITE-M/Rec-mau-M-24/mfG2561.rec
LAB: pau b  a0104 i0204 t     e0304 N0404 pau
REC: pau ch i0103 g     i0203 r     u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2562.lab vs ../HVITE-M/Rec-mau-M-24/mfG2562.rec
LAB: pau         t a0102 b i0202 pau
REC: pau k a0103 g i0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2565.lab vs ../HVITE-M/Rec-mau-M-24/mfG2565.rec
LAB: pau       s e0103 b i0203 r o0303 pau
REC: pau u0104 k e0204 t o0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2568.lab vs ../HVITE-M/Rec-mau-M-24/mfG2568.rec
LAB: pau e0104 N0204 ch    o0304 u0404 pau

REC: pau k     o0103 N0203 d     o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2569.lab vs ../HVITE-M/Rec-mau-M-24/mfG2569.rec
LAB: pau         sh    i0102 r     i0202 pau
REC: pau k i0104 N0204 s     e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2570.lab vs ../HVITE-M/Rec-mau-M-24/mfG2570.rec
LAB: pau n a0103 m     i0203 d     a0303 pau
REC: pau k a0104 i0204 r     a0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2571.lab vs ../HVITE-M/Rec-mau-M-24/mfG2571.rec
LAB: pau ch i0104 N0204 b o0304 ts u0404 pau
REC: pau k  i0104 N0204 m o0304 ts u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2572.lab vs ../HVITE-M/Rec-mau-M-24/mfG2572.rec
LAB: pau    ch    u0104 u0204 d o0304 k u0404 pau
REC: pau sh i0104 r     i0204 z o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2573.lab vs ../HVITE-M/Rec-mau-M-24/mfG2573.rec
LAB: pau s e0104 q0204 k a0304 k u0404 pau
REC: pau s e0104 N0204 t a0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2574.lab vs ../HVITE-M/Rec-mau-M-24/mfG2574.rec
LAB: pau sh i0103 k a0203 k u0303 pau
REC: pau sh i0103 k a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2575.lab vs ../HVITE-M/Rec-mau-M-24/mfG2575.rec
LAB: pau   s     o0102 b     a0202 pau
REC: pau g o0103 r     a0203 N0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2576.lab vs ../HVITE-M/Rec-mau-M-24/mfG2576.rec
LAB: pau    i0103 b a0203 r u0303 pau
REC: pau sh i0103 b a0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2578.lab vs ../HVITE-M/Rec-mau-M-24/mfG2578.rec
LAB: pau t  o0102 ch         i0202 pau
REC: pau ch o0103 ch i0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2580.lab vs ../HVITE-M/Rec-mau-M-24/mfG2580.rec
LAB: pau   m     u0102 k i0202 pau
REC: pau k u0103 i0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2581.lab vs ../HVITE-M/Rec-mau-M-24/mfG2581.rec
LAB: pau         k o0104 u0204 r i0304 ts u0404 pau
REC: pau m u0102 k                        u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2582.lab vs ../HVITE-M/Rec-mau-M-24/mfG2582.rec
LAB: pau k a0104 ts u0204 y o0304 u0404 pau
REC: pau t a0104 d  a0204 y o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2584.lab vs ../HVITE-M/Rec-mau-M-24/mfG2584.rec
LAB: pau   s     a0104 q0204 s u0304 r u0404 pau

REC: pau h i0103 t     a0203 s         u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2585.lab vs ../HVITE-M/Rec-mau-M-24/mfG2585.rec
LAB: pau   u0104 N0204 m e0304 i0404 pau
REC: pau b u0104 N0204 m e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2586.lab vs ../HVITE-M/Rec-mau-M-24/mfG2586.rec
LAB: pau b o0104 u0204 ry o0304 k u0404 pau
REC: pau d o0104 u0204 ry o0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2587.lab vs ../HVITE-M/Rec-mau-M-24/mfG2587.rec
LAB: pau i0104 ch    i0204 i0304 ch i0404 pau
REC: pau d     a0104 i0204 i0304 ch i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2588.lab vs ../HVITE-M/Rec-mau-M-24/mfG2588.rec
LAB: pau b     u0103 q0203 k a0303 pau
REC: pau a0103 t     e0203 n a0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2589.lab vs ../HVITE-M/Rec-mau-M-24/mfG2589.rec
LAB: pau          h a0102 b a0202 pau
REC: pau sh a0103 g a0203 m u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2590.lab vs ../HVITE-M/Rec-mau-M-24/mfG2590.rec
LAB: pau t e0104 q0204 k i0304 N0404 pau
REC: pau g a0104 i0204 j i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2591.lab vs ../HVITE-M/Rec-mau-M-24/mfG2591.rec
LAB: pau y  o0103 h o0203 d o0303 pau
REC: pau ry o0103   u0203 d o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2592.lab vs ../HVITE-M/Rec-mau-M-24/mfG2592.rec
LAB: pau   s     a0102 g i0202 pau
REC: pau k a0103 N0203 b u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2594.lab vs ../HVITE-M/Rec-mau-M-24/mfG2594.rec
LAB: pau       e0102 r u0202 pau
REC: pau u0103 e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2595.lab vs ../HVITE-M/Rec-mau-M-24/mfG2595.rec
LAB: pau k o0104 k u0204 r i0304 ts u0404 pau
REC: pau                 r i0102 ts u0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2597.lab vs ../HVITE-M/Rec-mau-M-24/mfG2597.rec
LAB: pau h       a0103 i0203 ry    o0303 pau
REC: pau h a0104 i0204 ky    u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2598.lab vs ../HVITE-M/Rec-mau-M-24/mfG2598.rec
LAB: pau h a0104   ts    u0204 i0304 k u0404 pau
REC: pau s a0104 b i0204 r     e0304 r u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2599.lab vs ../HVITE-M/Rec-mau-M-24/mfG2599.rec
LAB: pau   u0102 k         u0202 pau

REC: pau b a0103 k u0203 r o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2600.lab vs ../HVITE-M/Rec-mau-M-24/mfG2600.rec
LAB: pau t e0104 q0204 k  a0304 k u0404 pau
REC: pau   j     i0103 sh a0203 k u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2601.lab vs ../HVITE-M/Rec-mau-M-24/mfG2601.rec
LAB: pau           a0102 m e0202 pau
REC: pau b a0103 r i0203 k i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2602.lab vs ../HVITE-M/Rec-mau-M-24/mfG2602.rec
LAB: pau           a0103 ts u0203 i0303 pau
REC: pau g a0104 s o0204 r  i0304 N0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2603.lab vs ../HVITE-M/Rec-mau-M-24/mfG2603.rec
LAB: pau   a0104 q0204 p a0304 k u0404 pau
REC: pau h a0104 N0204 p u0304 k u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2604.lab vs ../HVITE-M/Rec-mau-M-24/mfG2604.rec
LAB: pau k o0104 u0204 s e0304 i0404 pau
REC: pau g o0104 u0204 s e0304 i0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2606.lab vs ../HVITE-M/Rec-mau-M-24/mfG2606.rec
LAB: pau sh o0104 u0204 sh o0304 u0404 pau
REC: pau ky o0104 u0204 ch o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2607.lab vs ../HVITE-M/Rec-mau-M-24/mfG2607.rec
LAB: pau m  i0103   e0203 r u0303 pau
REC: pau sh i0103 g e0203 r u0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2608.lab vs ../HVITE-M/Rec-mau-M-24/mfG2608.rec
LAB: pau k e0104 i0204 s o0304 ts u0404 pau
REC: pau s e0104 q0204 t o0304 k  u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2610.lab vs ../HVITE-M/Rec-mau-M-24/mfG2610.rec
LAB: pau g a0104 N0204 j o0304 u0404 pau
REC: pau k a0104 i0204 j o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2611.lab vs ../HVITE-M/Rec-mau-M-24/mfG2611.rec
LAB: pau h o0102 r               i0202 pau
REC: pau k u0104 r i0204 i0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2612.lab vs ../HVITE-M/Rec-mau-M-24/mfG2612.rec
LAB: pau y o0104 q0204 ky u0304 u0404 pau
REC: pau y o0104 u0204 ky u0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2613.lab vs ../HVITE-M/Rec-mau-M-24/mfG2613.rec
LAB: pau f u0103 a0203 N0303 pau
REC: pau   b     a0102 N0202 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2616.lab vs ../HVITE-M/Rec-mau-M-24/mfG2616.rec
LAB: pau n i0103 k u0203 m u0303 pau

REC: pau g i0103   N0203 m i0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2617.lab vs ../HVITE-M/Rec-mau-M-24/mfG2617.rec
LAB: pau   n     e0102 r u0202 pau
REC: pau g e0103 N0203 g o0303 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2618.lab vs ../HVITE-M/Rec-mau-M-24/mfG2618.rec
LAB: pau          y a0103 k e0203 d o0303 pau
REC: pau ts u0104 m e0204 k o0304 m u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2619.lab vs ../HVITE-M/Rec-mau-M-24/mfG2619.rec
LAB: pau ky u0104 u0204 k u0304 ts    u0404 pau
REC: pau ry u0104 u0204   ch    o0304 u0404 pau
Aligned transcription: ../HVITE-M/Rec-mau-M-24/mfG2620.lab vs ../HVITE-M/Rec-mau-M-24/mfG2620.rec
LAB: pau s u0102 k i0202 pau
REC: pau b u0102 k i0202 pau
====================== HTK Results Analysis =======================
 Date: Tue Nov 18 12:25:41 2003
 Ref : mlf_scriptmau-M
 Rec : ../HVITE-M/Rec-mau-M-24/mfG0002.rec
     : ../HVITE-M/Rec-mau-M-24/mfG0003.rec
     : ../HVITE-M/Rec-mau-M-24/mfG0004.rec
     : ../HVITE-M/Rec-mau-M-24/mfG0005.rec
     : ../HVITE-M/Rec-mau-M-24/mfG0006.rec
------------------------ Overall Results --------------------------
SENT: %Correct=25.39 [H=666, S=1954, N=2620]
WORD: %Corr=65.28, Acc=56.13 [H=13326, D=1077, S=6010, I=1868, N=20413]
===================================================================